A Complete Bibliography of Publications in Reviews of Modern Physics (1960–1969)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.06

Title word cross-reference

(α, 2α) [BHK^{+}65]. (d, n)2α [JBJ^{+}65]. (∞, ν) [GP64]. (p, 2p) [PB65, Rio65]. (p, 2p)n [BKF65]. (p, 3α) [Swa65]. (p, α)2α [KKR65]. (p, pd) [DSF65]. + [KK65]. +d → 2α + n [BB65]. 0.25 [VBB68]. 0.5 [Duc67]. 1.2 [Duc67]. 10.5 [PB65]. 135 [EGLM65]. 1500 [Moy61]. 163 [Deh65]. 1s^{2} [RSW60]. 1s^{2}2s [RSW60]. 1s^{2}2s^{2} [RSW60]. 210 [Tho67]. 2p [AWY65]. 31 [BKF65]. 310 [RCF^{+}61]. 4 [Mey65]. 5 [VBB68]. 50 [Hos67b]. 500 [Moy61]. 600 [Kaz67a]. 650 [Kaz67a]. 675 [KKR65]. ^{0} [EFG^{+}61]. ^{1} [SK63]. ^{10} [AD65, EWM^{+}65, WEHM65]. ^{11} [BLBM63, KKR65, Swa65]. ^{11}B(p, α)2α [Deh65]. ^{12} [AD65, Phi65, YH65]. ^{12}C → 3α [Deh65]. ^{13} [AD65]. ^{1}H(d, 2p)n [SBJP65]. ^{2}H(p, 2p)n [SBJP65]. ^{3} [AD65, AWY65, BT65, BB62, EWM^{+}65, GO65, MEHW65b, MEHW65a, WEHM65, YJE^{+}65]. ^{4} [AWY65, BT65, DSF65]. ^{40} [BHK^{+}65]. ^{6} [DSF65, NMC65, YJE^{+}65]. ^{62} [PB65]. ^{63} [PB65]. ^{7} [BB65, JBJ^{+}65]. ^{8} [AD65, MEHW65b]. ^{9} [MEHW65b, MEHW65a]. _{1} [Cal63]. _{2} [FB60c, HS63, Hij60, Hon68, Hyl68, KR60a, PBS68, Ric60, TB69, VBB68]. 2,3 [Cal63]. _{2}^{+} [Hij60]. _{3} [CCM64, CC64, Gol64, Hon68, Weg64]. _{4,5} [Cal63]. NN [BBD^{+}67, Duc67]. a + A → C^{∗} → b_{1} + b_{2} + b_{3} [Kra65]. α [AAA69, AD65, MEHW65b, Phi65, YH65]. ααα [MEHW65a, WEHM65]. d [Bec64]. e/h [TPL69, TPL73]. ep [MT69]. I = 1 [Hel67]. K_{2}^{0} + p → Λ^{0} + Π^{+} + Π^{0} [Ada61]. Λ^{0} [EFG^{+}61]. μp [MT69]. n [AH68, BY64]. ν_{e} [LS64a]. p [BT65, Cat67, GL61, Hos67b, Mac67, Moy61, RCF^{+}61, SK62, YH65, GL61]. pαα [YJE^{+}65]. pααα [EWM^{+}65]. Π [GL61, DP63]. Π^{+} [RCF^{+}61, SK62]. Π^{−} [Moy61]. R [DW64]. → [KK65]. → 3α [NMC65]. → αα [MEHW65b]. S [Che62, CD67, Sta62, Hor68]. SU(3) [Bar68]. SU_{3} [MJC64, Mos62].

-Band [Hor68]. -Electron [DP63]. -Group [Bec64]. -keV [Deh65, KKR65]. -Matrix [Che62, CD67, DW64, Sta62]. -Type [AH68].

1 [Per66, SK63]. 1959 [Par60]. 1963 [Ano63, Löw63a].

2 [NMC65].

32 [Tay61].

Abelian [Sch64d]. Above [Arn67, Kaz67b]. Abrikosov [AK64]. Absorption [Azá63, GL64, Mar64a, Ost61, Sho67]. Absorption-Line [Sho67]. Abstracts [Ano67a, Ano67b, Ano67c, Ano67d, Ano67e]. Ac [SMR64]. Accelerated [Rom63]. Accelerator [Mar61a, Mar66, Mar61b]. Accurate [KR60a]. across [Bur60]. Activity [Was60]. Addendum [FC69]. Adiabatic [Ago60, Pre63]. Adjoint [CJ67, Jos67, Jos68]. Adsorbed [Man67]. Advances [McW60]. Ages [And62]. Air [Bol64]. Algebra [Löw67]. Algebras [GG69]. Alkali [LA68, VS63]. Alkali-Noble [LA68]. Alloy [LCW68, Mei64]. Alloys [Bec64, BHMM64, Col64, GL64, JB64, JHC64, KY68, LA68, MCJR64, Mat61, Par64, SCGP64]. Alpha [AD65, Duc65, EP61]. Alpha-Particle [AD65]. Alternant [De 63]. Alternating [Par64]. Alternation [THH60]. Aluminum [HK64]. Ammonia [Lor60, MJ68, Tho68b]. among [Ara68]. Amorphous [MF64b]. Amplitude [HLS60, Kaz67b, Pai60]. Amplitudes [Ama67]. Analogs [SF64]. Analogy [Bro60a]. Analyses [Ada65, Tam65]. Analysis [Ara60, BKR62, DSF65, Hos67a, Jac65a, KLJ67, Kaz67b, Mac62, MHB67, Mos60, PP63, Per67, RCF^{+}61, Sch61, Sch68, Sin63b, VK68]. Analytical [Mac69, RSW60]. Angle [BT64, CN68, Sco63]. Angular [CCM64, Duc65, Kra65, Löw64, MEBL67, MEHW65b, RB67b, She63, SV66, SBJP65]. Anisotropic [Kap60]. Anisotropy [DP64]. Annihilation [GL61]. Announcement [Ano68]. Anomalies [HR68]. Anomalous [Bai68, Cut61, Gol64, KY68]. Antimatter [Alf65]. Antinuclear [Phi67]. Antiproton [Lyn61]. Antiproton-Proton [Lyn61]. Any [NG68].

Apparent [KA65]. Application [De 63, KW63, Shi64]. Applications [GL60, Löw67, Mos62]. Approach [Che61]. Approximate [Mus67, Nes61]. Approximately [Boy60]. Approximation [BvdH62, Cas64, Coo67, DW64, HM65, LPLH63, Mey60]. Approximations [Mul60]. Arbitrary [RM60]. Areas [ST65]. arranged [Löw63a]. Aspects [Bre62, Mar64a, Mös64, PP60, Sel65, Tho67, Bre63]. Assisted [KTB64]. Associated [Han68]. Association [FT60, GL60, KI60]. Associations [Bir64, Pre63]. Astrophysics [Bie60, Dal67a, Hue61]. Atmosphere [Cho61b, Jac63, Cho61a]. Atmospheres [Dal67b]. Atmospheric [Bol64]. Atom [BI66a, BI66b, Dal67a]. Atom-Atom [Dal67a]. Atomic [BB67, Con68a, Dal63b, ER63, FC68, FC69, FJMS66, Ger61, Hur63, I’H63, KD66, Löw63a, MPS63, MOK64, Mul60, Pre63, RSW60, SB67, Sla63a]. Atoms [Ara60, BS62, Gom63, Gre63, Hyl63, KD66, KS63b, McW63, MS68, RW64, Rud68, SK63, Sla63b, Tho63, MS69]. Atoms-The [Sla63b]. Attached [Cab60]. Attenuation [LS64b]. Auroral [Ves60]. Average [BvdH62, Mol64]. Axial [Hei64]. Axiomatic [Jaf69, Sta62]. Axiomatics [Bun67]. Azines [GK63b, GK63a].

B [KKR65, AD65, Ano67a, EWM^{+}65, Str64, Swa65]. Balance [Leh60]. Band [Bec64, Gol64, Hor68, LCW68, LA68, SPB69, Vas68, Wan62, Weg64]. Band-Edge [LCW68]. Band-Theoretic [LA68]. Bands [CP68, KC68]. Baryons [Nie68]. Based [DW64]. Basis [AK60, HVW65, Sch64a]. Be [AD65, Boy60, MEHW65b, BB66b, Dir62, MEHW65b, MEHW65a, WEHM65]. Bearing [Pol62]. Behavior [Bai68]. below [KKR65]. Berkeley [Che67]. Beta [Dan68, RHW68, Wei61, Wu64]. Beta- [RHW68]. Beta-Ray [Dan68]. Between [GBG64, Adk64, Alf60, Bie60, Duc67, Ger68, Hue61, JS61, JS62, KP64b, Row64, Van62, VBB68, Wad64, Whe62]. Bi [LCW68]. Bimesityl [Was60]. Binary [SD61, SM65]. Binding [CSR^{+}62, Hur60, VS63]. Biochemistry [PK63, PP60]. Biological [Löw63b]. Biophysics [Kot63, PK63]. Birthday [Ano62, DPSY64]. Bodies [Cho61a, Cho61b, Ima60, Sea60, Ste60]. Body [Bra67, Deh65, DP69, Lud60, Nie68, Phi64, RB67a, Sch61]. Bogoliubov [ST65]. Bohm [Tut68]. Boltzmann [HV66]. Bombs [Bel67]. Bond [FM60, HS63, HNLD63, Rue62, THH60]. Born [Bor64, HM65]. Boston [Nes60]. Boulder [Par60]. Bound [Mar64a, MS64a]. Boundary [De 64, Jun60, Nap60]. Bounds [BF63, Wei63]. Brackets [Jos64b]. Bragg [SPB69]. Brain [Blo62]. Branching [SPRM61]. Breakup [Phi64]. Bremsstrahlung [BvdH62, Pal68]. Brillouin [Gri68, Gri69, Mer64]. Broadened [Sto69]. Broadening [Coo67]. Brueckner [Day67]. Bub [Tut68]. Buckling [Fri60].

C [AD65, AD65, Ano67b, BBD^{+}67, BLBM63, Duc67, Phi65, YH65]. Ca [BHK^{+}65]. Calculated [Str64]. Calculating [GY69, Her62]. Calculation [BLBM63, Chr64, FB60b, Pre63, Tam65, War65]. Calculations [AK60, Bar63, Cou60, FB60c, Fre60, FKC60, GBD60, Hur63, I’H63, Nes63, RÖ63, VM60]. Calibration [Mar66]. Calibrations [Mar61a, Mar61b]. Can [BB66b]. Canonical [And64, FB60a]. Capacity [PLG64]. Capture [Ber68, MG65, RF60]. Carathéodory [Red68]. Carbon [HW67, HW68, Hur60]. Cerenkov [Bol64]. CERN [Wet61]. CH [FB60c]. Chain [DW62]. Chalcogenide [KY68]. Chance [Haw64]. Changes [Boy60]. Channel [Tam65]. Channels [Lyk60]. Character [LA68]. Characteristics [Sca64]. Characterization [Mor64]. Charge [Ing61, LS64a]. Charge-Conjugation [Ing61]. Charged [Bro60b, Kle68, Sco63]. Charging [Bai68]. Chemical [BC60, HS63, HNLD63, Pit63, Rue62, Sch68]. Chemistry [CD65]. Chew [Car65]. Chew-Low [Car65]. Chiral [GG69]. Choice [Haw64]. Classical [JS61, JS62, Lax66a, Lax66b, Pla61, Rue64, ST65]. Classification [Bar68]. Clebsch [dS65, MJC64, dS63]. Close [AB65]. Close-Packed [AB65]. Closed [HL68]. Closed-Shell [HL68]. Cluster [Bra67, Phi64, Rue64]. CO [Ric60, TB69]. Coefficient [Eyr62, GP64]. Coefficients [EHD69, MJC64, Mos62, TH60, dS63, dS65]. Coherence [MW65, TPL69, TPL73]. Collapse [Tut68]. Collective [BS67, DM64, Tam65]. Collision [Alf60, Dal67a, Dal67b, DW64, FKP60]. Collision-Free [FKP60]. Collisions [BAC61, SK62, Van64]. Colorado [Par60]. Column [SR62]. Combinations [Mul60]. Comet [Hue61]. Commencements [Fer60]. Comment [EH68, Sle60, Wil60, Hue61]. Comments [BB68, Max69]. Comparative [SB67, SLV62]. Comparison [Mey65, Str64]. Competition [PB65]. Compilation [CSR^{+}62]. Complete [BC60]. Completely [Bur60]. Completeness [Ing61]. Complex [Str66]. Composite [SDH64]. Compound [Bre64, Fow64a, Mol64, Por64, SV66]. Compound-Nucleus [Mol64]. Compounds [PP64]. Compression [Kol60]. Computer [Nes63]. Concept [Sew64, Yan62]. Conceptual [HVW65, Tis63]. concerning [Ada61, BDRA60, Lor60]. Conclusions [Mat61]. Condensations [Bir64]. Conditions [And64, Dir62]. Conducting [Has60, Ima60, MF64a, Ste60, Zhi60]. Conduction [KP64b]. Conductivity [Car61, CC64, DLLR64, KT64, Mcc66, Ost61, Res60, SG63, Tay60]. Conference [Ano64k, BS64, Fow64a, GMW67, Par60]. Configuration [MWY60]. Configurational [FB60a]. Configurations [FB60c, RSW60]. Confinement [Leh60]. Conformal [FRW62]. Conjugated [Dew63, Lie60, Rue63]. Conjugation [Ing61]. Consequences [Jor62, Tin64]. Conservation [Fle60a, Fle60b]. Conserved [Wu64]. Considerations [Ada61, And66, War68]. Consistent [KC68, LF67, Mul60, RSW60, Roo60]. Constants [BDRA60, CD65, Köp67, Löw62b, MJ68, TPL69, TPL73]. Constructed [Löw64]. Construction [Boy60]. Containing [SMR64]. Continuous [Str64]. Contributed [Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano64h, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e]. Contributions [Dre61a]. Convection [Nak60]. Conventional [GMGL64]. Convergence [Gil63]. Conversion [Erb66]. Convolution [Löw67]. Coordinate [And64].

Coordinates [Str66]. Core [BLBM63]. Corepresentations [BD68]. Corrections [Frö60, GY69, Max69, MT69]. Correlated [Gil63, KR60b, RW60]. Correlation [ACG63, CP68, Cat67, Con68a, GBG64, Hor68, Löw62a, MEHW65b, MP63, Rue64, She63, Sil64, Sla63b, SSW63]. Correlations [Ara68, GL61, Ing61, Jac65b, KC68, Lit64, SBJP65]. Cosmic [Bol64]. Cosmic-Ray [Bol64]. Cosmology [Dic62]. Coster [Cal63]. Coulomb [FM60, KRS60]. Coupled [BKR62, Jos64a, Tam65]. Coupled-Channel [Tam65]. Coupling [GK63a, GK63b, Köp67, Ric60, Tal62]. Couplings [Her62]. Covariance [Jos65]. Covariant [Fle60a, Fle60b]. Coverages [Man67]. Critical [Bal64a, Bar62a, BGR64, CCM64, DVB64, EL64, GC64, JHC64, KGH^{+}67, Mei64, RAG64, Tho67, Tox64]. Critique [FT60, Her62]. Cross [Bel67, FVV61, HW67, HW68, KD66, Mol64, Moo60, RCF^{+}61, Wil64]. Cross-Section [KD66, Moo60, RCF^{+}61]. Crystal [Ewa65, Mar64a, MV68, War68]. Crystals [Aiz62, AS60, BC64, Gri68, Gri69, Hoc62, MS64a, Mus67, Pal68, Sla65, Wal65, Zer67]. Cu [PB65]. Cubic [AC65]. Current [EL64, Fis64, Kun61, Mei64, Ros60, Wu64]. Currents [Bai68, Bar62a, CCM64, GC64, Mit64, SJH64, SR64c]. Curved [Fro65]. Curves [CH60, Cun64, KLS64, Mos60, VS63]. Cusp [WSL^{+}61].

D [Ano67c, BKF65, KK65]. Damping [Wad64]. Data [BLBM63, Bre67, KLJ67, Kaz67b, KD66, MG65, PP63, RHW68, RBGB^{+}64, RBGB^{+}65, RBGP^{+}67, RBSBG^{+}68, TB69]. Decay [Deh65, Jac65b, Kap60, Phi65, SBJP65, Swa65, Wei61, Wu64]. Decays [Nie68]. Decomposition [BT64]. Decompositions [IN66]. Defects [KT64, Liv64]. Defined [RB67b]. Definitions [Bre64]. Deformations [MPS63, Pip64a]. Deformed [Dav65, She60a]. Degenerate [EL68]. Delay [Eis64, Phi64]. Dense [EHD69, HF68]. Densities [Kun61]. Density [And63, BDRA60, CLS63, Col63, Coo67, Hub61, McW60]. Dependence [CCM64, Mei64, RAG64, Sil64, Tox64]. Dependences [Fis64]. Dependent [Ber60a, KA65, MB64]. Deposited [MF64a]. Derivation [Ito65]. Derivatives [Was60]. Described [BM68]. Determination [Bai68, HW63, Kaz67b, Mac67, TPL69, TPL73]. Determining [LS64a]. Deuteron [BAC61, Car65, EGLM65]. Deuterons [Udo65]. Development [Alf65]. Developments [CNO64, Dre61b, DW62, Hit69, Kie64]. Diagnostics [Ber60b]. Diagonal [Yan62]. Diagrammatic [War65]. Diatomic [KD66, KW63, KMKI60, Nes60, Ran60a, Ran60b]. Dielectric [Bai68, MJ68, O’K63]. Different [De 63, Men64a, Mey65, Mos63]. Differential [Whe62]. Diffraction [BC64, Cas64, PP63, SPB69]. Diffusion [Hoh62, SD61, SR62, SM65, SM66, TH60]. Dilute [MCJR64]. Dimensional [BM68, Hav64, Jun60, SML64, She60b, SPB69]. Dirac [Jor62]. Direct [Deh65, PM64, Ros60, Sch65]. Direct-Current [Ros60]. Discharge [Ber60b]. Discharging [Bai68]. Discrete [Mar62]. Discrete-Symmetry [Mar62]. Discussion [CLS63, För63, Gre63, Har60, HNLD63, Hir63, I’H63, JN63, LPLH63, Mat63, McW63, Mos63, O’K63, OA63, Pit63, PK63, Rue63, Sin63a, Sla63a, Wei63]. Disintegration [Car65, Udo65]. Dispersion [Fub61, HW63, Lyt65, Mos60, Sch65, SPB69]. Distances [Her62, Tam67]. Distorted [HM65]. Distorted-Wave [HM65]. Distribution [BC62, FC68, FC69, Mou66]. Distributions [Duc65, RB67b, SV66]. Divalent [Vas68]. DNA [Löw63b]. Does [SC64]. Domain [Aha62]. Done [Mic67]. Double [BAC61]. Drag [Jac63]. Due [KA65, HR60, KT64]. Dynamic [Bar62c, BNM^{+}64, Gra60, Lüs60, Nap60, SW60, Tat60, Zhi60]. Dynamical [BC64, Che61, CMS67, HL60, JS61, JS62, Mar64a, Pai66]. Dynamically [Mcc66]. Dynamics [Cha60, Ger60, Jun60, She60b, Spi60, Tay61, Wal65, Wan62].

Early [Hyl63]. Earth [CNO64, Ell64, Kie64, SMR64, Jac63]. Earths [Mös64]. Edge [LCW68]. Edges [Azá63]. Editorial [Ano69, BBJ^{+}69]. Effect [Ano64k, BGR64, Bro60a, Des64, Ell64, PP64, Pos67, RW64, SJH64, SF64, Wad64]. Effective [GG69, GY69, Tal62]. Effects [Bal64a, CP68, De 64, HJ62, Jas60, Jér68, KB67a, KB67b, Löw62a, Mer64, Mit64, Sei62, SCGP64, SSW63, Tol63, Zhi60]. Egil [Löw63a]. Einstein [Bas65]. Elastic [Mac67, MT69, Moy61, YH65]. Elasticity [Wal65]. Electric [Ber60b, Bur60, HMW63, Wan62]. Electric-Discharge [Ber60b]. Electrical [Hon68, Kle62, MF64b, Tay60]. Electrodynamic [KB67a, KB67b, Pla61]. Electrodynamics [CN67, Ito65, TPL69, TPL73]. Electromagnetic [CD67, Erb66, Fin64, Ros60, Sch64d]. Electron [Ber68, BT64, BvdH62, CN67, DP63, FG61, Frö60, GP66, GO65, HS63, HW67, HW68, Hyl63, Kes69, KD66, KS63b, LI66, MS68, MOK64, Pal68, RÖ63, RF60, Rud68, SK63, Sch64c, Sin63b, SPB69, MS69]. Electron-Diffraction [SPB69]. Electron-Volt [HW67]. Electronic [BDRA60, BHMM64, CH60, Col64, FKC60, Gol60, GBD60, Gre63, Hur60, Hyl68, KR60a, Kot63, Nes63, PP60, Roo60, Sla63b, Weg64]. Electrons [Ara68, BS62, DEM68, Ito65, KP64b, MP63, Phe68, Wan62]. Element [Bec64, BY64]. Elementary [Che62, Fro65, JF62, Mel60, Mel62, Roo63, RBGB^{+}64]. Elements [Day67, JHC64, RB67b]. Elimination [Hur63]. Embedded [Ne’65]. Embedding [Fri65, RN65, Ros65]. Emission [JBJ^{+}65, Ost61]. Emitting [AD65]. Emphasis [Ham62]. Empirical [SLV62, Wil64]. Energies [ACG63, Bre67, Che67, CSR^{+}62, Gol60, Hel67, HW67, KLJ67, Lea67, Mar66, Sel65, HW68]. Energy [Bar68, BB67, BB65, BS62, Cal63, CP68, Che61, Chi69, Ded62, Ded63, DP64, Dom69, Dre61a, Erb66, FM60, FKC60, GO65, GY69, Hij60, Hur60, JF62, KA65, Kaz67b, KC68, Kes69, Köp67, Lan60, Leh60, LCW68, Mar61a, Mar61b, Moy61, Pal68, Phe68, Pre63, RW64, SDH64, Ser64, Sim67, SBJP65, SPB69, Tat60, Van64, VS63, Vas68, Vog62, Wad64, Wil64, Wip64]. Energy-Angular [SBJP65]. Energy-Band [LCW68]. Energy-Dependent [KA65]. entire [Tay61]. EPR [SSW63]. Equation [Bab63, Bal60, Pra63]. Equations [BM68, HL60, Mac66, Mac69, Mar64b, Pla61, Row68]. Equations-of-Motion [Row68]. Equilibria [FR60]. Equilibrium [Ago60, CFH60, Mcc66, Wol60]. Equivalence [Fre60, ST65]. Errata [BW61, Cho61a, Ded63, JS62, Mar61b, MGC63a, Mel62, Tay61]. Erratum [AM67, Ano64i, BDFL62a, Bre63, Bur63a, CN64, CJS63a, Fle60a, Fow64b, GK63a, Gri69, HW68, KB67a, KS64, MS69, Sci64a, TPL73, Web60, dS65]. Error [BF63]. Errors [Hur63]. Eu [KY68]. Eu-Chalcogenide [KY68]. Euclidean [Fri65, Ros65]. Eugene [Ano62]. Euler [BT64]. Euler-Angle [BT64]. Eutectic [SCGP64]. eV [VBB68]. Evaluation [KP66]. Even [Eic64, SV66, She60a]. Even- [SV66]. Even-Even [Eic64, She60a]. Events [BBS63]. Evidence [BBS63, PM64, RHW68]. Evolution [Dic62]. Exact [GL60, Mar64b, RM60]. Exchange [Löw62a]. Excitation [MS68, MS69, MOK64, Phe68]. Excitations [JF62, VK68]. Excited [CH60, CD63, KS63b, Mey65]. Excluded [BB66b]. Existence [Coh64a]. Expansion [Löw60]. Expansions [Bra67, HNLD63]. Expectation [BF63, Wei63]. Experiment [EFG^{+}61, KGH^{+}67, LF67]. Experimental [DS68, Don65, Kaz67a, Mac69, Mic67, RF60, Tho67, KD66]. Experiments [Che67, CFH60, Cso65, DH60, Lap67, RCF^{+}61, Sco62, Sim67, Sle60, Wil59, Wil60]. Explicit [Mar62]. Extended [Azá63, Löw60, Row68]. External [Pai60]. Extrapolation [Car65].

F [Ano67e, Hij60]. Factors [BvdH62, HMW63, HV66, SD67]. Far [GL64]. Fast [Cin66, Sco63]. Fast-Neutron [Cin66]. Fermi [Sch63, Tel62, Vas68, Wu64]. Fermion [And63, CLS63, Col63, Cso65, Nes61]. Fermion-Fermion [Cso65]. Fermions [SML64]. Ferrihemoprotein [GBG64]. Ferroelectricity [Aiz62]. Ferromagnetic [SMR64]. Few [Don65]. Few-Nucleon [Don65]. Field [Bea64, Bro60b, Bur60, CCM64, CS64, Des64, Dir62, DH60, DK62, DEM68, EL64, Fin64, Fis64, GG69, GMGL64, HLS60, Hak64, Hei64, Hoh62, Jaf69, JHC64, KM60, Kle68, MS64b, Mul60, Nak60, NT64, Pai60, Par64, RSW60, Roo60, RAG64, Sch60, Sch64d, Sew64, Wer64]. Fields [Bal64a, Bar62a, BW57, DVB64, Erb66, Kun61, MW65, Mar64b, Mar62, Sch64d, Tay60, Tox64, Udo65, Wan62, BW61]. Film [Man67]. Films [Bal64a, Dou64, GC64, HT64, LC64, Mei64, MF64a, MKQ64, Mit64, NT64, SR64a, Tin64, Tox64]. Final [Ada61, Ada65, BT65, Lyn61, Sel65]. Final-State [Ada61, Ada65, BT65, Sel65]. Fine [Azá63]. Finite [BM68, HLS60, Pip64a, Res60, Tay60]. Finite- [BM68]. Finitely [Zhi60]. First [Fer60, Ran60b, Wei61]. First-Forbidden [Wei61]. First-Row [Ran60b]. Fits [Bre67]. Fixed [BT64, Lyt65]. Fixed-Nucleus [BT64]. Flight [Udo65]. Florida [Löw63a]. Flow [And66, HR60, Lud60, Lyk60, Mey60, Ros60, Sea60]. Flows [Gra60, Ima60]. Fluctuations [Lax60]. Fluid [Ger60, Gra60, Has60, Jun60, Lüs60, Lyk60, Mou66, Nap60, Spi60, Tat60, Tay60, Tay61, Zhi60]. Fluids [Ger68, Ima60, Ste60, Tol63]. Fluorescence [FJMS66]. Fluorite [Zer67].

Flux [AK64, SC64, SR64c]. Fluxoid [Tin64]. Fock [LPLH63, Löw60, MB64, Sla63b]. Forbidden [Wei61]. Force [BDRA60, Bur60, CH60, RHW68, Sch60]. Force-Free [Sch60]. Forces [Cha67, Dal63a, Hir63, KS63a, Phi67, SM65, SM66, TH60, VM60, Ing67, KS64]. Form [HMW63, Pri63, RÖ63, SD67]. Formal [Pet67]. Formaldehyde [GBD60]. Formalism [DK62, JS61, JS62]. Formalisms [And64]. Forms [Hub61]. Formula [Wil64]. Formulas [BvdH62, McK63]. Formulation [Coo67]. Formulations [Hav64]. Foundation [Bas65]. Foundations [CN61, CN64]. Four [BKR62, Hav64]. Four-Dimensional [Hav64]. Four-Layer [BKR62]. Fourier [PP63]. Frames [Rom63]. Free [AHJW60, CN67, FKP60, JM66, Mar62, Sch60, SDH64, Wyd64]. Free-Electron [CN67]. Free-Surface [AHJW60]. Frequencies [Bai68, Ost61]. Frequency [Mcc66, VK68]. Frontiers [Whe62]. Fully [Ost61, Pai60]. Function [BC62, Löw60, Nes63, Sla65]. Functional [Bru61]. Functioning [Blo62]. Functions [AK60, Gil63, Har63, KR60a, Men64c, Mus67, Ran60b, RÖ63, Ric60, RSW60, Rue64, Sch64b, SB67, Sil64, Sin63b, SLV62]. Fundamental [CD65, Red68, TPL69, TPL73]. Further [War68]. Future [Con68b, Whi67].

Ga [Weg64]. Galactic [Dic62]. Galaxies [BBS63]. Galvanomagnetic [Dre61b]. Gamma [Mar64a, MS64a, PB65, RB67b, RHW68, SV66]. Gamma-Ray [RHW68, SV66]. Gap [DP64, Gol64, RW64, Wad64]. Gas [Alf60, BvdH62, Bur60, CN67, KL60, Ost61, SD61, SM65, SM66, She60b]. Gas-Ionizing [KL60]. Gaseous [Ago60]. Gases [Cab60, EHD69, Pol64, Rue64, SG63, ST65]. Gauge [Sch64d]. Gaunt [BvdH62]. Gaussian [Har63]. General [Ara60, Har60, MEBL67, Mös64, Sci64a, Bas65, Ber61, DK62, Hav64, Pen65, Sci64b, Whe62]. Generalizations [She60b]. Generalized [Jos65, Men64c, Pra63, Pri63, RC67]. Generally [Fle60a, Fle60b]. Geomagnetic [Cha60]. Geometric [HVW65]. Geometrodynamics [Whe61]. Geometry [Whe62]. Geophysical [Jas60, Jor62]. GeV [Duc67]. Ginzburg [Mar64b]. Goldstone [Day67]. Gordan [dS65, MJC64, dS63]. Graphical [MEBL67]. Graphites [Kle62]. Gravitational [Ago60, And64, BW57, BW61, Fow64b, Fow64c]. Gravity [Tol63]. Green [FT60, KI60]. Grey [SK68a]. Ground [KR60b, KRS60, MWY60, MP63, RW60, SK63, SSW63]. Ground-State [SSW63]. Group [AH68, BI66a, BI66b, Bar62b, Bec64, JS61, Löw67, Mos62, JS62, Mic67]. Groups [Alt63, AC65, BDFL62a, BDFL62b, BD68, IN66, Mcv65, Sla65]. Gudder [BB68]. Gyromagnetic [HJ62, Sco62].

H [GO65, HS63, KR60a]. H. [FT60]. Hadrons [Chi69]. Hall [Bro60a]. Hamiltonian [CJS63a, CJS63b, Gar66, Ito65]. Hard [AK64, KHS64, SR64b]. Harmonics [AC65, AB65, FW63]. Hartree [LPLH63, Löw60, MB64, Sla63b]. Harwell [Per67, Whi67]. HCHO [FB60b].

Heat [CS64, Nak60, PLG64, SCGP64]. Heitler [Her62]. Held [Par60]. Helium [And66, AM66, AM67, BS67, JF62, Man67, Mey65, MP63]. Heliumlike [RW60]. Hemoglobins [Kot63]. Hexagonal [SK68b, AB65]. Hidden [Bel66, BB66a, BB66b, BB68, Gud68, JP68, Tut68]. High [Bar68, Bea64, Che61, Chi69, Ded62, Ded63, Des64, Dom69, Dre61a, Erb66, GO65, Hak64, Kun61, Lea67, Mcc66, MS64b, Pal68, Sel65, Ser64, Sle60, Van64, Wil59, Wil60, Wil64]. High- [Che61]. High-Energy [Bar68, Chi69, Ded62, Ded63, Dre61a, Erb66, GO65, Pal68, Ser64, Van64, Wil64]. High-Field [Bea64, Des64, Hak64, MS64b]. High-Frequency [Mcc66]. High-Temperature [Sle60, Wil59, Wil60]. Higher [FVV61]. History [Bru67]. Hollow [LS60]. Homonuclear [KMKI60]. Honor [Löw63a]. Hubbard [EH68, EL68]. Hydride [VS63]. Hydrocarbons [De 63, Lie60]. Hydrodynamic [Gri68, Gri69]. Hydrodynamics [AM66, AM67]. Hydrogen [BCF^{+}61, BI66a, BI66b, BS62, GY69, Ham62, KR60b, KW63, MWY60]. Hydrogenic [Men64c]. Hydromagnetic [CFH60, FR60, HLS60, HR60, Kap60, Res60, Wol60]. Hydroxides [GBG64]. Hylleraas [Löw63a]. Hyperfine [BLBM63, McD63, Mös64, OA63]. Hyperfragment [CSR^{+}62]. Hyperon [AFL61]. Hyperons [BCF^{+}61, SS61]. Hypersonic [Mey60]. Hypochromism [RC67]. Hypothesis [Jor62, O’K63]. Hysteresis [Bal64a, Liv64, SR64b].

Ice [O’K63]. Idealized [Cha60]. Ideas [Lap67]. II [AB65, AM66, Ano64b, BI66b, Ber64, BKR62, CP68, CCM64, CNO64, CJ67, DLLR64, Goo64, JF62, KS63b, Liv64, Ran60b, AM67]. III [Ano64c, CS64, HS63, Jos68, Lax66a, SK63]. Impact [Coo67, KD66, MS68, MS69, Rud68]. Impedance [MF64a, Wal64]. Implications [Dic62, Lor60, Löw63b, TPL69, TPL73]. Imply [SC64]. Importance [BLBM63]. Important [TB69]. Improved [DP63, Ric60]. Impure [DEM68, Jér68, Mik68, Wip64]. Impurities [BGR64, Coh64b]. Impurity [DS68]. Including [Tol63]. Increasing [Bro60b]. Indium [Mei64]. Indium-Tin [Mei64]. Individual [Lan60]. Inelastic [EGLM65, Hos67a, MT69, She63, SV66]. Inert [SG63]. Infinite [BM68]. Infinite-Dimensional [BM68]. Influence [Coh64b, MS64b, Udo65]. Information [Mcc65, Rio65]. Infrared [GL64, Ran62, VK68, Zer67]. Inhomogeneously [Sto69]. Initio [AK60]. Inner [Sew64]. Inorganic [Mas65]. INS [Hos67b]. Instability [Han68, LP67]. Insulator [Mot68]. Integral [Gar66]. Integrals [Bru61, JN63, Men64c]. Integration [Che63]. Intense [Erb66]. Intensities [Str64]. Interacting [CJS63a, CJS63b, Van64]. Interaction [AAA69, Ama67, BT65, BDFL62a, BDFL62b, BW57, Dre61a, FB60a, GMW67, Hel67, Kaz67a, Ker61, KP64b, KRS60, Lea67, LF67, McD63, MWY60, Mor67, Mos67, Nig63, Sel65, Sla67, Tam67, Won67, Wu64, BW61]. Interactions [Abr61, Ada61, ABB^{+}61, Che61, Che62, Cho61b, CD67, Coh64b, Dal61, Fin64, Fub61, GS67, Kar60, Lie60, Mcc66, Mös64, SPRM61, SK62, Tal62, Van62, Wet61, Cho61a]. Interatomic [Cha67]. Interface [Ger68]. Intermediate [Bre67, Che67, MF60]. Intermolecular [Dal63a, Hir63, SM65, SM66, VM60]. Internal [FKS61, NG68]. International [Ano64k, Tay61, BS64, Löw63a]. Internuclear [SLV62]. Interphase [Wip64]. Interpolative [McK63]. Interpretation [Gil63, KK65, Shi64]. Interstellar [Bir64, DG66]. Interstitial [Des64]. Introduction [Ano61, Ano67f, FS60, Kot63, Löw63a, RN65, Tho68a]. Introductory [Goo65, Par60]. Invariance [Cso65, CJS63a, CJS63b, FRW62, HVW65]. Invariant [Boy60, BY64]. Investigations [Don65, Kaz67a]. Inviscid [Lud60]. Invited [Fow64a]. Involving [Dal67b]. Ionic [Aiz62]. Ionization [KD66, Lor60, Rud68]. Ionized [Bur60, Cab60, Ost61, Pai60]. Ionizing [Gro65, KL60]. Ionospheric [Ves60]. Ions [KD66, KP64b, Van62]. Ir [MGC^{+}64]. Iron [Kot63]. Irradiation [Sei62]. Irreversible [KM60, ST65]. Ising [Bru67, SML64]. Island [Löw63a]. Isolated [Coo67]. Isomer [Shi64]. Isometric [Fri65]. Isothermal [Mac66]. Isotopic [SM66]. Isotropic [HLS60]. issue [Tay61]. IV [AH68, Ano64d, Lax66b].

J. [DPSY64]. January [Ano63, Löw63a]. Jauch [BB66b, BB68]. Jet [LS60]. Jets [Jun60]. Josephson [Fis64, SJH64]. Junction [Sca64]. June [Par60].

K-Nucleon [Ker61]. Kapitza [Pol69]. keV [Deh65, KKR65]. Kind [CS64, DM64, DLLR64, Gor64a, KLS64]. Kinematics [Ded62, Ded63]. Kinetic [Sch68]. Knight [CGJY64]. Knock [Rio65]. Knock-Out [Rio65]. Knowing [Pol62]. Knowledge [CD65]. Koopman [Lor60]. Kronig [Cal63].

L [Cal63, Cal63]. Ladder [CJ67, Jos67, Jos68]. Lagrangians [GG69]. Lambda [WSL^{+}61]. Laminar [Lyk60]. Landau [Cun64, Mar64b]. Langevin [Lax66b]. Large [Her62]. Laser [TB69]. Lattice [AC65, AB65, Han68, KT64, Mar64a, PLG64, Wal65]. Laue [SPB69]. Laws [Fle60a, Fle60b]. Layer [BKR62, Jas60]. Layers [Jun60, Man67, Nap60]. LCAO [Ran60b]. Lead [DVB64, GL64, LS64b, PLG64, SCGP64, Wad64]. Lecture [Jos64b]. Leiden [Gor64b]. Leipnik [FT60, KI60]. Lenz [Bru67]. Leptonic [Nie68]. Leptons [Dom69]. Levels [BB67, Cal63, CFMM63, GY69, Hij60, HM65, Lan60]. Li [JBJ^{+}65, NMC65, BB65, DSF65, Ric60, YJE^{+}65]. Lie [JS61, JS62]. Lie-Group [JS62]. Light [Ewa65, Gol63, Gri68, Hay63, MF60, Mou66]. Line [HV66, Sho67]. Linear [Ado63, CN61, CN64, Mul60]. Linearized [Ost60, RM60]. Lines [AK64, Coo67, DM64, Sch60, Sto69]. Linked [Bra67]. Linked-Cluster [Bra67]. Liquid [BB62, BS67, CFH60, Yan62]. Liquids [PP63]. List [Ano64j, Tho68a]. Local [Fle60b, FKC60, Hor68, Wer64, Fle60a]. Local-Energy [FKC60]. Localizability [BM68, Wig62]. Localization [DEM68]. Localized [Ara68, DS68, ER63]. Location [CFMM63]. London [Goo64, Her62]. Long [Cha67, Lit64, TH60, THH60, Yan62]. Long-Range [Cha67, Lit64, TH60, Yan62]. Low [Bai68, BNM^{+}64, BB65, BS62, Car65, Car61, Che61, Hel67, Hoh62, KLJ67, Kes69, Köp67, Lud60, Man67, Phe68, Sim67, SPB69, Vog62]. Low-Energy [BS62, Che61, Kes69, Köp67, Phe68, Sim67, SPB69]. Low-Temperature [BNM^{+}64, Hoh62]. Lower [Wei63]. Luminescence [Hoc62].

M [Cal63]. Mach [Sch64a]. Macroscopic [AM66, AM67, TPL69, TPL73]. Magnetic [BD68, Bro60b, Bur60, CP68, CS64, DS68, DH60, DLLR64, DEM68, Erb66, Fer60, Fis64, GBG64, GP66, HLS60, Ham62, HK64, HMW63, HJ62, Hei64, Hoh62, JB64, KM60, KLS64, Kle68, KP64b, Kol60, Kun61, Liv64, Lud60, Mar64b, Nak60, NT64, NK64, Pai60, RM60, SW60, Tay60, Van62, Ves60, Wan62, Wer64]. Magnetism [Ell64, MCJR64]. Magnetization [Bea64, Hit64, MKQ64, KLS64]. Magnetized [Alf60, Leh60]. Magneto [Ger60, Gra60, Jun60, Lüs60, Lyk60, Nap60, She60b, Spi60, Tat60, Tay61, Zhi60]. Magneto-Fluid [Ger60, Gra60, Jun60, Lüs60, Lyk60, Nap60, Spi60, Tat60, Tay61, Zhi60]. Magneto-Gas [She60b]. Magnetoaerodynamics [RM60]. Magnetodynamic [Ago60]. Magnetohydrodynamic [FKP60, Has60, KL60, Mey60]. Magnetohydrodynamic-Hypersonic [Mey60]. Magnetohydrodynamics [AHJW60]. Main [Str64]. Main-Sequence [Str64]. Manifolds [Fri65]. Many [Bra67, Frö60, Löw63c, Nes61, SML64, Sin63b]. Many-Body [Bra67]. Many-Electron [Frö60, Sin63b]. Many-Fermion [Nes61]. Many-Particle [Löw63c]. Markoff [Lax66a]. Masers [Web59, Web60]. Mass [Ago60, EL68, GGJ^{+}69, RHW68, SV66]. Mass-14 [RHW68]. Massive [Bir64, Fow64c, Fow64b]. Mathematical [BC60]. Matrices [And63, CLS63, Col63]. Matrix [Che62, CD67, Coo67, DW64, GL60, Mac67, McM61, McW60, RB67b, Sta62]. Matrix-Spinor [GL60]. Matter [Bir64, Day67, RB67a]. Max [Bor64]. Maxwellian [BvdH62]. Mean [Wyd64]. Measurement [BB66a]. Measurements [Bai68, BBD^{+}67, Ber60a, DP64, Duc67, Eis64, Hos67b, HW67, Rom63, SCGP64, SMR64, HW68]. Measuring [Bur63a, Bur63b]. Mechanic [Lyk60]. Mechanical [Bro60a, Hub61, SW60, VM60, Wig62]. Mechanics [Bel66, BB66a, BB66b, CN68, FT60, GL60, Gud68, Hav64, Hyl63, Ing61, JS61, JS62, Kar60, Löw62b, Löw63a, Löw67, Par60, Str66, JS62]. Mechanism [KKR65]. Mechanisms [Adl68, Zup65]. Mechanized [Bar63]. Medium [DG66, Kap60, Zhi60]. Melting [Bab63]. Mercury [AHJW60, HF68, LS60, PLG64]. Meson [BAC61, GS67, Köp67, Won67]. Meson-Nucleon [Köp67]. Mesons [SS61]. Message [Bor64]. Metagalaxy [Alf65]. Metal [Adl68, AH68, BGR64, BHMM64, CGJY64, Han68, HF68, Jér68, LA68, Mik68, Mot68, SK68b, SMR64, Tho68b]. Metal-Ammonia [Tho68b]. Metal-Insulator [Mot68]. Metal-Nonmental [Adl68]. Metal-Nonmetal [HF68, Jér68, Mik68, Tho68b]. Metal-Semiconductor [Han68]. Metal-to-Semiconductor [SK68b]. Metallic [AS60, Ara68, Hyl68, LA68]. Metals [BGR64, Col64, DS68, Geb64, Sei62, TW63, Van62, Vas68]. Meteorite [And62]. Method [Car65, Cas64, De 63, DW64, FKC60, GL60, Her62, KI60, LS64a, MEBL67, Pre63, Row68, Sin63b, Sla63b]. Methods [Ara60, Gar66, Lax66b, Nes61, SW60, Sch68]. MeV [RCF^{+}61, BKF65, EGLM65, Hos67b, Kaz67a, Moy61, PB65, Tho67]. MeV-Preliminary [Hos67b]. Microwave [MJ68, SR64a]. Microwaves [SJH64]. Mills [Fin64]. Mixed [DLLR64, SPB69]. Mixtures [SD61, SG63, SM65, SM66]. MO [Ran60b, JB64]. Mobility [Man67]. Mod [Tay61]. Model [Blo62, Bru67, EL68, GY69, Hit69, Hor68, Mik68, Phi64, Row68, SML64, dS63, dS65]. Models [Ing62]. Moderately [Cha67, EHD69]. Modern [Con68b]. Modes [DM64, Har67]. Molecular [AS60, Bar63, Ber60a, Boy60, Cou60, ER63, För63, Fre60, Hoc62, Hur63, I’H63, JN63, Kar60, KP66, Löw62a, Löw63a, Mul60, Par60, Ran60a, Ran60b, Ric60]. Molecule [Boy60, Ham62, KR60a, KR60b, KW63, MWY60, Nes60, Ran60a]. Molecules [Ara60, BDRA60, Ham62, Har63, HL68, JN63, KD66, KW63, KMKI60, MB64, Mus67, Phe68, Pre63, Ran60b, Sch63, Tel62, VS63]. Molybdenum [Mat61]. Moment [LS64a]. Momenta [Kra65]. Moments [BLBM63, KP66]. Momentum [Löw64, MEBL67]. Monoxide [Hur60]. Mössbauer [Ano64k, CNO64, Ell64, Kie64, PP64, SF64]. Motion [Ago60, AK64, Bro60b, Bur60, Ger68, Kle68, Löw62b, Pla61, Row68, Ste60, Whe61]. Motions [AM66, AM67, LCW68, Pai60]. Mott [EH68]. Moving [Cho61a, Cho61b]. Multi [Adk64]. Multi-Particle [Adk64]. Multicomponent [SG63]. Multiple [Sch61, Sco63]. Multiple-Scattering [Sch61].

N [TB69, Bre67, GMW67, GS67, KK65, Ric60]. Narrow [CP68, Hor68, KC68]. Natural [HNLD63, RÖ63, HS63]. Nature [HS63, Hon68, Hyl68, Rue62, VS63]. Nb [CCM64, CC64, CS64, Gol64, Hit64]. NbMo [HGB64]. Near [KGH^{+}67, FVV61, LCW68, Tho67]. Neutral [Sch60]. Neutrinos [BW57, BW61]. Neutron [Cin66, CFMM63, HW67, HW68, MG65, Sim67, Sla67]. Neutron-Carbon [HW67, HW68]. Neutron-Neutron [Sla67]. Neutron-Proton [HW67, HW68, Sim67]. Newtonian [Hav64, Mey60]. Ni [PB65]. Niobium [DP64, LC64, RAG64]. NiS [SK68b]. Nitrene [SSW63]. NMR [Jér68]. Noble [LA68]. Noise [Lax66a, Lax66b]. Non [CMS67, Sch64d]. Non-Abelian [Sch64d]. Non-Dynamical [CMS67]. Nonadiabatic [KW63]. Nonequilibrium [Lax60]. Nonforward [Chi69]. Nonhomogeneous [Sil64]. Nonideal [EL64]. Nonionized [Alf60]. Nonisotropic [HLS60]. Nonlinear [FW63, KB67a, KB67b, Lax66a, Mit64, Pet67, War65]. Nonlinearities [SR64a]. Nonmental [Adl68]. Nonmetal [HF68, Jér68, Mik68, Tho68b]. Nonrelativistic [BvdH62, Ito65]. Nontransition [BGR64]. Normal [GBD60, Har67, Löw62b, MV68, MF64a, War68, Wyd64]. Note [Goo65, Mel62, Par60, Van62]. Nuclear [Bar62c, BNM^{+}64, BLBM63, Bra67, Day67, DP69, Fes64, Fow64a, GGJ^{+}69, GS67, GY69, HK64, Ing62, JM66, Kra65, Lan60, Moo60, Mös64, Mos67, MOK64, NK64, Per66, RB67a, Ric60, Rio65, RHW68, SW60, Sch65, Tam65, Udo65, Vog62, Wil64, Ing67, BLBM63]. Nuclear-Mass [GGJ^{+}69]. Nucleation [Aha62]. Nuclei [BBS63, Dav65, Eic64, GP66, Gol63, Hay63, KS63a, KS64, MF60, Mar64a, MPS63, MS64a, SV66, She60a, Tal62, Tam65, Van62]. Nucleon [AAY65, Ama67, Arn67, Bre62, Bre63, Che67, Don65, Duc65, Fub61, HW63, HMW63, Kaz67a, KLJ67, Kaz67b, Ker61, Köp67, LP67, Lea67, LF67, Lyt65, McK63, Mic67, Mor67, Mos67, Nig63, She63, SV66, SD67, Tam67, Tho67, Won67]. Nucleon-Nucleon [Ama67, Arn67, Bre62, Bre63, Che67, Kaz67a, KLJ67, LP67, Lea67, LF67, Mic67, Mos67, SD67, Tam67, Tho67, Won67]. Nucleons [Hel67]. Nucleosynthesis [Bel67]. Nucleus [BT64, Mol64]. Number [Lud60]. Numbers [RM60].

O [Hon68]. Observables [Ber61]. Observation [KB67a, KB67b]. Observational [Sch64a]. Observed [AD65, Str64]. Obtainable [Mcc65]. Occurrence [BBS63]. Octet [dS63, dS65]. October [Tay61]. Odd [SV66]. Odd-Mass [SV66]. Off [Yan62]. Off-Diagonal [Yan62]. One [Boy60, DK62, She60b]. One-Dimensional [She60b]. Onset [Hor68]. Open [Roo60]. Operational [Bas65]. Operator [EL68, Hub61, Pri63]. Operators [Ado63, BM68, CJ67, FG61, Jos67, Jos68, Löw63c, Löw64, Mar62, Sin63a]. Oppenheimer [DPSY64]. Optical [CE63, FW63, MW65, MHB67, SF64, VBB68, War65, Was60, Zer67]. Optics [Ewa65]. Orbit [GK63b, GK63a]. Orbital [CGJY64, Ric60, RF60]. Orbitals [Boy60, De 63, ER63, HS63, KR60b, Mos63, Mul60, RW60]. Order [SK63, Yan62]. Organic [Hoc62]. Orsay [Mic67]. Orthogonality [Che63, CD63]. Oscillating [HR60]. Oscillations [Ost60]. Oscillator [FC68, FC69]. other [Rio65]. Our [CD65]. Overview [GMW67]. Oxford [Men64b]. Oxides [Adl68]. Ozone [Ber60a].

P [KK65, AD65, BT65, Mcc65]. Packed [AB65]. Pair [ACG63, Con68a, KC68, MOK69, Pal68, Sil64]. Papers [Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano64h, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e, Fow64a, Par60]. Parahydrogen [VK68]. Paramagnetic [RW64]. Paramagnetism [CGJY64, HL68]. Parameter [Bre67]. Parameters [Bab63, Cal63, Cat67, HW63, LCW68]. Partial [FVV61, Mac62]. Participants [Ano63, Ano64j, Tho68a]. Particle [Adk64, AD65, Bar68, Bro60b, CFMM63, DK62, Jas60, Kle68, Löw63c, MEHW65b, Mor67, Ne’65, Pai66, Pra63, Zup65]. Particle-Particle [MEHW65b]. Particles [BY64, Che62, CJS63b, Dal61, Ded62, EP61, Fro65, Ing61, KRS60, Mel60, Mel62, NG68, Phi65, Roo63, RBGB^{+}64, RBGB^{+}65, RBGP^{+}67, RBSBG^{+}68, Sco63, Tam65, Van64, CJS63a, Ded63]. Particular [Ham62]. Passage [DH60]. Past [Con68b, Ima60, Lud60, Sea60]. Path [Gar66]. Path-Integral [Gar66]. Paths [Wyd64]. Paul [Ano62]. Pauli [Kra65]. Peak [Des64]. Peratization [KP64a]. Perceptron [BKR62, Blo62]. Perfect [BC64]. Performance [SR62]. Periodic [Sch60]. Peripheral [Dre61a, Jac65b]. Perpendicular [HLS60]. Persistent [DK62, FKS61, SR64c]. Perturbation [EL68, KS63b, Mus67, Pri63, SK63, Sin63a, War65]. Phase [Arn67, Bre67, BC62, BY64, CN68, Fer60, HW63, Hos67b, Hyl68, KLJ67, Kaz67b, McK63, NT64, Per67, RB67b, TPL69, TPL73, Wal67, RCF^{+}61]. Phase-Defined [RB67b]. Phase-Parameter [Bre67]. Phase-Shift [Arn67, KLJ67, Kaz67b, Per67, RCF^{+}61]. Phase-Space [BY64]. Phases [Yan62]. Phenomena [BC60, Dre61b, FW63, HJ62, KGH^{+}67, KY68, Ves60]. Phenomenological [Bre67, Hos67a, Lea66]. Phenomenology [Chi69]. Philosophy [Pol62]. Phonon [KTB64]. Phonon-Assisted [KTB64]. Phosphors [Mas65]. Photodisintegration [Hay63]. Photoemission [PBS68]. Photons [MOK69]. Photoproduction [CMS67]. Phys [Tay61]. Physical [Bun67, BY64, Rue62, Sci64a, Sci64b, TPL69, TPL73]. Physics [Bie60, Bru61, CD65, FRW62, GS67, Hue61, Mcv65, Mor67, Pai66, Tis63, Con68b]. pi [Wal67]. Pinch [Fri60]. Pion [AFL61, Ama67, ABB^{+}61, Arn67, Fub61, HW63, Kaz67b, LP67, Lyt65, McK63, SPRM61, SD67, SK62]. Pion-Hyperon [AFL61]. Pion-Nucleon [Fub61, HW63, Lyt65, McK63, SD67]. Pion-Pion [ABB^{+}61, SPRM61, SK62]. Pion-Production [Ama67, Arn67, Kaz67b]. Pions [BCF^{+}61]. Piron [BB66b, BB68]. Planar [GC64]. Plane [HR60, Pen65]. Planetary [Dal67b]. Plasma [Alf60, Bal60, Bie60, Cha60, FKP60, HLS60, HL60, Hoh62, Hue61, Kov60, Leh60, LI66, Mcc66, Ost60, Pai60, Sle60, TH60, Wil59, Wil60]. Plasmas [KM60, Kol60]. Point [Alt63, KT64]. Points [KGH^{+}67]. Poisson [Jos64b]. Polar [Ves60]. Polarization [Aiz62, Bar62c, BNM^{+}64, BLBM63, Cso65, Duc67, FG61, FKS61, Kes69, McM61, RCF^{+}61, Sim67]. Polarized [Dom69, Lap67]. Pole [Hit69, Lea66]. Poles [Bar68]. POLYATOM [Bar63]. Polyatomic [Har63]. Polyenes [THH60]. Polymers [SW60]. Polynomial [Che63]. Polytropes [Fow64b, Fow64c]. Position [BM68]. Positive [BCF^{+}61]. Positrons [BS62]. Possibilities [KB67b, KB67a]. Possible [Cso65, HR68, LS64a]. Potential [GP64, GY69, Lor60, MHB67, SLV62, VS63]. Potentials [KP64a]. Precipitation [TW63]. Precision [Ran62]. Predictions [BC60]. Preliminary [Hos67b]. Presence [Bur60, Nak60]. Present [Cou60, DP69, Gom63, Lea66, Whi67]. Pressure [Bab63, HLS60, MS64b, OBL^{+}64, Vas68]. Prewar [Men64b]. Prigogine [ST65]. Principle [Kra65, Sch64a]. Principles [Cso65, HVW65, Mel60, Mel62]. Probability [Per66]. Problem [Bel66, BT64, BB66a, Bra67, CD63, DP69, FT60, GL60, KI60, RB67a, RN65, Sch61, SML64, Whe61]. Problems [Ada65, BC60, Cas64, Cha60, Dal67b, Gra60, Jac65a, KMKI60, Kot63, Pip64b, PK63, Pol62, Spi60, Whe62]. Procedure [FB60a]. Proceedings [Löw63a, Tay61]. Process [Bur63a, Bur63b, NG68, WEHM65]. Processes [CMS67, Dal67a, Dal67b, Dre61a, Erb66, Ger61, Gly67, KA65, KM60, Lax66a, ST65, TB69]. Production [Ama67, Arn67, BCF^{+}61, BAC61, EFG^{+}61, Jac65b, Kaz67b, MOK69, Pal68, SPRM61, WSL^{+}61]. Products [Alt63]. Professor [Löw63a]. Profiles [Sho67]. Program [Ran60a]. Programs [Nes63]. Progress [EFG^{+}61, Fub61]. Project [Nes60]. Projection [Löw62b, Löw64]. Proof [BB66b]. Propagation [Kap60]. Properties [And63, BSBGP^{+}69, BHMM64, CS64, Cun64, DLLR64, Eic64, GBG64, Hak64, Ham62, HNLD63, Jér68, KLS64, Kle62, LI66, MW65, MV68, Mat63, MF64b, Per66, VBB68, War68, Zer67]. Property [Pen65, Rue64]. Propoerties [Por64]. Proposed [BB66a]. Proton [BBD^{+}67, BAC61, Car65, EGLM65, Hos67a, HW67, HW68, Köp67, Löw63b, Lyn61, O’K63, PB65, Ser64, Sim67]. Proton-Deuteron [BAC61, EGLM65]. Proton-Proton [BBD^{+}67, Hos67a, Köp67, Ser64]. Proton-Tunneling [O’K63]. Proximity [SCGP64]. Pseudo [KA65, Ros65]. Pseudo-Euclidean [Ros65]. Pseudo-Resonances [KA65]. Pseudoreversible [Hit64]. Pulsation [Chr64]. Pumps [Ros60]. Pure [Wip64]. Pyroelectricity [Aiz62]. Pyrolytic [Kle62].

Quadratures [Che63]. Quadrupole [KP66, Ric60]. Quantization [AM66, BC62, NG68, SC64, SR64c, Tin64, AM67]. Quantized [PM64]. Quantum [BM68, Bel66, BB66a, BB66b, BC60, Bur63a, Bur63b, CN68, Dew63, Dir62, Dre61b, FB60b, FB60c, Gud68, Hub61, Hyl63, Ing61, Ito65, JS61, JS62, KB67a, KB67b, Kar60, Kot63, Löw62b, Löw63a, Löw63c, Löw67, Mar62, Nes61, Par60, Str66, TPL69, TPL73, VM60, Wig62, Yan62]. Quantum-Mechanical [Hub61]. Quasi [JM66, Mey60, YH65]. Quasi-Elastic [YH65]. Quasi-Free [JM66]. Quasi-Newtonian [Mey60]. Quasifree [Jac65a]. Quasiparticle [Wad64]. Quaternions [Ras64].

R. [FT60]. Radial [Men64c]. Radiation [Bol64, Fow64c, Fow64b]. Radiative [Max69, MT69, Pla61]. Radio [Ost61]. Raman [VK68]. Range [Bal64b, Cha67, FB60c, Lit64, TH60, VM60, Yan62]. Rapid [Gil63]. Rapidly [Cho61a, Cho61b]. Rare [CNO64, Ell64, Kie64, Mös64, Pol64, SMR64]. Rare-Earth [CNO64, Kie64]. Rate [Eyr62, Gly67]. Rates [Cal63, Dic62]. Ratio [Sco62, Wad64]. Ratios [SPRM61]. Ray [Azá63, BB67, Bea67, Bol64, Dan68, PP63, RHW68, SV66]. Rays [BC64, Ewa65, Mar64a, MS64a, RB67b]. Re [JB64]. Reaction [Ada61, AWY65, BB65, BKF65, Deh65, DSF65, EWM^{+}65, Eyr62, JBJ^{+}65, KKR65, KK65, Löw63c, MEHW65b, MEHW65a, Moo60, NMC65, PB65, Pla61, Sin63a, Swa65, WEHM65, YJE^{+}65]. Reactions [AD65, BHK^{+}65, Cin66, CFMM63, Fes64, Kra65, MF60, Mcc65, Mey65, Rio65, Sch65, SBJP65, Vog62, Zup65]. Reactivity [Pit63]. Reactors [DW62]. Real [RÖ63]. Realistic [Ing67]. Reconsidered [Gud68]. Reduced [Lan60, RB67b, Sch64b]. Reducible [Gra60]. Reduction [KT64].

Reevaluation [BB67]. Reference [Rom63]. Reflection [AS60]. Reflections [SPB69]. Refractivity [Sho67]. Refutation [BB66b]. Regge [Bar68, Hit69, Lea66]. Region [Gri68, Hos67a, Kaz67b, Moy61, Gri69]. Regions [BY64]. Related [HJ62]. Relating [Bab63, Cha60]. Relation [Ell64, Hav64, JS61, JS62]. Relations [Bie60, GGJ^{+}69, HW63, Hue61, Lyt65, Sch65]. Relationships [DS68, HL60]. Relative [Ger68, LCW68]. Relativistic [CJS63b, Dir62, Fow64b, Fow64c, Frö60, Mac62, Ros65, CJS63a]. Relativity [Bas65, Ber61, Hav64, Pen65, Ras64, Sci64a, Sci64b, Whe62]. Relaxation [HK64, Hub61, MB63, NK64, O’K63, Sch68, TB69]. Remarkable [Pen65]. Remarks [Ber60b, Gor64a, Hyl68, Res60, Sea60]. Reminiscences [Hyl63]. Reply [BB68, Wil60]. Report [EFG^{+}61]. Representation [Arn67, Mar62, McM61]. Representations [Bar62b, IN66]. Reproduction [Mor64]. Rescattering [KA65]. Research [Mor67]. Residual [DVB64, KS63a, KS64]. Resistance [Pol69]. Resistive [JB64, KHS64]. Resistivity [DVB64, RAG64]. Resonance [Deh65, FVV61, HK64, KKR65, Mar64a, McD63, NK64, OA63, SPRM61, SW60, Sto69]. Resonances [Ada65, AFL61, Jac65b, KA65, MHB67, Sal61]. Resonant [MS64a, Roo63, RBGB^{+}64, RBGB^{+}65, RBGP^{+}67, RBSBG^{+}68]. Response [Pet67]. Responsible [Coh64b]. Restriction [Fre60]. Resultant [GGJ^{+}69]. Results [Ber68, BDRA60, Hos67b, Mic67, Nak60, RF60]. Retardation [MB63]. Rev [Tay61]. Revealed [Jac63]. Review [BSBGP^{+}69, Coh68, Mac69, Sco62, Sle60, Tho67, Wal67, Wil59, Wil60]. Reviews [Con68b]. Revisited [JP68]. Reynolds [Lud60, RM60]. Rh [MGC^{+}64]. RIAS [Nes60]. Riemann [Lan62]. Riemannian [Fri65, Ros65]. Robert [DPSY64]. Role [FM60, Fre60, HV66]. Rotating [AM66, AM67, Sle60, Wil59, Wil60]. Rotating-Plasma [Sle60, Wil59, Wil60]. Rotation [Bar62b, CE63, Tol63]. Rotational [Ago60, Phe68, VK68]. Rotations [Dav65]. Rotatory [Mos60]. Row [Ran60b]. Rules [Dal63b].

S [FT60, SK63]. Sagnac [Pos67]. Salts [Mös64]. Sanibel [Löw63a, Ano63]. Satellite [Jac63]. Sb [LCW68]. Scattered [Mou66]. Scattering [Arn67, Bar68, BBD^{+}67, Bre62, Bre63, BS62, Cat67, Che67, Chi69, Cso65, Dom69, EP61, EGLM65, Ger61, GP66, Gri68, Gri69, GO65, Hos67a, Jac65a, JM66, KLJ67, Kaz67b, Kes69, Köp67, LP67, Mac67, McK63, Mic67, MT69, MS64a, MOK64, Moy61, Sch61, Sco63, Ser64, She63, Sho67, SD67, Tam65, Tho67, YH65]. SCF [Ran60b]. Schemes [Löw60, Tal62]. Science [BS64]. Scope [Ran60a]. Screened [Mcc66]. Screening [Cal63]. Search [Aha62]. Second [BC62, CS64, DM64, DLLR64, Gor64a, KLS64, NG68]. Second-Kind [DLLR64, KLS64]. Second-Quantization [NG68]. Section [KD66, Moo60, RCF^{+}61]. Sections [Bel67, FVV61, HW67, HW68, Mol64, Wil64]. Seen [Gor64b, Men64b]. Selected [Ran60b]. Self [CJ67, Jos67, Jos68, KC68, Mor64, Mul60, RSW60, Roo60]. Self-Adjoint [CJ67, Jos67, Jos68]. Self-Consistent [KC68, Mul60, RSW60, Roo60]. Self-Reproduction [Mor64]. Semiclassical [CN67, Hub61]. Semiconductor [AH68, HR68, Han68, LCW68, SK68b]. Semiconductor-to-Metal [AH68]. Semiconductors [AH68, Coh64a, DEM68, Jér68, Mik68, MF64b]. Semidirect [Alt63]. Semiempirical [Lie60]. Semimetal [HR68, LCW68]. Semimetal-Semiconductor [HR68, LCW68]. Seminar [RN65]. Sensitivity [Cal63]. Sequence [Str64]. Sequential [JBJ^{+}65, KA65, Phi65, SBJP65, Swa65]. Series [BKR62]. Series-Coupled [BKR62]. Session [Ano64a, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano64h, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e]. Set [GGJ^{+}69]. Several [Don65]. Shape [HV66]. Shapes [Dan68, Sto69]. Shared [Abr61]. Shell [HL68, Row68]. Shells [Roo60]. Shift [Arn67, CGJY64, KLJ67, Kaz67b, Per67, RCF^{+}61]. Shifts [HW63, Hos67b, McK63, Shi64, Wal67]. Shock [Ber60b, Cab60, DH60, FKP60, Ger60, Gro65, KL60, Lüs60]. Shock-Tube [Ber60b]. Shock-Wave [Ger60]. Short [Bal64b, Kot63, VM60]. Short-Range [Bal64b, VM60]. Showers [Bol64]. Si [Weg64]. Sigma [EFG^{+}61, WSL^{+}61]. Sigma-K [EFG^{+}61]. Similar [Mas65]. Simon [Bab63]. Simple [BDFL62b, DSF65, KS63a, KS64, Lie60, Mac66, Mou66, Zup65, BDFL62a]. Since [Red68]. Single [CFMM63, Pra63]. Single-Particle [CFMM63, Pra63]. Singular [KP64a]. Singularities [Sca64]. Sixth [SK63]. Sixth-Order [SK63]. Sixtieth [Ano62, DPSY64]. Slabs [Mar64b]. Slater [Mul60]. Slater-Type [Mul60]. Slowly [Bro60b]. Small [BDRA60, Pai60, Pip64a, Sco63, Tam67]. Small-Angle [Sco63]. Sn [CCM64, CC64, Gol64, MKQ64]. Sodium [CFH60, MJ68]. Sodium-Ammonia [MJ68]. Solid [BB62, CFH60, Des64, JHC64, Löw62a, Man67, Mcc66, Pol64, VK68]. Solid-State [Löw62a, Mcc66]. Solids [Car61, Gly67, Pip64a, SF64, Sto69]. Soluble [SML64]. Solutes [SMR64]. Solution [BB66a, GL60, JHC64, KI60, Mar64b]. Solutions [Des64, MJ68, RM60, Tho68b]. Some [Ada61, AHJW60, Ber60b, Boy60, BDRA60, Cun64, Dal67b, DS68, Eic64, Gor64a, HJ62, Hyl68, JB64, KK65, KS63b, KMKI60, Lie60, Lor60, Mac66, Mac69, McW60, MS64b, Nak60, Per66, Pol62, PP60, RM60, Res60, Sal61, Sea60, Sel65, She60b, Zhi60, Mos62]. Space [BC62, BY64, Fro65, Ne’65, Sla65]. Space-Time [Ne’65]. Spaces [Fri65, Ros65]. Spatial [MP63, Sil64]. Special [Hav64]. Specific [SCGP64, CS64]. Spectra [Dan68, För63, Sla63a, SSW63, Str64, Tat60, VK68]. Spectral [FC68, Mou66, FC69]. Spectroscopic [EWM^{+}65, GBG64]. Spectroscopy [Ran62]. Spectrum [JF62]. Speculations [Sal61, SK68a]. Spherical [KS63a, KS64]. Spin [Bal64b, Cat67, GK63b, Her62, Kes69, Löw62a, McD63, NG68, OA63, GK63a].

Spin-Correlation [Cat67]. Spin-Orbit [GK63b, GK63a]. Spinor [GL60]. Spins [De 63, Mos63, Van62]. Splitting [Lan62, OA63]. Stability [FR60, Ger68, LS60, Sch60, Sch63, Tay60, Tel62]. Stars [Bel67, Fow64b, Fow64c, Str64]. State [Ada61, Ada65, BT65, Coh64a, Cou60, DLLR64, Gom63, GBD60, Hyl68, KR60b, KRS60, Kot63, Lax60, Löw62a, Mac66, Mac69, MWY60, Mcc66, Mey65, MP63, Pol64, RW60, SK63, Sel65, SDH64, SSW63, Wyd64]. States [AD65, Bre64, CH60, CD63, DS68, DK62, Fow64a, KHS64, KS63b, Lyn61, Por64, Roo63, RBGB^{+}64, RBGB^{+}65, RBGP^{+}67, RBSBG^{+}68, She60a]. Static [KGH^{+}67]. Stationary [Cab60, FR60]. Statistical [Bru61, FT60, Gol60, Gom63, GL60, McW63, Por64, Tho63]. Status [DP69, Lea66]. Steady [Gra60, Lax60]. Stellar [Bir64, Chr64, Dic62, MG65]. Stimulation [Mas65]. Stochastic [Ado63]. Storm [Fer60]. Storms [Cha60]. Strange [Dal61]. Stratified [Tol63]. Strength [Sch64b]. Strengths [FC68, FC69]. Stressed [Wal65]. Stripping [CFMM63, HM65, MF60]. Strong [Abr61, BDFL62a, BDFL62b, Che61, Che62, Dal61, Gro65, KM60, Wet61]. Strongly [Van64]. Structure [Azá63, Bec64, Ber60a, BLBM63, CLS63, Col63, Col64, Con68a, Cou60, CMS67, Dal63b, Ger60, Gol63, GBD60, Gre63, GO65, Hur60, JM66, MF60, Mos67, Ran60a, Ran60b, Sci64a, Sci64b, Sla63b, Smi64, SPB69, Tis63, Weg64]. Strucure [Rio65]. Studies [AD65, AHJW60, CNO64, Dal67a, EWM^{+}65, Kie64, Ran60a, Ran60b, Sel65]. Study [BHK^{+}65, BB65, EFG^{+}61, KS63b, LA68, Nie68, SK63, SB67, SLV62]. Substituent [SSW63]. Substitutional [Des64]. Substrates [MF64a]. Substructure [Smi64]. Sudden [Fer60]. Sulfide [Mas65]. Sulfides [Adl68]. Sum [Dal63b]. Summary [Coh68, GMW67, Ran60a]. Superconducting [Coh64a, Dou64, DVB64, GL64, Gol64, HK64, HT64, Hei64, LS64b, Mat61, MF64a, Mit64, NT64, NK64, Par64, PLG64, Sca64, SDH64, SJH64, SR64a, Tin64, Tox64, Wyd64]. Superconductivity [AK64, Ber64, BHMM64, Coh64b, Col64, Geb64, Gor64a, Gor64b, HGB64, Kun61, Lit64, LC64, MGC63b, MGC^{+}64, Men64a, Men64b, Mer64, MS64b, OBL^{+}64, Pip64b, SC64, Sew64, BS64, MGC63a]. Superconductor [Hak64, PM64, Wer64]. Superconductors [Adk64, Bal64b, Bar62a, Bea64, CGJY64, CS64, De 64, DM64, Des64, DLLR64, EL64, Goo64, Jos64a, KHS64, KLS64, KTB64, KT64, Liv64, MS64b, RW64, Row64, SR64b, Sil64, SMR64, TPL69, TPL73, Wal64, Yan62]. Superfluid [And66]. Superposition [MB63]. Superstructure [Smi64]. Surface [AHJW60, MF64a, Vas68, Wal64, Wip64]. Surfaces [Ing62, Pre63, SPB69]. Survey [AAA69, Cin66, Eic64, Mic67, Per66, TB69, Zup65]. Susceptibilities [War65]. Susceptibility [Bal64b, RC67, SMR64]. Sutherland [GP64]. Symmetric [BT64]. Symmetries [BDFL62b, Ne’65, BDFL62a]. Symmetry [GG69, MV68, Mar62, Mat63, Mcv65, Mel60, Mel62, NG68, Pai66, Sla65, War68]. Symmteries [Abr61]. Symposium [Ano63, Löw63a, Tay61]. Synchrocyclotron [Whi67]. System [AAY65, Bar63, HGB64, LCW68, Lyn61, Lyt65, TB69]. Systematic [AAA69, Per66]. Systematics [MPS63]. Systems [BM68, Dew63, Don65, Duc65, Frö60, KRS60, Löw63c, MGC^{+}64, Nes61, RÖ63, RW60, Roo60, Rue63, Wig62].

Table [GGJ^{+}69]. Tables [MJC64, Roo63]. Tacit [Pol62]. Tails [Hue61]. Tangential [Mar64b]. Tantalum [HT64, Mei64, Wip64]. Target [Lap67]. Targets [Kes69]. Technetium [Mat61]. Technique [Löw62b]. Technology [DW62]. Temperature [Bab63, BNM^{+}64, BGR64, Fis64, Hoh62, Mei64, RAG64, Sle60, Tox64, Wad64, Wil59, Wil60]. Temperatures [Car61, MG65]. Tensor [Lan62, RHW68]. Terms [RB67b, Tam65]. Terrestrial [Cho61a, Cho61b]. Test [Car65, Cso65]. Their [BD68, Col64, Hav64, Mul60, VS63]. Theorem [Lor60]. Theorems [Löw60]. Theoretic [GS67, LA68]. Theoretical [Aha62, Dal67a, KK65, SSW63]. Theoretically [Str64]. Theories [CJS63a, CJS63b, DK62, Fle60b, Fle60a, RC67]. Theory [AAY65, And64, AK64, Azá63, BI66a, BI66b, Bas65, BB66a, Bre62, Bur63b, Car61, Che62, CD67, Con68a, Day67, DP63, Dew63, Dir62, Dre61b, DW62, DW64, EH68, EHD69, EL68, Eyr62, Fer60, Fes64, FM60, Fub61, GMGL64, Gol60, Gom63, Ham62, Hav64, HJ62, Hub61, Jaf69, KGH^{+}67, Kap60, KW63, KMKI60, Lea66, Lie60, LF67, Löw62a, Löw63c, Mar62, McW60, McW63, MB64, Moo60, MS64a, Mus67, Nes61, Ost60, Pet67, Pri63, RC67, Roo60, Rud68, SR62, Sch64c, Sco63, Sew64, Sho67, Sin63a, Sta62, ST65, Swa65, Tel62, Tho63, Tol63, VK68, Vog62, War65, Was60, Wer64, Wol60, Won67, Zer67, Bre63, Bur63a, Tut68]. There [Gor64b]. Thermal [Car61, CC64, CS64, DLLR64, HL60, KLS64, KT64, SD61, SR62, SG63, SM65, SM66, DLLR64, Mcc66]. Thermal-Diffusion [SR62]. Thermodynamic [Hak64, LI66, Mor64]. Thermodynamics [Red68]. Theta [Ing61]. Thin [Dou64, GC64, LC64, Mit64, NT64, Pal68, PM64, SR64a, Tox64]. Third [Ano64k, GP64, SPRM61]. Thomas [Sch63, Tel62]. Three [AAY65, Deh65, DP69, Duc65, KRS60, Nie68, Phi64, Phi65, RB67a, Sch61, SPB69, Zup65]. Three-Alpha [Duc65]. Three-Body [Deh65, DP69, Nie68, Phi64, RB67a, Sch61]. Three-Dimensional [SPB69]. Three-Nucleon [AAY65, Duc65]. Three-Particle [Zup65]. Threshold [Arn67, EFG^{+}61, Kaz67b, WSL^{+}61]. Thresholds [Cut61]. Ti [Hon68]. Time [Ber60a, Eis64, MB64, Ne’65, Phi64, Rom63, Tut68]. Time-Dependent [Ber60a, MB64]. Tin [Bal64a, Mei64, Par64, SK68a, SCGP64]. Tin-Lead [SCGP64]. Tokyo [Hos67b]. Tool [Mor67]. Topical [Fow64a]. Torque [Hei64]. Total [HW67, Löw60]. Trained [Hei64]. Trajectory [GMGL64]. Transistion [HR68, Jér68]. Transition [Adl68, AH68, Bec64, BGR64, BHMM64, Cal63, CGJY64, Col64, Geb64, HF68, Hon68, JHC64, LCW68, Lyk60, Mik68, Mot68, NT64, Par64, Per66, SK68b, SMR64, Tho68b, Wad64, EH68]. Transition-Metal [Adl68]. Transitions [Adl68, Han68, Hyl68, JB64, SW60, Tin64]. Transmission [Eyr62]. Transport [Bal60, Dre61b, EHD69, HL60, KY68, Nak60]. Transverse [Hei64, MKQ64]. Trapped [Jas60]. Treated [Löw62b]. Treatment [BS67, FT60, Mat63]. Triplet [SSW63]. Tube [Ber60b]. Tunnel [Sca64]. Tunneling [Adk64, Fis64, KTB64, Löw63b, O’K63, Row64, Sch64c, SJH64]. Turbulence [Kov60, Tat60]. Turbulent [Lyk60, Nap60]. Twisted [Tay60]. Two [BT64, Ger68, HS63, Hel67, Hyl63, Jun60, KS63b, Mey65, Mor67, Nig63, RÖ63, SK63, SML64]. Two-Dimensional [Jun60, SML64]. Two-Electron [BT64, HS63, Hyl63, KS63b, RÖ63, SK63]. Two-Nucleon [Mor67, Nig63]. Type [AH68, Mul60, Ber64, Goo64, Liv64]. Types [Men64a].

Ultrasonic [DP64, LS64b]. Unified [Che61, Fes64]. Uniform [Ago60, Kle68]. Unitary [IN66]. Universal [Wu64]. Universe [Har67]. University [Löw63a, Par60, Nes60]. Unpedagogical [Jos64b]. Unpolarized [Kes69]. Unsolved [Pip64b]. until [Gor64b]. Upper [Jac63, JHC64, RAG64, Wei63]. Use [HVW65]. Using [Coo67, Mus67, Pre63, TPL69, TPL73, War65].

V [Ano64e, Sca64, Weg64]. Vacancies [TW63]. Vacuum [GMGL64, KB67b, KB67a]. Valence [FM60]. Valence-Bond [FM60]. Values [BF63, Wei63]. Vanadium [NK64]. Vapor [HF68]. Variable [BB66a, Lyt65, Tut68]. Variables [Bel66, BB66b, BB68, BY64, CN68, Gud68, JP68]. Variation [DVB64]. Variational [FB60b, FB60c, SB67]. Variations [Cal63, Jac63]. Various [Ara60, Ros65]. Vector [Sch64d, Wu64]. VI [Ano64f]. Vibrational [Phe68, She60a, TB69, VK68]. Vibrations [Dav65, Har67, MV68, War68]. Vibronic [Lie60]. Vicinity [KKR65]. VII [Ano64g]. VIII [Ano64h]. Violent [BBS63]. Virial [GP64]. Virtual [HM65]. Viscoelastic [Pip64a]. Viscoelasticity [CN61, CN64]. Viscous [Has60]. Visible [Ewa65]. VO [Hyl68, PBS68, VBB68]. Volt [HW67, HW68]. Volume [BY64]. Vortex [DM64]. Vortices [PM64].

Wakes [Has60]. Wave [BM68, DH60, FKP60, Ger60, Gil63, Har63, HM65, KR60a, Löw60, Löw63c, Mac62, Mus67, Nes63, Pai60, Ran60b, RÖ63, Ric60, SB67, Sin63b, Sla65]. Wave-Function [Nes63, Sla65]. Wavefunctions [Löw64]. Wavelengths [Bea67]. Waves [Cab60, Ger60, Gro65, HLS60, Kap60, KL60, Lüs60, Pen65, Res60, Tol63]. Weak [Abr61, Kar60]. Weakly [Ara68]. Well [MHB67]. Whither [Jaf69]. Widths [Lan60, Sch64b]. Wigner [Ano62, BC62, Mos62]. Wires [Hei64, SDH64]. without [Che62, MOK64, Pai60]. Work [DG66, KK65, Men64b, Mic67, Wet61, Whi67].

X [Azá63, BC64, BB67, Bea67, Ewa65, PP63]. X-Ray [Azá63, BB67, Bea67, PP63]. Xenon [PP64].

Yang [Fin64]. Yields [FJMS66].

Zinc [Mas65]. Zone [Mer64].

References

Afzal:1969:SSI

[AAA69] S. A. Afzal, A. A. Z. Ahmad, and S. Ali. Systematic survey of the α–α interaction. Reviews of Modern Physics, 41 (1):247–273, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.247;   http://
rmp.aps.org/abstract/RMP/v41/i1/p247_1.

Aaron:1965:TTN

[AAY65] R. Aaron, R. D. Amado, and Y. Y. Yam. Theory of the three-nucleon system. Reviews of Modern Physics, 37(3): 516–518, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.516;   http://
rmp.aps.org/abstract/RMP/v37/i3/p516_1.

Altmann:1965:LHI

[AB65] S. L. Altmann and C. J. Bradley. Lattice harmonics II. Hexagonal close-packed lattice. Reviews of Modern Physics, 37(1):33–45, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.37.33;   http://
rmp.aps.org/abstract/RMP/v37/i1/p33_1.

Anderson:1961:PPI

[ABB^{+}61] Jerry A. Anderson, Vo X. Bang, Philip G. Burke, D. Duane Carmony, and Norbert Schmitz. Pion-pion interactions. Reviews of Modern Physics, 33(3):431–434, July 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.431;   http://rmp.aps.org/abstract/RMP/
v33/i3/p431_1.

Abraham:1961:SSS

[Abr61] Pais Abraham. On symmteries shared by strong and weak interactions. Reviews of Modern Physics, 33(3):493–497, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.493;   http://rmp.aps.org/abstract/RMP/
v33/i3/p493_1.

Altmann:1965:LHC

[AC65] S. L. Altmann and A. P. Cracknell. Lattice harmonics I. Cubic groups. Reviews of Modern Physics, 37(1):19–32, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.37.19;   http://rmp.aps.org/abstract/
RMP/v37/i1/p19_1.

Allen:1963:PCE

[ACG63] Leland C. Allen, Enrico Clementi, and H. M. Gladney. Pair correlation energies. Reviews of Modern Physics, 35(3): 465–473, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.465;   http://
rmp.aps.org/abstract/RMP/v35/i3/p465_1.

Alburger:1965:SAP

[AD65] D. E. Alburger and P. F. Donovan. Studies of alpha-particle emitting states of C^{12} observed in the reactions B^{10} (He^{3}, p)C^{12}(α)Be^{8} and C^{13} (He^{3}, α)C^{12}(α)Be^{8}. Reviews of Modern Physics, 37(3):448–449, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.448;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p448_1.

Adair:1961:SCC

[Ada61] Robert K. Adair. Some considerations concerning final-state interactions and the reaction K_{2}^{0} + p → Λ^{0} + Π^{+} + Π^{0}. Reviews of Modern Physics, 33(3):406–415, July 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.406;   http://rmp.aps.org/abstract/RMP/
v33/i3/p406_1.

Adair:1965:PAF

[Ada65] Robert K. Adair. Problems in the analyses of final-state resonances. Reviews of Modern Physics, 37(3):473–483, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.473;   http://rmp.aps.org/abstract/RMP/
v37/i3/p473_1.

Adkins:1964:MPT

[Adk64] C. J. Adkins. Multi-particle tunneling between superconductors. Reviews of Modern Physics, 36(1):211–213, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.211;   http://rmp.aps.org/abstract/RMP/
v36/i1/p211_1.

Adler:1968:MMN

[Adl68] David Adler. Mechanisms for metal-nonmental transitions in transition-metal oxides and sulfides. Reviews of Modern Physics, 40(4):714–736, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.714;   http://
rmp.aps.org/abstract/RMP/v40/i4/p714_1.

Adomian:1963:LSO

[Ado63] G. Adomian. Linear stochastic operators. Reviews of Modern Physics, 35(1):185–207, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.185;   http:
//rmp.aps.org/abstract/RMP/v35/i1/p185_1.

Alston:1961:PHR

[AFL61] Margaret H. Alston and Massimiliano Ferro-Luzzi. Pion-hyperon resonances. Reviews of Modern Physics, 33(3):416–426, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.416;   http://rmp.aps.org/abstract/RMP/
v33/i3/p416_1.

Agostinelli:1960:MAE

[Ago60] Cataldo Agostinelli. On the magnetodynamic and adiabatic equilibrium of a gaseous mass with a uniform rotational and gravitational motion. Reviews of Modern Physics, 32(4):941– 946, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.941;   http://
rmp.aps.org/abstract/RMP/v32/i4/p941_1.

Alexander:1968:SMT

[AH68] Michael N. Alexander and Donald F. Holcomb. Semiconductorto-metal transition in n-Type Group IV semiconductors. Reviews of Modern Physics, 40(4):815–829, October 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.815;   http://rmp.aps.org/abstract/RMP/
v40/i4/p815_1.

Aharoni:1962:TSD

[Aha62] Amikam Aharoni. Theoretical search for domain nucleation. Reviews of Modern Physics, 34(2):227–238, April 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.227;   http://rmp.aps.org/abstract/RMP/
v34/i2/p227_1.

Alpher:1960:SSF

[AHJW60] R. A. Alpher, H. Hurwitz, Jr., R. H. Johnson, and D. R. White. Some studies of free-surface mercury magnetohydrodynamics. Reviews of Modern Physics, 32(4):758–769, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.758;   http://rmp.aps.org/abstract/RMP/
v32/i4/p758_1.

Aizu:1962:PPF

[Aiz62] Kêitsiro Aizu. Polarization, pyroelectricity, and ferroelectricity of ionic crystals. Reviews of Modern Physics, 34(3): 550–576, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.550;   http://
rmp.aps.org/abstract/RMP/v34/i3/p550_1.

Allen:1960:BFI

[AK60] Leland C. Allen and Arnold M. Karo. Basis functions for ab initio calculations. Reviews of Modern Physics, 32(2):

275–285, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.275;   http://
rmp.aps.org/abstract/RMP/v32/i2/p275_1.

Anderson:1964:HST

[AK64] P. W. Anderson and Y. B. Kim. Hard superconductivity: Theory of the motion of Abrikosov flux lines. Reviews of Modern Physics, 36(1):39–43, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.39;   http://rmp.aps.org/abstract/RMP/v36/i1/p39_1.

Alfven:1960:CBN

[Alf60] H. Alfvén. Collision between a nonionized gas and a magnetized plasma. Reviews of Modern Physics, 32(4):710– 713, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.710;   http://
rmp.aps.org/abstract/RMP/v32/i4/p710_1.

Alfven:1965:ADM

[Alf65] H. Alfvén. Antimatter and the development of the metagalaxy. Reviews of Modern Physics, 37(4):652–665, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.652;   http://rmp.aps.org/abstract/RMP/
v37/i4/p652_1.

Altmann:1963:SPP

[Alt63] S. L. Altmann. Semidirect products and point groups. Reviews of Modern Physics, 35(3):641–644, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.641;   http://rmp.aps.org/abstract/RMP/
v35/i3/p641_1.

Andronikashvili:1966:QMM

[AM66] E. L. Andronikashvili and Yu. G. Mamaladze. Quantization of macroscopic motions and hydrodynamics of rotating helium II. Reviews of Modern Physics, 38(4):567–625, October 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.38.567;   http://rmp.aps.org/abstract/RMP/

v38/i4/p567_1. See erratum [AM67].

Andronikashvili:1967:EQM

[AM67] E. L. Andronikashvili and Yu. G. Mamaladze. Erratum: Quantization of Macroscopic Motions and Hydrodynamics of Rotating Helium II. Reviews of Modern Physics, 39(2):494, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.494;   http://rmp.aps.org/abstract/RMP/
v39/i2/p494_1. See [AM66].

Amaldi:1967:PPA

[Ama67] U. Amaldi, Jr. Pion-production amplitudes and nucleonnucleon interaction. Reviews of Modern Physics, 39(3): 649–656, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.649;   http://
rmp.aps.org/abstract/RMP/v39/i3/p649_1.

Anders:1962:MA

[And62] Edward Anders. Meteorite ages. Reviews of Modern Physics, 34(2):287–325, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.287;   http://
rmp.aps.org/abstract/RMP/v34/i2/p287_1.

Ando:1963:PFD

[And63] Tsuyoshi Ando. Properties of fermion density matrices. Reviews of Modern Physics, 35(3):690–702, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.690;   http://rmp.aps.org/abstract/RMP/
v35/i3/p690_1.

Anderson:1964:CCC

[And64] James L. Anderson. Coordinate conditions and canonical formalisms in gravitational theory. Reviews of Modern Physics, 36 (4):929–938, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.929;   http://
rmp.aps.org/abstract/RMP/v36/i4/p929_1.

Anderson:1966:CFS

[And66] P. W. Anderson. Considerations on the flow of superfluid helium. Reviews of Modern Physics, 38(2):298–310, April 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.298;   http://rmp.aps.org/abstract/RMP/
v38/i2/p298_1.

Anonymous:1961:I

[Ano61] Anonymous. Introduction. Reviews of Modern Physics, 33 (3):355, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.355;   http://
rmp.aps.org/abstract/RMP/v33/i3/p355_1.

Anonymous:1962:EPW

[Ano62] Anonymous. To Eugene Paul Wigner on his sixtieth birthday. Reviews of Modern Physics, 34(4):587–591, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.587;   http://rmp.aps.org/abstract/RMP/
v34/i4/p587_1.

Anonymous:1963:PSS

[Ano63] Anonymous. Participants in the Sanibel Symposium 14–19 January 1963. Reviews of Modern Physics, 35(3):416–420, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.416;   http://rmp.aps.org/abstract/RMP/
v35/i3/p416_1.

Anonymous:1964:CPSa

[Ano64a] Anonymous. Contributed papers for Session I. Reviews of Modern Physics, 36(1):351–352, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.351;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p351_1.

Anonymous:1964:CPSb

[Ano64b] Anonymous. Contributed papers for Session II. Reviews of Modern Physics, 36(1):357–361, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.357;   http://rmp.aps.org/abstract/RMP/
v36/i1/p357_1.

Anonymous:1964:CPSc

[Ano64c]  Anonymous.  Contributed papers for Session III.  Re-

views of Modern Physics, 36(1):391–393, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.391;   http://rmp.aps.org/abstract/RMP/
v36/i1/p391_1.

Anonymous:1964:CPSd

[Ano64d]  Anonymous.  Contributed papers for Session IV.  Re-

views of Modern Physics, 36(1):394–398, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.394;   http://rmp.aps.org/abstract/RMP/
v36/i1/p394_2.

Anonymous:1964:CPSe

[Ano64e] Anonymous. Contributed papers for Session V. Reviews of Modern Physics, 36(1):406–409, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.406;   http://rmp.aps.org/abstract/RMP/
v36/i1/p406_1.

Anonymous:1964:CPSf

[Ano64f]  Anonymous.  Contributed papers for Session VI.  Re-

views of Modern Physics, 36(1):437–440, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.437;   http://rmp.aps.org/abstract/RMP/
v36/i1/p437_1.

Anonymous:1964:CPSg

[Ano64g] Anonymous. Contributed papers for Session VII. Reviews of Modern Physics, 36(1):458–463, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.458;   http://rmp.aps.org/abstract/RMP/
v36/i1/p458_1.

Anonymous:1964:CPSh

[Ano64h] Anonymous. Contributed papers for Session VIII. Reviews of Modern Physics, 36(1):469–471, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.469;   http://rmp.aps.org/abstract/RMP/
v36/i1/p469_1.

Anonymous:1964:E

[Ano64i] Anonymous. Erratum. Reviews of Modern Physics, 36 (3):880, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.880;   http://
rmp.aps.org/abstract/RMP/v36/i3/p880_1.

Anonymous:1964:LP

[Ano64j] Anonymous. List of participants. Reviews of Modern Physics, 36(1):335–338, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.335;   http://
rmp.aps.org/abstract/RMP/v36/i1/p335_1.

Anonymous:1964:TIC

[Ano64k] Anonymous. Third international conference on the Mössbauer effect. Reviews of Modern Physics, 36(1):333–334, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.333;   http://rmp.aps.org/abstract/RMP/
v36/i1/p333_1.

Anonymous:1967:ACPa

[Ano67a] Anonymous. Abstracts of contributed papers in Session B. Reviews of Modern Physics, 39(3):574, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.574;   http://rmp.aps.org/abstract/RMP/
v39/i3/p574_1.

Anonymous:1967:ACPb

[Ano67b] Anonymous. Abstracts of contributed papers in Session C. Reviews of Modern Physics, 39(3):592–593, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.592;   http://rmp.aps.org/abstract/RMP/
v39/i3/p592_1.

Anonymous:1967:ACPc

[Ano67c] Anonymous. Abstracts of contributed papers in Session D. Reviews of Modern Physics, 39(3):648, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.648;   http://rmp.aps.org/abstract/RMP/
v39/i3/p648_1.

Anonymous:1967:ACPd

[Ano67d] Anonymous. Abstracts of contributed papers in Session E. Reviews of Modern Physics, 39(3):699, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.699;   http://rmp.aps.org/abstract/RMP/
v39/i3/p699_1.

Anonymous:1967:ACPe

[Ano67e] Anonymous. Abstracts of contributed papers in Session F. Reviews of Modern Physics, 39(3):716–717, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.716;   http://rmp.aps.org/abstract/RMP/
v39/i3/p716_1.

Anonymous:1967:I

[Ano67f] Anonymous. Introduction. Reviews of Modern Physics, 39 (3):495–497, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.495;   http://
rmp.aps.org/abstract/RMP/v39/i3/p495_1.

Anonymous:1968:A

[Ano68] Anonymous. Announcement. Reviews of Modern Physics, 40(2):237, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.237;   http://
rmp.aps.org/abstract/RMP/v40/i2/p237_1.

Anonymous:1969:E

[Ano69] Anonymous. Editorial. Reviews of Modern Physics, 41(4): 575, October 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.575;   http://
rmp.aps.org/abstract/RMP/v41/i4/p575_1.

Arai:1960:GAV

[Ara60] Tadashi Arai. General analysis of various methods of atoms in molecules. Reviews of Modern Physics, 32(2):370–400, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.370;   http://rmp.aps.org/abstract/RMP/
v32/i2/p370_1.

Arai:1968:CAL

[Ara68] Tadashi Arai. Correlations among localized and weakly metallic electrons. Reviews of Modern Physics, 40(4):800– 801, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.800;   http://
rmp.aps.org/abstract/RMP/v40/i4/p800_1.

Arndt:1967:PSR

[Arn67] Richard A. Arndt. Phase-shift representation for nucleon-nucleon scattering above pion-production threshold. Reviews of Modern Physics, 39(3):710–715, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.710;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p710_1.

Anex:1960:MRM

[AS60] Basil G. Anex and William T. Simpson. Metallic reflection from molecular crystals. Reviews of Modern Physics, 32 (2):466–476, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.466;   http://
rmp.aps.org/abstract/RMP/v32/i2/p466_1.

Aldridge:1965:R

[AWY65] J. P. Aldridge III, B. H. Wildenthal, and D. H. Youngblood. The He^{3}(He^{3}, 2p)He^{4} reaction. Reviews of Modern Physics, 37(3):430–432, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.430;   http://
rmp.aps.org/abstract/RMP/v37/i3/p430_1.

Azaroff:1963:TEF

[Azá63] Leonid V. Azároff. Theory of extended fine structure of X-ray absorption edges. Reviews of Modern Physics, 35(4):1012–1021, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.35.1012;   http://rmp.aps.org/
abstract/RMP/v35/i4/p1012_1.

Babb:1963:PSE

[Bab63] Stanley E. Babb, Jr. Parameters in the Simon equation relating pressure and melting temperature. Reviews of Modern Physics, 35(2):400–413, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.400;   http://
rmp.aps.org/abstract/RMP/v35/i2/p400_1.

Booth:1961:DMP

[BAC61] Norman E. Booth, Alexander Abashian, and Kenneth M. Crowe. Double meson production in proton-deuteron collisions. Reviews of Modern Physics, 33(3):393–394, July 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.393;   http://rmp.aps.org/abstract/RMP/
v33/i3/p393_1.

Baird:1968:DDB

[Bai68] M. E. Baird. Determination of dielectric behavior at low frequencies from measurements of anomalous charging and discharging currents. Reviews of Modern Physics, 40(1):219– 227, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.219;   http://
rmp.aps.org/abstract/RMP/v40/i1/p219_1.

Balescu:1960:TEP

[Bal60] R. Balescu. Transport equation of a plasma. Reviews of Modern Physics, 32(4):719–721, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.719;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p719_1.

Baldwin:1964:CFH

[Bal64a] J. P. Baldwin. Critical fields and hysteresis effects in tin films. Reviews of Modern Physics, 36(1):317–320, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.317;   http://rmp.aps.org/abstract/RMP/
v36/i1/p317_1.

Baltensperger:1964:SRS

[Bal64b] W. Baltensperger. Short-range spin susceptibility of superconductors. Reviews of Modern Physics, 36(1):157–159, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.157;   http://rmp.aps.org/abstract/RMP/
v36/i1/p157_1.

Bardeen:1962:CFC

[Bar62a] John Bardeen. Critical fields and currents in superconductors. Reviews of Modern Physics, 34(4):667–681, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.667;   http://rmp.aps.org/abstract/RMP/
v34/i4/p667_1.

Bargmann:1962:RRG

[Bar62b] V. Bargmann. On the representations of the rotation group. Reviews of Modern Physics, 34(4):829–845, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.829;   http://rmp.aps.org/abstract/RMP/
v34/i4/p829_1.

Barker:1962:DNP

[Bar62c] W. A. Barker. Dynamic nuclear polarization. Reviews of Modern Physics, 34(2):173–185, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.173;   http:
//rmp.aps.org/abstract/RMP/v34/i2/p173_1.

Barnett:1963:MMC

[Bar63] Michael P. Barnett. Mechanized molecular calculations — the POLYATOM system. Reviews of Modern Physics, 35 (3):571–572, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.571;   http://
rmp.aps.org/abstract/RMP/v35/i3/p571_1.

Barger:1968:RPP

[Bar68] V. Barger. Regge poles and SU(3): Particle classification and high-energy scattering. Reviews of Modern Physics, 40 (1):129–139, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.129;   http://
rmp.aps.org/abstract/RMP/v40/i1/p129_1.

Basri:1965:OFE

[Bas65] Saul A. Basri. Operational foundation of Einstein’s General Theory of Relativity. Reviews of Modern Physics, 37 (2):288–315, April 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.288;   http://
rmp.aps.org/abstract/RMP/v37/i2/p288_1.

Bernardes:1962:LS

[BB62] N. Bernardes and D. F. Brewer. Liquid and solid He^{3}. Reviews of Modern Physics, 34(2):190–214, April 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.190;   http://rmp.aps.org/abstract/RMP/
v34/i2/p190_1.

Bilwes:1965:SLR

[BB65] R. Bilwes and B. Bourotte. Study of the Li^{7} +d → 2α + n reaction at low energy. Reviews of Modern Physics, 37 (3):458–459, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.458;   http://
rmp.aps.org/abstract/RMP/v37/i3/p458_1.

Bohm:1966:PSM

[BB66a] D. Bohm and J. Bub. A proposed solution of the measurement problem in quantum mechanics by a hidden variable theory. Reviews of Modern Physics, 38(3):453–469, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.453;   http://rmp.aps.org/abstract/RMP/
v38/i3/p453_1.

Bohm:1966:RPJ

[BB66b] D. Bohm and J. Bub. A refutation of the proof by Jauch and Piron that hidden variables can be excluded in quantum mechanics. Reviews of Modern Physics, 38(3):470–475, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.470;   http://rmp.aps.org/abstract/RMP/

v38/i3/p470_1. See errata [JP68, BB68].

Bearden:1967:RXR

[BB67] J. A. Bearden and A. F. Burr. Reevaluation of X-ray atomic energy levels. Reviews of Modern Physics, 39(1): 125–142, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.125;   http://
rmp.aps.org/abstract/RMP/v39/i1/p125_1.

Bohm:1968:HVR

[BB68] D. Bohm and J. Bub. On hidden variables — a reply to comments by Jauch and Piron and by Gudder. Reviews of Modern Physics, 40(1):235–236, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.235;   http:
//rmp.aps.org/abstract/RMP/v40/i1/p235_1. See [BB66b,

JP68].

Beretvas:1967:MCP

[BBD^{+}67] A. Beretvas, N. E. Booth, C. Dolnick, R. J. Esterling, R. E. Hill, J. Scheid, D. Sherden, and A. Yokosawa. Measurements of C_{NN} in proton-proton scattering. Reviews of Modern Physics, 39(3):536–537, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.536;   http://
rmp.aps.org/abstract/RMP/v39/i3/p536_1.

Branscomb:1969:E

[BBJ^{+}69] Lewis M. Branscomb, S. J. Buchsbaum, J. D. Jackson, A. K. Kerman, J. Krumhansl, and R. Novick. Editorial. Reviews of Modern Physics, 41(1):1–2, January 1969. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
41.1;   http://rmp.aps.org/abstract/RMP/v41/i1/p1_1.

Burbidge:1963:EOV

[BBS63] G. R. Burbidge, E. M. Burbidge, and A. R. Sandage. Evidence for the occurrence of violent events in the nuclei of galaxies. Reviews of Modern Physics, 35(4):947–972, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.947;   http://rmp.aps.org/abstract/RMP/
v35/i4/p947_1.

Boys:1960:MPC

[BC60] S. F. Boys and G. B. Cook. Mathematical problems in the complete quantum predictions of chemical phenomena. Reviews of Modern Physics, 32(2):285–295, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.285;   http://rmp.aps.org/abstract/RMP/
v32/i2/p285_1.

Brittin:1962:WDF

[BC62] Wesley E. Brittin and Willard R. Chappell. The Wigner distribution function and second quantization in phase space. Reviews of Modern Physics, 34(4):620–627, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.620;   http://rmp.aps.org/abstract/RMP/
v34/i4/p620_1.

Batterman:1964:DDX

[BC64] Boris W. Batterman and Henderson Cole. Dynamical diffraction of X rays by perfect crystals. Reviews of Modern Physics, 36(3):681–717, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.681;   http://
rmp.aps.org/abstract/RMP/v36/i3/p681_1.

Baltay:1961:PHH

[BCF^{+}61] C. Baltay, H. Courant, W. J. Fickinger, E. C. Fowler, H. L. Kraybill, J. Sandweiss, J. R. Sanford, D. L. Stonehill, and H. D. Taft. Production of hyperons in hydrogen by positive pions. Reviews of Modern Physics, 33(3):374–381, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.374;   http://rmp.aps.org/abstract/RMP/
v33/i3/p374_1.

Bradley:1968:MGT

[BD68] C. J. Bradley and B. L. Davies. Magnetic groups and their corepresentations. Reviews of Modern Physics, 40(2): 359–379, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.359;   http://
rmp.aps.org/abstract/RMP/v40/i2/p359_1.

Behrends:1962:ESG

[BDFL62a] R. E. Behrends, J. Dreitlein, C. Fronsdal, and B. W. Lee. Erratum: Simple Groups and Strong Interaction Symmetries. Reviews of Modern Physics, 34(3):584, July 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.584;   http://rmp.aps.org/abstract/RMP/

v34/i3/p584_1. See [BDFL62b].

Behrends:1962:SGS

[BDFL62b] R. E. Behrends, J. Dreitlein, C. Fronsdal, and W. Lee. Simple groups and strong interaction symmetries. Reviews of Modern Physics, 34(1):1–40, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.34.1;
 http://rmp.aps.org/abstract/RMP/v34/i1/p1_1. See erra-

tum [BDFL62a].

Bratoz:1960:SRR

[BDRA60] S. Bratoz, R. Daudel, M. Roux, and M. Allavena. Some recent results concerning the electronic density and the force constants of small molecules. Reviews of Modern Physics, 32 (2):412–417, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.412;   http://
rmp.aps.org/abstract/RMP/v32/i2/p412_1.

Bean:1964:MHF

[Bea64] Charles P. Bean. Magnetization of high-field superconductors. Reviews of Modern Physics, 36(1):31–39, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.31;   http://rmp.aps.org/abstract/RMP/v36/i1/p31_1.

Bearden:1967:XRW

[Bea67] J. A. Bearden. X-ray wavelengths. Reviews of Modern Physics, 39(1):78–124, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.39.78;   http://
rmp.aps.org/abstract/RMP/v39/i1/p78_1.

Beck:1964:BSG

[Bec64] Paul A. Beck. Band structure of the d-group transition element alloys. Reviews of Modern Physics, 36(1):394, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.394;   http://rmp.aps.org/abstract/RMP/
v36/i1/p394_1.

Bell:1966:PHV

[Bel66] John S. Bell. On the problem of hidden variables in quantum mechanics. Reviews of Modern Physics, 38(3):447–452, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.447;   http://rmp.aps.org/abstract/RMP/
v38/i3/p447_1.

Bell:1967:CSN

[Bel67] George I. Bell. Cross sections for nucleosynthesis in stars and bombs. Reviews of Modern Physics, 39(1):59–68, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.39.59;   http://rmp.aps.org/abstract/
RMP/v39/i1/p59_1.

Berry:1960:TDM

[Ber60a] R. Stephen Berry. Time-dependent measurements and molecular structure: Ozone. Reviews of Modern Physics, 32(2): 447–454, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.447;   http://
rmp.aps.org/abstract/RMP/v32/i2/p447_1.

Bershader:1960:SRE

[Ber60b] Daniel Bershader. Some remarks on electric-discharge shocktube diagnostics. Reviews of Modern Physics, 32(4):780– 784, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.780;   http://
rmp.aps.org/abstract/RMP/v32/i4/p780_1.

Bergmann:1961:OGR

[Ber61] Peter G. Bergmann. Observables in General Relativity. Reviews of Modern Physics, 33(4):510–514, October 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.510;   http://rmp.aps.org/abstract/RMP/
v33/i4/p510_1.

Berlincourt:1964:TIS

[Ber64] T. G. Berlincourt. Type II superconductivity. Reviews of Modern Physics, 36(1):19–26, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.19;   http://rmp.aps.org/abstract/RMP/v36/i1/p19_1.

Berenyi:1968:RRE

[Ber68] D. Berényi. Recent results on electron capture. Reviews of Modern Physics, 40(2):390–398, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.390;   http:
//rmp.aps.org/abstract/RMP/v40/i2/p390_1.

Bazley:1963:EBE

[BF63] Norman W. Bazley and David W. Fox. Error bounds for expectation values. Reviews of Modern Physics, 35(3):712–715, July 1963.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.712;   http://rmp.aps.org/abstract/RMP/
v35/i3/p712_1.

Boato:1964:ETM

[BGR64] G. Boato, G. Gallinaro, and C. Rizzuto. Effect of transition metal impurities on the critical temperature of nontransition metals. Reviews of Modern Physics, 36(1):162–164, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.162;   http://rmp.aps.org/abstract/RMP/
v36/i1/p162_1.

Bauer:1965:SCR

[BHK^{+}65] R. W. Bauer, G. Heymann, W. Kossler, N. S. Wall, and C. R. Gruhn. Study of the Ca^{40}(α, 2α) reactions. Reviews of Modern Physics, 37(3):369–371, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.369;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p369_1.

Bucher:1964:SEP

[BHMM64] E. Bucher, F. Heiniger, J. Muheim, and J. Muller. Superconductivity and electronic properties of transition metal alloys. Reviews of Modern Physics, 36(1):146–149, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.146;   http://rmp.aps.org/abstract/RMP/
v36/i1/p146_1.

Bander:1966:GTHa

[BI66a] M. Bander and C. Itzykson. Group theory and the hydrogen atom (I). Reviews of Modern Physics, 38(2):330–345, April 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.330;   http://rmp.aps.org/abstract/RMP/
v38/i2/p330_1.

Bander:1966:GTHb

[BI66b] M. Bander and C. Itzykson. Group theory and the hydrogen atom (II). Reviews of Modern Physics, 38(2):346–358, April 1966.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.346;   http://rmp.aps.org/abstract/RMP/
v38/i2/p346_1.

Biermann:1960:RBP

[Bie60] L. Biermann. Relations between plasma physics and astrophysics. Reviews of Modern Physics, 32(4):1008–1011, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1008;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1008_1.

Bird:1964:MCI

[Bir64] Joseph F. Bird. Massive condensations in interstellar matter and stellar associations. Reviews of Modern Physics, 36 (3):717–747, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.717;   http://
rmp.aps.org/abstract/RMP/v36/i3/p717_1.

Bunch:1965:DRM

[BKF65] S. M. Bunch, C. C. Kim, and H. H. Forster. D(p, 2p)n reaction at 31 MeV. Reviews of Modern Physics, 37(3):528–530, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.528;   http://rmp.aps.org/abstract/RMP/
v37/i3/p528_1.

Block:1962:AFL

[BKR62] H. D. Block, B. W. Knight, Jr., and F. Rosenblatt. Analysis of a four-layer series-coupled perceptron. II. Reviews of Modern Physics, 34(1):135–142, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.135;   http:
//rmp.aps.org/abstract/RMP/v34/i1/p135_1.

Bessis:1963:ICP

[BLBM63] N. Bessis, H. Lefebvre-Brion, and C. M. Moser. Importance of core polarization on the calculation of nuclear moments from hyperfine structure data. Nuclear moments of C^{11}. Reviews of Modern Physics, 35(3):548–550, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.548;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p548_1.

Block:1962:PMB

[Blo62] H. D. Block. The perceptron: a model for brain functioning. I. Reviews of Modern Physics, 34(1):123–135, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.123;   http://rmp.aps.org/abstract/RMP/
v34/i1/p123_1.

Barut:1968:POL

[BM68] A. O. Barut and S. Malin. Position operators and localizability of quantum systems described by finite- and infinitedimensional wave equations. Reviews of Modern Physics, 40 (3):632–651, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.632;   http://
rmp.aps.org/abstract/RMP/v40/i3/p632_1.

Barker:1964:LTD

[BNM^{+}64] W. A. Barker, Mohan Lal Narchal, Sushil Misra, Albert J. Bevolo, Francis X. Haas, Claire M. Birkbeck, Barbara M. Kuehne, and John C. Witte. Low-temperature dynamic nuclear polarization. Reviews of Modern Physics, 36(3):872–879, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.872;   http://rmp.aps.org/abstract/RMP/
v36/i3/p872_1.

Boley:1964:ACR

[Bol64]  Forrest I. Boley.  Atmospheric Cerenkov radiation from

cosmic-ray air showers. Reviews of Modern Physics, 36(3): 792–808, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.792;   http://
rmp.aps.org/abstract/RMP/v36/i3/p792_1.

Born:1964:MMB

[Bor64] Max Born. Message from Max Born. Reviews of Modern Physics, 36(2):509, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.509;   http:/
/rmp.aps.org/abstract/RMP/v36/i2/p509_1. Open letter to

Robert Oppenheimer.

Boys:1960:CSM

[Boy60] S. F. Boys. Construction of some molecular orbitals to be approximately invariant for changes from one molecule to another. Reviews of Modern Physics, 32(2):296–299, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.296;   http://rmp.aps.org/abstract/RMP/
v32/i2/p296_1.

Brandow:1967:LCE

[Bra67] Baird H. Brandow. Linked-cluster expansions for the nuclear many-body problem. Reviews of Modern Physics, 39 (4):771–828, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.771;   http://
rmp.aps.org/abstract/RMP/v39/i4/p771_1.

Breit:1962:ANN

[Bre62] G. Breit. Aspects of nucleon-nucleon scattering theory. Reviews of Modern Physics, 34(4):766–812, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.766;   http://rmp.aps.org/abstract/RMP/

v34/i4/p766_1. See errata [Bre63].

Breit:1963:EAN

[Bre63] G. Breit. Erratum: Aspects of Nucleon-Nucleon Scattering Theory. Reviews of Modern Physics, 35(3):734, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.734;   http://rmp.aps.org/abstract/RMP/
v35/i3/p734_2. See [Bre62].

Breit:1964:DCS

[Bre64] G. Breit. Definitions of compound states. Reviews of Modern Physics, 36(4):1071–1075, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.36.1071;
http://rmp.aps.org/abstract/RMP/v36/i4/p1071_1.

Breit:1967:PPP

[Bre67] G. Breit. Phenomenological phase-parameter fits to N–N data at intermediate energies. Reviews of Modern Physics, 39(3):560–570, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.560;   http://
rmp.aps.org/abstract/RMP/v39/i3/p560_1.

Broer:1960:MAH

[Bro60a] L. J. F. Broer. Mechanical analogy of the Hall effect. Reviews of Modern Physics, 32(4):888, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.888;   http://rmp.aps.org/abstract/RMP/
v32/i4/p888_1.

Broer:1960:MCP

[Bro60b] L. J. F. Broer. Motion of a charged particle in a slowly increasing magnetic field. Reviews of Modern Physics, 32(4): 742–743, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.742;   http://
rmp.aps.org/abstract/RMP/v32/i4/p742_1.

Brush:1961:FIS

[Bru61] S. G. Brush. Functional integrals and statistical physics. Reviews of Modern Physics, 33(1):79–92, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
33.79;   http://rmp.aps.org/abstract/RMP/v33/i1/p79_1.

Brush:1967:HLI

[Bru67] Stephen G. Brush. History of the Lenz–Ising model. Reviews of Modern Physics, 39(4):883–893, October 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.883;   http://rmp.aps.org/abstract/RMP/
v39/i4/p883_1.

Burke:1962:LES

[BS62] P. G. Burke and Kenneth Smith. The low-energy scattering of electrons and positrons by hydrogen atoms. Reviews of Modern Physics, 34(3):458–502, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.458;   http:
//rmp.aps.org/abstract/RMP/v34/i3/p458_1.

Bardeen:1964:ICS

[BS64] John Bardeen and Roland W. Schmitt. International Conference on the Science of Superconductivity. Reviews of Modern Physics, 36(1):1–2, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.1;   http://rmp.aps.org/abstract/RMP/v36/i1/p1_1.

Bohm:1967:CTL

[BS67] D. Bohm and B. Salt. Collective treatment of liquid helium. Reviews of Modern Physics, 39(4):894–910, October 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.894;   http://rmp.aps.org/abstract/RMP/
v39/i4/p894_1.

Barash-Schmidt:1969:RPP

[BSBGP^{+}69] Naomi Barash-Schmidt, Angela Barbaro-Galtieri, Leroy R. Price, Arthur H. Rosenfeld, Paul Söding, Charles G. Wohl, Matts Roos, and Gianni Conforto (Particle Data Group). Review of particle properties. Reviews of Modern Physics, 41(1): 109–192, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.109;   http://
rmp.aps.org/abstract/RMP/v41/i1/p109_1.

Bhatia:1964:SEA

[BT64] A. K. Bhatia and A. Temkin. Symmetric Euler-angle decomposition of the two-electron fixed-nucleus problem. Reviews of Modern Physics, 36(4):1050–1064, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.36.1050;
http://rmp.aps.org/abstract/RMP/v36/i4/p1050_1.

Bacher:1965:FSI

[BT65] A. D. Bacher and T. A. Tombrello. p–He^{4} final-state interaction in He^{3}(He^{3}, 2 p)He^{4}. Reviews of Modern Physics, 37(3):433–436, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.433;   http://
rmp.aps.org/abstract/RMP/v37/i3/p433_1.

Bunge:1967:PA

[Bun67] Mario Bunge. Physical axiomatics. Reviews of Modern Physics, 39(2):463–474, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.463;   http://
rmp.aps.org/abstract/RMP/v39/i2/p463_1.

Burgers:1960:MCI

[Bur60] J. M. Burgers. Motion of a completely ionized gas across a magnetic field in the presence of an electric force. Reviews of Modern Physics, 32(4):868–880, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.868;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p868_1.

Burgers:1963:EMP

[Bur63a] J. M. Burgers. Erratum: The Measuring Process in Quantum Theory. Reviews of Modern Physics, 35(4):1032, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.35.1032;   http://rmp.aps.org/
abstract/RMP/v35/i4/p1032_1. See [Bur63b].

Burgers:1963:MPQ

[Bur63b] J. M. Burgers. The measuring process in quantum theory. Reviews of Modern Physics, 35(1):145–150, January 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.145;   http://rmp.aps.org/abstract/RMP/

v35/i1/p145_1. See erratum [Bur63a].

Brussaard:1962:AFN

[BvdH62] P. J. Brussaard and H. C. van de Hulst. Approximation formulas for nonrelativistic Bremsstrahlung and average Gaunt factors for a Maxwellian electron gas. Reviews of Modern Physics, 34(3):507–520, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.507;   http://
rmp.aps.org/abstract/RMP/v34/i3/p507_1.

Brill:1957:ING

[BW57] Dieter R. Brill and John A. Wheeler. Interaction of neutrinos and gravitational fields. Reviews of Modern Physics, 29(3):465–479, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.29.465;   http:
//rmp.aps.org/abstract/RMP/v29/i3/p465_1.  See errata

[BW61].

Brill:1961:EIN

[BW61] Dieter R. Brill and John A. Wheeler. Errata: Interaction of Neutrinos and Gravitational Fields. Reviews of Modern Physics, 33(4):623–624, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.623;   http://
rmp.aps.org/abstract/RMP/v33/i4/p623_2. See [BW57].

Byers:1964:PRI

[BY64] N. Byers and C. N. Yang. Physical regions in invariant variables for n particles and the phase-space volume element. Reviews of Modern Physics, 36(2):595–609, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.595;   http://rmp.aps.org/abstract/RMP/
v36/i2/p595_1.

Cabannes:1960:ASS

[Cab60] H. Cabannes. Attached stationary shock waves in ionized gases. Reviews of Modern Physics, 32(4):973–976, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.973;   http://rmp.aps.org/abstract/RMP/
v32/i4/p973_1.

Callan:1963:SMC

[Cal63] Edwin J. Callan. Sensitivity of L_{1} L_{2,3} M_{4,5} Coster–Kronig transition rates to variations in energy levels and screening parameters. Reviews of Modern Physics, 35(3):524–527, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.35.524;   http://rmp.aps.org/abstract/RMP/
v35/i3/p524_1.

Carruthers:1961:TTC

[Car61] Peter Carruthers. Theory of thermal conductivity of solids at low temperatures. Reviews of Modern Physics, 33(1):92–138, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.33.92;   http://rmp.aps.org/abstract/
RMP/v33/i1/p92_1.

Carlson:1965:PDD

[Car65] Richard F. Carlson. Proton disintegration of the deuteron — a test of the chew-low extrapolation method. Reviews of Modern Physics, 37(3):531–532, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.531;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p531_1.

Case:1964:AMD

[Cas64] K. M. Case. An approximation method for diffraction problems. Reviews of Modern Physics, 36(2):669–679, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.669;   http://rmp.aps.org/abstract/RMP/
v36/i2/p669_1.

Catillon:1967:SCP

[Cat67] P. Catillon. Spin-correlation parameters in p–p scattering. Reviews of Modern Physics, 39(3):534–535, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.534;   http://rmp.aps.org/abstract/RMP/
v39/i3/p534_1.

Cody:1964:TCN

[CC64] G. D. Cody and R. W. Cohen. Thermal conductivity of Nb_{3}Sn. Reviews of Modern Physics, 36(1):121–123, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.121;   http://rmp.aps.org/abstract/RMP/
v36/i1/p121_1.

Cody:1964:FAD

[CCM64] G. D. Cody, G. W. Cullen, and J. P. Mcevoy, Jr. Field and angular dependence of critical currents in Nb_{3}Sn II. Reviews of Modern Physics, 36(1):95–97, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.95;   http://rmp.aps.org/abstract/RMP/v36/i1/p95_1.

Cohen:1963:OPE

[CD63] M. Cohen and A. Dalgarno. On the orthogonality problem for excited states. Reviews of Modern Physics, 35(3): 506–507, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.506;   http://
rmp.aps.org/abstract/RMP/v35/i3/p506_1.

Cohen:1965:OKF

[CD65] E. Richard Cohen and Jesse W. M. Dumond. Our knowledge of the fundamental constants of physics and chemistry in 1965. Reviews of Modern Physics, 37(4):537–594, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.537;   http://rmp.aps.org/abstract/RMP/
v37/i4/p537_1.

Chou:1967:MTE

[CD67] Tsu-Teh Chou and Max Dresden. S-matrix theory of electromagnetic interactions. Reviews of Modern Physics, 39(1): 143–166, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.143;   http://
rmp.aps.org/abstract/RMP/v39/i1/p143_1.

Caldwell:1963:OR

[CE63] Dennis J. Caldwell and Henry Eyring. Optical rotation. Reviews of Modern Physics, 35(3):577–586, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.577;   http://rmp.aps.org/abstract/RMP/
v35/i3/p577_1.

Colgate:1960:HEE

[CFH60] S. A. Colgate, H. P. Furth, and F. O. Halliday. Hydromagnetic equilibrium experiments with liquid and solid sodium. Reviews of Modern Physics, 32(4):744–747, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.744;   http://rmp.aps.org/abstract/RMP/
v32/i4/p744_1.

Cohen:1963:LNS

[CFMM63] B. L. Cohen, R. H. Fulmer, A. L. McCarthy, and P. Mukherjee. Location of neutron single-particle levels from stripping reactions. Reviews of Modern Physics, 35(2):332–334, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.332;   http://rmp.aps.org/abstract/RMP/
v35/i2/p332_1.

Clogston:1964:OPK

[CGJY64] A. M. Clogston, A. C. Gossard, V. Jaccarino, and Y. Yafet. Orbital paramagnetism and the Knight shift in transition metal superconductors. Reviews of Modern Physics, 36(1): 170–175, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.170;   http://
rmp.aps.org/abstract/RMP/v36/i1/p170_1.

Clinton:1960:FCE

[CH60] William L. Clinton and Walter C. Hamilton. Force curves for excited electronic states. Reviews of Modern Physics, 32 (2):422–425, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.422;   http://
rmp.aps.org/abstract/RMP/v32/i2/p422_1.

Chapman:1960:IPP

[Cha60] Sydney Chapman. Idealized problems of plasma dynamics relating to geomagnetic storms. Reviews of Modern Physics, 32 (4):919–933, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.919;   http://
rmp.aps.org/abstract/RMP/v32/i4/p919_1.

Chang:1967:MLR

[Cha67] Tai Yup Chang. Moderately long-range interatomic forces. Reviews of Modern Physics, 39(4):911–942, October 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.911;   http://rmp.aps.org/abstract/RMP/
v39/i4/p911_1.

Chew:1961:UDA

[Che61] Geoffrey F. Chew. A unified dynamical approach to high- and low-energy strong interactions. Reviews of Modern Physics, 33(3):467–470, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.467;   http://
rmp.aps.org/abstract/RMP/v33/i3/p467_1.

Chew:1962:MTS

[Che62] Geoffrey F. Chew. S-matrix theory of strong interactions without elementary particles. Reviews of Modern Physics, 34 (3):394–401, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.394;   http://
rmp.aps.org/abstract/RMP/v34/i3/p394_1.

Chen:1963:POI

[Che63] Tien Chi Chen. Polynomial orthogonality and integration quadratures. Reviews of Modern Physics, 35(3):569–570, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.569;   http://rmp.aps.org/abstract/RMP/
v35/i3/p569_1.

Cheng:1967:NNS

[Che67] David Cheng. Nucleon-nucleon scattering experiments at intermediate energies in Berkeley. Reviews of Modern Physics, 39(3):526–530, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.526;   http://
rmp.aps.org/abstract/RMP/v39/i3/p526_1.

Chiu:1969:NSH

[Chi69] Charles B. Chiu. Nonforward scattering of hadrons and highenergy phenomenology. Reviews of Modern Physics, 41(4):

640–668, October 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.640;   http://
rmp.aps.org/abstract/RMP/v41/i4/p640_1.

Chopra:1961:EIR

[Cho61a] K. P. Chopra. Errata: Interactions of Rapidly Moving Bodies in Terrestrial Atmosphere. Reviews of Modern Physics, 33 (4):624–625, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.624;   http://
rmp.aps.org/abstract/RMP/v33/i4/p624_1. See [Cho61b].

Chopra:1961:IRM

[Cho61b] K. P. Chopra. Interactions of rapidly moving bodies in terrestrial atmosphere. Reviews of Modern Physics, 33(2): 153–189, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.153;   http:
//rmp.aps.org/abstract/RMP/v33/i2/p153_1.  See errata

[Cho61a].

Christy:1964:CSP

[Chr64] Robert F. Christy. The calculation of stellar pulsation. Reviews of Modern Physics, 36(2):555–571, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.555;   http://rmp.aps.org/abstract/RMP/
v36/i2/p555_1.

Cindro:1966:SFN

[Cin66] N. Cindro. A survey of fast-neutron reactions. Reviews of Modern Physics, 38(3):391–446, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.38.391;   http:
//rmp.aps.org/abstract/RMP/v38/i3/p391_1.

Coulson:1967:SAL

[CJ67] C. A. Coulson and A. Joseph. Self-adjoint ladder operators. II. Reviews of Modern Physics, 39(4):838–849, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.39.838;   http://rmp.aps.org/abstract/RMP/
v39/i4/p838_1.

Currie:1963:ERI

[CJS63a] D. G. Currie, T. F. Jordan, and E. C. G. Sudarshan. Erratum: Relativistic Invariance and Hamiltonian Theories of Interacting Particles. Reviews of Modern Physics, 35(4):1032, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.35.1032;   http://rmp.aps.org/
abstract/RMP/v35/i4/p1032_2. See [CJS63b].

Currie:1963:RIH

[CJS63b] D. G. Currie, T. F. Jordan, and E. C. G. Sudarshan. Relativistic invariance and Hamiltonian theories of interacting particles. Reviews of Modern Physics, 35(2):350–375, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.350;   http://rmp.aps.org/abstract/RMP/

v35/i2/p350_1. See erratum [CJS63a].

Coleman:1963:DSF

[CLS63] A. J. Coleman, P.-O. Löwdin, and Darwin W. Smith. Discussion on “Structure of Fermion Density Matrices”. Reviews of Modern Physics, 35(3):687–689, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.687;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p687_1. See [Col63].

Csonka:1967:NDS

[CMS67] Paul L. Csonka, Michael J. Moravcsik, and Michael D. Scadron. Non-dynamical structure of photoproduction processes. Reviews of Modern Physics, 39(1):178–202, January 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.178;   http://rmp.aps.org/abstract/RMP/
v39/i1/p178_1.

Coleman:1961:FLV

[CN61] Bernard D. Coleman and Walter Noll. Foundations of linear viscoelasticity. Reviews of Modern Physics, 33(2):239– 249, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.33.239;   http:/
/rmp.aps.org/abstract/RMP/v33/i2/p239_1. See erratum

[CN64].

Coleman:1964:EFL

[CN64] B. D. Coleman and W. Noll. Erratum: Foundations of Linear Viscoelasticity. Reviews of Modern Physics, 36(4):1103, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.1103;   http://rmp.aps.org/
abstract/RMP/v36/i4/p1103_2. See [CN61].

Calkin:1967:ESF

[CN67] M. G. Calkin and P. J. Nicholson. Electrodynamics of a semiclassical free-electron gas. Reviews of Modern Physics, 39 (2):361–372, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.361;   http://
rmp.aps.org/abstract/RMP/v39/i2/p361_1.

Carruthers:1968:PAV

[CN68] P. Carruthers and Michael Martin Nieto. Phase and angle variables in quantum mechanics. Reviews of Modern Physics, 40(2):411–440, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.411;   http://
rmp.aps.org/abstract/RMP/v40/i2/p411_1.

Cohen:1964:RDR

[CNO64] S. G. Cohen, I. Nowik, and S. Ofer. Recent developments in rare-earth Mössbauer studies. II. Reviews of Modern Physics, 36(1):378–385, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.378;   http://
rmp.aps.org/abstract/RMP/v36/i1/p378_1.

Cohen:1964:ESS

[Coh64a] Marvin L. Cohen. The existence of a superconducting state in semiconductors. Reviews of Modern Physics, 36(1):240– 243, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.240;   http://
rmp.aps.org/abstract/RMP/v36/i1/p240_1.

Cohen:1964:III

[Coh64b] Morrel H. Cohen. Influence of impurities on the interactions responsible for superconductivity. Reviews of Modern Physics, 36 (1):243–246, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.243;   http://
rmp.aps.org/abstract/RMP/v36/i1/p243_1.

Cohen:1968:SR

[Coh68] Morrel H. Cohen. Summary and review. Reviews of Modern Physics, 40(4):839–844, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.839;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p839_1.

Coleman:1963:SFD

[Col63] A. J. Coleman. Structure of fermion density matrices. Reviews of Modern Physics, 35(3):668–686, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.668;   http://rmp.aps.org/abstract/RMP/

v35/i3/p668_1. See discussion [CLS63].

Coles:1964:ESS

[Col64] B. R. Coles. Electronic structure and superconductivity of transition metals and their alloys. Reviews of Modern Physics, 36 (1):139–145, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.139;   http://
rmp.aps.org/abstract/RMP/v36/i1/p139_1.

Condon:1968:PCT

[Con68a] E. U. Condon. On pair correlation in the theory of atomic structure. Reviews of Modern Physics, 40(4):872–875, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.872;   http://rmp.aps.org/abstract/RMP/
v40/i4/p872_1.

Condon:1968:PFR

[Con68b] E. U. Condon. The past and the future of the Reviews of Modern Physics. Reviews of Modern Physics, 40(4):876– 878, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.876;   http://
rmp.aps.org/abstract/RMP/v40/i4/p876_1.

Cooper:1967:BIL

[Coo67] J. Cooper. Broadening of isolated lines in the impact approximation using a density matrix formulation. Reviews of Modern Physics, 39(1):167–177, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.167;   http:
//rmp.aps.org/abstract/RMP/v39/i1/p167_1.

Coulson:1960:PSM

[Cou60] C. A. Coulson. Present state of molecular structure calculations. Reviews of Modern Physics, 32(2):170–177, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.170;   http://rmp.aps.org/abstract/RMP/
v32/i2/p170_1.

Caron:1968:CME

[CP68] Laurent G. Caron and George W. Pratt, Jr. Correlation and magnetic effects in narrow energy bands. II. Reviews of Modern Physics, 40(4):802–806, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.802;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p802_1.

Conville:1964:TMP

[CS64] T. Mc Conville and B. Serin. Thermal and magnetic properties of second kind superconductors. III. Specific heat of Nb in a magnetic field. Reviews of Modern Physics, 36(1): 112–116, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.112;   http://
rmp.aps.org/abstract/RMP/v36/i1/p112_1.

Csonka:1965:PPE

[Cso65] Paul L. Csonka. Possible polarization experiments in fermionfermion scattering as a test of invariance principles. Reviews of Modern Physics, 37(1):177–200, January 1965. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.177;   http://rmp.aps.org/abstract/RMP/
v37/i1/p177_1.

Crayton:1962:CHB

[CSR^{+}62] N. Crayton, R. Levi Setti, M. Raymund, O. Skjeggestad, D. Abeledo, R. G. Ammar, J. H. Roberts, and E. N. Shipley. Compilation of hyperfragment binding energies. Reviews of Modern Physics, 34(2):186–189, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.186;   http:
//rmp.aps.org/abstract/RMP/v34/i2/p186_1.

Cunningham:1964:SPL

[Cun64] John Cunningham. Some properties of the Landau curves. Reviews of Modern Physics, 36(3):833–843, July 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.833;   http://rmp.aps.org/abstract/RMP/
v36/i3/p833_1.

Cutkosky:1961:AT

[Cut61] R. E. Cutkosky. Anomalous thresholds. Reviews of Modern Physics, 33(3):448–455, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.448;   http:
//rmp.aps.org/abstract/RMP/v33/i3/p448_1.

Dalitz:1961:SIS

[Dal61] R. H. Dalitz. On the strong interactions of the strange particles. Reviews of Modern Physics, 33(3):471–492, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.471;   http://rmp.aps.org/abstract/RMP/
v33/i3/p471_1.

Dalgarno:1963:IF

[Dal63a] A. Dalgarno. Intermolecular forces. Reviews of Modern Physics, 35(3):611–616, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.611;   http://
rmp.aps.org/abstract/RMP/v35/i3/p611_1.

Dalgarno:1963:SRA

[Dal63b] A. Dalgarno. Sum rules and atomic structure. Reviews of Modern Physics, 35(3):522–524, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.522;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p522_1.

Dalgarno:1967:AAC

[Dal67a] A. Dalgarno. Atom-atom collision processes in astrophysics: Theoretical studies. Reviews of Modern Physics, 39(4):850– 857, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.850;   http://
rmp.aps.org/abstract/RMP/v39/i4/p850_1.

Dalgarno:1967:SPP

[Dal67b] A. Dalgarno. Some problems in planetary atmospheres involving collision processes. Reviews of Modern Physics, 39 (4):858–861, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.858;   http://
rmp.aps.org/abstract/RMP/v39/i4/p858_1.

Daniel:1968:SBR

[Dan68] H. Daniel. Shapes of beta-ray spectra. Reviews of Modern Physics, 40(3):659–672, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.659;   http:
//rmp.aps.org/abstract/RMP/v40/i3/p659_1.

Davidson:1965:RVD

[Dav65] J. P. Davidson. Rotations and vibrations in deformed nuclei. Reviews of Modern Physics, 37(1):105–158, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.37.105;   http://rmp.aps.org/abstract/RMP/
v37/i1/p105_1.

Day:1967:EBG

[Day67] B. D. Day. Elements of the Brueckner–Goldstone theory of nuclear matter. Reviews of Modern Physics, 39(4):719– 744, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.719;   http://
rmp.aps.org/abstract/RMP/v39/i4/p719_1.

DeHeer:1963:MDO

[De 63] J. De Heer. The method of different orbitals for different spins and its application to alternant hydrocarbons. Reviews of Modern Physics, 35(3):631–637, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.631;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p631_1.

DeGennes:1964:BES

[De 64] P. G. De Gennes. Boundary effects in superconductors. Reviews of Modern Physics, 36(1):225–237, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.225;   http://rmp.aps.org/abstract/RMP/
v36/i1/p225_1.

Dedrick:1962:KHE

[Ded62] K. G. Dedrick. Kinematics of high-energy particles. Reviews of Modern Physics, 34(3):429–442, July 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.429;   http://rmp.aps.org/abstract/RMP/

v34/i3/p429_1. See errata [Ded63].

Dedrick:1963:EKH

[Ded63] K. G. Dedrick. Errata: Kinematics of High-Energy Particles. Reviews of Modern Physics, 35(2):414, April 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.414;   http://rmp.aps.org/abstract/RMP/
v35/i2/p414_1. See [Ded62].

Dehnhard:1965:DTB

[Deh65] D. Dehnhard. Direct three-body decay ^{12}C → 3α in the reaction ^{11}B(p, α)2α at the 163-keV resonance. Reviews of Modern Physics, 37(3):450–452, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.450;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p450_1.

Durkan:1968:LEI

[DEM68] J. Durkan, R. J. Elliott, and N. H. March. Localization of electrons in impure semiconductors by a magnetic field. Reviews of Modern Physics, 40(4):812–815, October 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.812;   http://rmp.aps.org/abstract/RMP/
v40/i4/p812_1.

Desorbo:1964:PES

[Des64] W. Desorbo. The peak effect in substitutional and interstitial solid solutions of high-field superconductors. Reviews of Modern Physics, 36(1):90–94, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.90;   http://rmp.aps.org/abstract/RMP/v36/i1/p90_1.

Dewar:1963:QTC

[Dew63] Michael J. S. Dewar. Quantum theory of conjugated systems. Reviews of Modern Physics, 35(3):586–598, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.586;   http://rmp.aps.org/abstract/RMP/
v35/i3/p586_1.

Dieter:1966:RWI

[DG66] Nannielou H. Dieter and W. Miller Goss. Recent work on the interstellar medium. Reviews of Modern Physics, 38 (2):256–297, April 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.256;   http://
rmp.aps.org/abstract/RMP/v38/i2/p256_1.

Dolder:1960:EPS

[DH60] K. Dolder and R. Hide. Experiments on the passage of a shock wave through a magnetic field. Reviews of Modern Physics, 32 (4):770–779, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.770;   http://
rmp.aps.org/abstract/RMP/v32/i4/p770_1.

Dicke:1962:ICS

[Dic62] R. H. Dicke. Implications for cosmology of stellar and galactic evolution rates. Reviews of Modern Physics, 34(1):110– 122, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.110;   http://
rmp.aps.org/abstract/RMP/v34/i1/p110_1.

Dirac:1962:CQF

[Dir62] P. A. M. Dirac. The conditions for a quantum field theory to be relativistic. Reviews of Modern Physics, 34(4): 592–596, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.592;   http://
rmp.aps.org/abstract/RMP/v34/i4/p592_1.

Dresden:1962:FTP

[DK62] M. Dresden and Peter B. Kahn. Field theories with persistent one particle states. I. General formalism. Reviews of Modern Physics, 34(3):401–429, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.401;   http:
//rmp.aps.org/abstract/RMP/v34/i3/p401_1.

Dubeck:1964:MTP

[DLLR64] L. Dubeck, P. Lindenfeld, E. A. Lynton, and H. Rohrer. Magnetic and thermal properties of second-kind superconductors. II. Thermal conductivity in the mixed state. Reviews of Modern Physics, 36(1):110–112, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.110;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p110_1.

DeGennes:1964:CMV

[DM64] P. G. De Gennes and J. Matricon. Collective modes of vortex lines in superconductors of the second kind. Reviews of Modern Physics, 36(1):45–49, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.45;   http://rmp.aps.org/abstract/RMP/v36/i1/p45_1.

Dombey:1969:SPL

[Dom69] Norman Dombey. Scattering of polarized leptons at high energy. Reviews of Modern Physics, 41(1):236–246, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.41.236;   http://rmp.aps.org/abstract/RMP/
v41/i1/p236_1.

Donovan:1965:EIS

[Don65] Paul F. Donovan. Experimental investigations of several few-nucleon systems. Reviews of Modern Physics, 37(3): 501–511, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.501;   http://
rmp.aps.org/abstract/RMP/v37/i3/p501_1.

Douglass:1964:STF

[Dou64] D. H. Douglass, Jr. Superconducting thin films. Reviews of Modern Physics, 36(1):316, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.316;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p316_1.

DelRe:1963:TIE

[DP63] Giuseppe Del Re and Robert G. Parr. Toward an improved π-electron theory. Reviews of Modern Physics, 35 (3):604–610, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.604;   http://
rmp.aps.org/abstract/RMP/v35/i3/p604_1.

Dobbs:1964:AEG

[DP64] E. Roland Dobbs and John M. Perz. Anisotropy of the energy gap in niobium from ultrasonic measurements. Reviews of Modern Physics, 36(1):257–260, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.257;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p257_1.

Delves:1969:PSN

[DP69] L. M. Delves and A. C. Phillips. Present status of the nuclear three-body problem. Reviews of Modern Physics, 41 (3):497–530, July 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.497;   http://
rmp.aps.org/abstract/RMP/v41/i3/p497_1.

Dyson:1964:JRO

[DPSY64] Freeman Dyson, Abraham Pais, Bengt Stromgren, and Chen Ning Yang. To J. Robert Oppenheimer on his sixtieth birthday. Reviews of Modern Physics, 36(2):507–508, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.507;   http://rmp.aps.org/abstract/RMP/
v36/i2/p507_1.

Drell:1961:PCH

[Dre61a] S. D. Drell. Peripheral contributions to high-energy interaction processes. Reviews of Modern Physics, 33(3):458–466, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.458;   http://rmp.aps.org/abstract/RMP/
v33/i3/p458_1.

Dresden:1961:RDQ

[Dre61b] M. Dresden. Recent developments in the quantum theory of transport and galvanomagnetic phenomena. Reviews of Modern Physics, 33(2):265–342, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.265;   http:
//rmp.aps.org/abstract/RMP/v33/i2/p265_1.

deSwart:1963:OMC

[dS63] J. J. de Swart. The octet model and its Clebsch–Gordan coefficients. Reviews of Modern Physics, 35(4):916–939, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.35.916;   http://rmp.aps.org/abstract/RMP/

v35/i4/p916_1. See erratum [dS65].

deSwart:1965:EOM

[dS65] J. J. de Swart. Erratum: The Octet Model and Its Clebsch–Gordan Coefficients. Reviews of Modern Physics, 37 (2):326, April 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.326;   http://
rmp.aps.org/abstract/RMP/v37/i2/p326_1. See [dS63].

Daybell:1968:LMI

[DS68] M. D. Daybell and W. A. Steyert. Localized magnetic impurity states in metals: Some experimental relationships. Reviews of Modern Physics, 40(2):380–389, April 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.380;   http://rmp.aps.org/abstract/RMP/
v40/i2/p380_1.

Devins:1965:SAL

[DSF65] D. W. Devins, Bruce L. Scott, and H. H. Forster. A simple analysis of the Li^{6}(p, pd)He^{4} reaction. Reviews of Modern Physics, 37(3):396–398, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.396;   http://
rmp.aps.org/abstract/RMP/v37/i3/p396_1.

Duck:1965:ADT

[Duc65] Ian Duck. Angular distributions in three-alpha and three-nucleon systems. Reviews of Modern Physics, 37(3):418–429, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.418;   http://rmp.aps.org/abstract/RMP/
v37/i3/p418_1.

Ducros:1967:PCM

[Duc67] Y. Ducros. Polarization and C_{NN} measurements between 0.5 and 1.2 GeV. Reviews of Modern Physics, 39(3):531–533, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.531;   http://rmp.aps.org/abstract/RMP/
v39/i3/p531_1.

Druyvesteyn:1964:VCF

[DVB64] W. F. Druyvesteyn, D. J. Van Ooijen, and T. J. Berben. Variation of the critical fields of superconducting lead with the residual resistivity. Reviews of Modern Physics, 36(1):58–61, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.58;   http://rmp.aps.org/abstract/
RMP/v36/i1/p58_1.

Dresner:1962:RDT

[DW62] Lawrence Dresner and Alvin M. Weinberg. Recent developments in the theory and technology of chain reactors. Reviews of Modern Physics, 34(4):747–765, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.747;   http:
//rmp.aps.org/abstract/RMP/v34/i4/p747_1.

Duke:1964:AMC

[DW64] C. B. Duke and E. P. Wigner. Approximation method in collision theory based on R-matrix theory. Reviews of Modern Physics, 36(2):584–589, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.584;   http://
rmp.aps.org/abstract/RMP/v36/i2/p584_1.

Eisler:1961:PRE

[EFG^{+}61] F. Eisler, P. Franzini, J. M. Gaillard, A. Garfinkel, J. Keren, R. Plano, A. Prodell, and M. Schwartz. Progress report on an experiment to study Λ^{0}–K^{0} production at sigma-K threshold. Reviews of Modern Physics, 33(3):436–438, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.436;   http://rmp.aps.org/abstract/RMP/
v33/i3/p436_1.

Esten:1965:IPD

[EGLM65] M. J. Esten, T. C. Griffith, G. J. Lush, and A. J. Metheringham. Inelastic proton-deuteron scattering at 135 MeV. Reviews of Modern Physics, 37(3):533–536, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.533;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p533_1.

Edwards:1968:CHT

[EH68] D. M. Edwards and A. C. Hewson. Comment on Hubbard’s theory of the Mott Transition. Reviews of Modern Physics, 40 (4):810–811, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.810;   http://
rmp.aps.org/abstract/RMP/v40/i4/p810_1.

Ernst:1969:TTC

[EHD69] M. H. Ernst, L. K. Haines, and J. R. Dorfman. Theory of transport coefficients for moderately dense gases. Reviews of Modern Physics, 41(2):296–316, April 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.41.296;   http:
//rmp.aps.org/abstract/RMP/v41/i2/p296_1.

Eichler:1964:SSP

[Eic64] Eugene Eichler. A survey of some properties of even-even nuclei. Reviews of Modern Physics, 36(3):809–814, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.809;   http://rmp.aps.org/abstract/RMP/
v36/i3/p809_1.

Eisberg:1964:TDM

[Eis64] R. M. Eisberg. Time delay measurements. Reviews of Modern Physics, 36(4):1100–1102, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.36.1100;
http://rmp.aps.org/abstract/RMP/v36/i4/p1100_1.

ElBindari:1964:CCF

[EL64] A. El Bindari and M. M. Litvak. Critical current and field in nonideal superconductors. Reviews of Modern Physics, 36 (1):98–102, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.98;   http://
rmp.aps.org/abstract/RMP/v36/i1/p98_1.

Esterling:1968:DMO

[EL68] D. M. Esterling and Robert V. Lange. Degenerate mass operator perturbation theory in the Hubbard model. Reviews of Modern Physics, 40(4):796–799, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.796;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p796_1.

Elliott:1964:REM

[Ell64] R. J. Elliott. Rare earth magnetism in relation to the Mössbauer effect. Reviews of Modern Physics, 36(1):385–391, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.385;   http://rmp.aps.org/abstract/RMP/
v36/i1/p385_1.

Eisberg:1961:SAP

[EP61] R. M. Eisberg and C. E. Porter. Scattering of alpha particles. Reviews of Modern Physics, 33(2):190–230, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.190;   http://rmp.aps.org/abstract/RMP/
v33/i2/p190_1.

Edmiston:1963:LAM

[ER63] Clyde Edmiston and Klaus Ruedenberg. Localized atomic and molecular orbitals. Reviews of Modern Physics, 35(3): 457–464, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.457;   http://
rmp.aps.org/abstract/RMP/v35/i3/p457_1.

Erber:1966:HEE

[Erb66] Thomas Erber. High-energy electromagnetic conversion processes in intense magnetic fields. Reviews of Modern Physics, 38(4):626–659, October 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.626;   http://
rmp.aps.org/abstract/RMP/v38/i4/p626_1.

Ewald:1965:COV

[Ewa65] P. P. Ewald. Crystal optics for visible light and X rays. Reviews of Modern Physics, 37(1):46–56, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
37.46;   http://rmp.aps.org/abstract/RMP/v37/i1/p46_1.

Etter:1965:SSB

[EWM^{+}65] J. E. Etter, M. A. Waggoner, C. Moazed, H. D. Holmgren, and C. Han. Spectroscopic studies with the B^{10}(He^{3}, pααα) reaction. Reviews of Modern Physics, 37(3):444–447, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.444;   http://rmp.aps.org/abstract/RMP/
v37/i3/p444_1.

Eyring:1962:TCR

[Eyr62] Henry Eyring. The transmission coefficient in reaction rate theory. Reviews of Modern Physics, 34(4):616–619, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.616;   http://rmp.aps.org/abstract/RMP/
v34/i4/p616_1.

Foster:1960:CCI

[FB60a] J. M. Foster and S. F. Boys. Canonical configurational interaction procedure. Reviews of Modern Physics, 32(2): 300–302, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.300;   http://
rmp.aps.org/abstract/RMP/v32/i2/p300_1.

Foster:1960:QVCa

[FB60b] J. M. Foster and S. F. Boys. A quantum variational calculation for HCHO. Reviews of Modern Physics, 32(2):303–304, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.303;   http://rmp.aps.org/abstract/RMP/
v32/i2/p303_1.

Foster:1960:QVCb

[FB60c] J. M. Foster and S. F. Boys. Quantum variational calculations for a range of CH_{2} configurations. Reviews of Modern Physics, 32(2):305–307, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.305;   http://
rmp.aps.org/abstract/RMP/v32/i2/p305_1.

Fano:1968:SDA

[FC68] U. Fano and J. W. Cooper. Spectral distribution of atomic oscillator strengths. Reviews of Modern Physics, 40(3):441– 507, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.441;   http://
rmp.aps.org/abstract/RMP/v40/i3/p441_1. See addendum

[FC69].

Fano:1969:ASD

[FC69] U. Fano and J. W. Cooper. Addendum: Spectral Distribution of Atomic Oscillator Strengths. Reviews of Modern Physics, 41 (4):724–725, October 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.724;   http://
rmp.aps.org/abstract/RMP/v41/i4/p724_1. See [FC68].

Ferraro:1960:TSC

[Fer60] V. C. A. Ferraro. Theory of sudden commencements and of the first phase of a magnetic storm. Reviews of Modern Physics, 32(4):934–940, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.934;   http://
rmp.aps.org/abstract/RMP/v32/i4/p934_1.

Feshbach:1964:UTN

[Fes64] Herman Feshbach. Unified theory of nuclear reactions. Reviews of Modern Physics, 36(4):1076–1078, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.1076;   http://rmp.aps.org/abstract/RMP/v36/i4/p1076_
1.

Fradkin:1961:EPO

[FG61] D. M. Fradkin and R. H. Good, Jr. Electron polarization operators. Reviews of Modern Physics, 33(2):343–352, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.343;   http://rmp.aps.org/abstract/RMP/
v33/i2/p343_1.

Finkelstein:1964:EIY

[Fin64] R. J. Finkelstein. Electromagnetic interactions of a Yang–Mills field. Reviews of Modern Physics, 36(2):632–640, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.632;   http://rmp.aps.org/abstract/RMP/
v36/i2/p632_1.

Fiske:1964:TMF

[Fis64] Milan D. Fiske. Temperature and magnetic field dependences of the Josephson tunneling current. Reviews of Modern Physics, 36(1):221–222, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.221;   http://
rmp.aps.org/abstract/RMP/v36/i1/p221_1.

Fink:1966:AFY

[FJMS66] R. W. Fink, R. C. Jopson, Hans Mark, and C. D. Swift. Atomic fluorescence yields. Reviews of Modern Physics, 38 (3):513–540, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.513;   http://
rmp.aps.org/abstract/RMP/v38/i3/p513_1.

Frost:1960:LEM

[FKC60] Arthur A. Frost, Reid E. Kellogg, and Earl C. Curtis. Localenergy method in electronic energy calculations. Reviews of Modern Physics, 32(2):313–317, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.313;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p313_1.

Fishman:1960:MSW

[FKP60] F. J. Fishman, A. R. Kantrowitz, and H. E. Petschek. Magnetohydrodynamic shock wave in a collision-free plasma. Reviews of Modern Physics, 32(4):959–966, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.959;   http://rmp.aps.org/abstract/RMP/
v32/i4/p959_1.

Freeman:1961:PIP

[FKS61] J. R. Freeman, H. P. Kallmann, and M. Silver. Persistent internal polarization. Reviews of Modern Physics, 33 (4):553–573, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.553;   http://
rmp.aps.org/abstract/RMP/v33/i4/p553_1.

Fletcher:1960:ELC

[Fle60a] John George Fletcher. Erratum: Local Conservation Laws in Generally Covariant Theories. Reviews of Modern Physics, 32(4): 1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1033;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1033_3. See [Fle60b].

Fletcher:1960:LCL

[Fle60b] John George Fletcher. Local conservation laws in generally covariant theories. Reviews of Modern Physics, 32(1):65–87, January 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.32.65;   http://rmp.aps.org/abstract/

RMP/v32/i1/p65_1. See erratum [Fle60a].

Fraga:1960:RCE

[FM60] Serafin Fraga and Robert S. Mulliken. Role of Coulomb energy in the valence-bond theory. Reviews of Modern Physics, 32(2):254–265, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.254;   http://
rmp.aps.org/abstract/RMP/v32/i2/p254_1.

Forster:1963:DMS

[För63] Theodore Förster. Discussion on molecular spectra. Reviews of Modern Physics, 35(3):572–575, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.572;   http://rmp.aps.org/abstract/RMP/
v35/i3/p572_1.

Fowler:1964:IPT

[Fow64a] J. L. Fowler. Invited papers from the topical conference on compound nuclear states. Reviews of Modern Physics, 36(4):1065, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.1065;   http://rmp.aps.org/
abstract/RMP/v36/i4/p1065_1.

Fowler:1964:EMS

[Fow64b] William A. Fowler. Erratum: Massive Stars, Relativistic Polytropes, and Gravitational Radiation. Reviews of Modern Physics, 36(4):1104, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.36.1104;   http://rmp.
aps.org/abstract/RMP/v36/i4/p1104_1. See [Fow64c].

Fowler:1964:MSR

[Fow64c] William A. Fowler. Massive stars, relativistic polytropes, and gravitational radiation. Reviews of Modern Physics, 36 (2):545–555, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.545;   http:
//rmp.aps.org/abstract/RMP/v36/i2/p545_1. See erratum

[Fow64b].

Frieman:1960:HSS

[FR60] E. Frieman and Manuel Rotenberg. On hydromagnetic stability of stationary equilibria. Reviews of Modern Physics, 32 (4):898–902, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.898;   http://
rmp.aps.org/abstract/RMP/v32/i4/p898_1.

Freeman:1960:RER

[Fre60] A. J. Freeman. On the role of the equivalence restriction in molecular calculations. Reviews of Modern Physics, 32 (2):273–274, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.273;   http://
rmp.aps.org/abstract/RMP/v32/i2/p273_1.

Friedrichs:1960:PB

[Fri60] K. O. Friedrichs. Pinch buckling. Reviews of Modern Physics, 32(4):889–897, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.889;   http://
rmp.aps.org/abstract/RMP/v32/i4/p889_1.

Friedman:1965:IER

[Fri65] Avner Friedman. Isometric embedding of Riemannian manifolds into Euclidean spaces. Reviews of Modern Physics, 37 (1):201–203, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.201;   http://
rmp.aps.org/abstract/RMP/v37/i1/p201_2.

Froman:1960:RCM

[Frö60] Anders Fröman. Relativistic corrections in many-electron systems. Reviews of Modern Physics, 32(2):317–321, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.317;   http://rmp.aps.org/abstract/RMP/
v32/i2/p317_1.

Fronsdal:1965:EPC

[Fro65] C. Fronsdal. Elementary particles in a curved space. Reviews of Modern Physics, 37(1):221–224, January 1965. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.221;   http://rmp.aps.org/abstract/RMP/
v37/i1/p221_1.

Fulton:1962:CIP

[FRW62] T. Fulton, F. Rohrlich, and L. Witten. Conformal invariance in physics. Reviews of Modern Physics, 34(3):442–457, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.442;   http://rmp.aps.org/abstract/RMP/
v34/i3/p442_1.

Frenkiel:1960:I

[FS60] François N. Frenkiel and William R. Sears. Introduction. Reviews of Modern Physics, 32(4):695, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.695;   http://rmp.aps.org/abstract/RMP/
v32/i4/p695_1.

Fisher:1960:APS

[FT60] Michael E. Fisher and H. N. V. Temperley. Association problem in statistical mechanics — critique of the treatment of H. S. Green and R. Leipnik. Reviews of Modern Physics, 32(4):1029– 1031, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1029;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1029_1.

Fubini:1961:RPD

[Fub61] Sergio Fubini. Recent progress in the dispersion theory of pion-nucleon interactions. Reviews of Modern Physics, 33 (3):455–457, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.455;   http://
rmp.aps.org/abstract/RMP/v33/i3/p455_1.

Falk-Vairant:1961:PCS

[FVV61] P. Falk-Vairant and G. Valladas. Partial cross sections near the higher resonance. Reviews of Modern Physics, 33(3): 362–367, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.362;   http://
rmp.aps.org/abstract/RMP/v33/i3/p362_1.

Franken:1963:OHN

[FW63] P. A. Franken and J. F. Ward. Optical harmonics and nonlinear phenomena. Reviews of Modern Physics, 35(1):23–39, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.35.23;   http://rmp.aps.org/abstract/
RMP/v35/i1/p23_1.

Garrod:1966:HPI

[Gar66] Claude Garrod. Hamiltonian path-integral methods. Reviews of Modern Physics, 38(3):483–494, July 1966. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.483;   http://rmp.aps.org/abstract/RMP/
v38/i3/p483_1.

Goodfriend:1960:CES

[GBD60] P. L. Goodfriend, F. W. Birss, and A. B. F. Duncan. Calculations on the electronic structure of the normal state of formaldehyde. Reviews of Modern Physics, 32(2):307–311, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.307;   http://rmp.aps.org/abstract/RMP/
v32/i2/p307_1.

George:1964:FHC

[GBG64] P. George, J. Beetlestone, and J. S. Griffith. Ferrihemoprotein hydroxides: a correlation between magnetic and spectroscopic properties. Reviews of Modern Physics, 36(1):441–458, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.441;   http://rmp.aps.org/abstract/RMP/
v36/i1/p441_1.

Glover:1964:CCT

[GC64] Rolfe E. Glover III and Howard T. Coffey. Critical currents in thin planar films. Reviews of Modern Physics, 36 (1):299–302, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.299;   http://
rmp.aps.org/abstract/RMP/v36/i1/p299_1.

Geballe:1964:STM

[Geb64] T. H. Geballe. Superconductivity in the transition metals. Reviews of Modern Physics, 36(1):134–138, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.134;   http://rmp.aps.org/abstract/RMP/
v36/i1/p134_1.

Germain:1960:SWS

[Ger60] P. Germain. Shock waves and shock-wave structure in magneto-fluid dynamics. Reviews of Modern Physics, 32(4): 951–958, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.951;   http://
rmp.aps.org/abstract/RMP/v32/i4/p951_1.

Gerjuoy:1961:ASP

[Ger61] E. Gerjuoy. Atomic scattering processes. Reviews of Modern Physics, 33(4):544–553, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.544;   http:
//rmp.aps.org/abstract/RMP/v33/i4/p544_1.

Gerwin:1968:SIB

[Ger68] Richard A. Gerwin. Stability of the interface between two fluids in relative motion. Reviews of Modern Physics, 40 (3):652–658, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.652;   http://
rmp.aps.org/abstract/RMP/v40/i3/p652_1.

Gasiorowicz:1969:ELF

[GG69] S. Gasiorowicz and D. A. Geffen. Effective Lagrangians and field algebras with chiral symmetry. Reviews of Modern Physics, 41(3):531–573, July 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.531;   http://
rmp.aps.org/abstract/RMP/v41/i3/p531_1.

Garvey:1969:SNM

[GGJ^{+}69] G. T. Garvey, W. J. Gerace, R. L. Jaffe, I. Talmi, and I. Kelson. Set of nuclear-mass relations and a resultant mass table. Reviews of Modern Physics, 41(4):S1–S80, October 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
41.S1;   http://rmp.aps.org/abstract/RMP/v41/i4/pS1_1.

Gilbert:1963:IRC

[Gil63] T. L. Gilbert. Interpretation of the rapid convergence of correlated wave functions. Reviews of Modern Physics, 35 (3):491–494, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.491;   http://
rmp.aps.org/abstract/RMP/v35/i3/p491_1.

Goodman:1963:ESO

[GK63a] Lionel Goodman and V. G. Krishna. Erratum: Spin-Orbit Coupling in Azines. Reviews of Modern Physics, 35(3):735, July 1963.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.735;   http://rmp.aps.org/abstract/RMP/
v35/i3/p735_1. See [GK63b].

Goodman:1963:SOC

[GK63b] Lionel Goodman and V. G. Krishna. Spin-orbit coupling in azines. Reviews of Modern Physics, 35(3):541–548, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.541;   http://rmp.aps.org/abstract/RMP/

v35/i3/p541_1. See erratum [GK63a].

Green:1960:ESA

[GL60] H. S. Green and Roy Leipnik. Exact solution of the association problem by a matrix-spinor method with applications to statistical mechanics. Reviews of Modern Physics, 32(1): 129–141, January 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.129;   http://
rmp.aps.org/abstract/RMP/v32/i1/p129_1.

Goldhaber:1961:CA

[GL61] Gerson Goldhaber and Wonyong Lee. Π–Π correlations in p̄–p annihilation. Reviews of Modern Physics, 33(3):402–406, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.402;   http://rmp.aps.org/abstract/RMP/
v33/i3/p402_1.

Ginsberg:1964:FIA

[GL64] D. M. Ginsberg and J. D. Leslie. Far infrared absorption in superconducting lead alloys. Reviews of Modern Physics, 36 (1):198–199, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.198;   http://
rmp.aps.org/abstract/RMP/v36/i1/p198_1.

Glyde:1967:RPS

[Gly67] H. R. Glyde. Rate processes in solids. Reviews of Modern Physics, 39(2):373–382, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.373;   http://
rmp.aps.org/abstract/RMP/v39/i2/p373_1.

Gell-Mann:1964:VTC

[GMGL64] M. Gell-Mann, M. L. Goldberger, and F. E. Low. The vacuum trajectory in conventional field theory. Reviews of Modern Physics, 36(2):640–649, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.640;   http:
//rmp.aps.org/abstract/RMP/v36/i2/p640_1.

Green:1967:IOC

[GMW67] Alex E. S. Green, Malcolm H. Macgregor, and Richard Wilson. The N–N interaction, 1967 — an overview and conference summary. Reviews of Modern Physics, 39(3):498–508, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.498;   http://rmp.aps.org/abstract/RMP/
v39/i3/p498_1.

Griffy:1965:SHH

[GO65] T. A. Griffy and R. J. Oakes. Structure of He^{3} and H^{3} from high-energy electron scattering. Reviews of Modern Physics, 37(3):402–405, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.402;   http://
rmp.aps.org/abstract/RMP/v37/i3/p402_1.

Golden:1960:STE

[Gol60] Sidney Golden. Statistical theory of electronic energies. Reviews of Modern Physics, 32(2):322–327, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.322;   http://rmp.aps.org/abstract/RMP/
v32/i2/p322_1.

Goldhammer:1963:SLN

[Gol63] Paul Goldhammer. The structure of light nuclei. Reviews of Modern Physics, 35(1):40–107, January 1963. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
35.40;   http://rmp.aps.org/abstract/RMP/v35/i1/p40_1.

Goldstein:1964:ABG

[Gol64] Y. Goldstein. Anomalous band gap in superconducting Nb_{3}Sn. Reviews of Modern Physics, 36(1):213–215, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.213;   http://rmp.aps.org/abstract/RMP/
v36/i1/p213_1.

Gombas:1963:PSS

[Gom63] P. Gombás. Present state of the statistical theory of atoms. Reviews of Modern Physics, 35(3):512–516, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.512;   http://rmp.aps.org/abstract/RMP/
v35/i3/p512_1.

Goodman:1964:TIL

[Goo64] B. B. Goodman. Type II or London superconductors. Reviews of Modern Physics, 36(1):12–19, January 1964. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.12;   http://rmp.aps.org/abstract/RMP/v36/i1/p12_1.

Goodman:1965:IN

[Goo65] Charles D. Goodman. Introductory note. Reviews of Modern Physics, 37(3):327–329, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.327;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p327_1.

Gorter:1964:SRS

[Gor64a] C. J. Gorter. Some remarks on superconductivity of the second kind. Reviews of Modern Physics, 36(1):27–31, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.27;   http://rmp.aps.org/abstract/
RMP/v36/i1/p27_1.

Gorter:1964:SUL

[Gor64b] C. J. Gorter. Superconductivity until 1940 in Leiden and as seen from there. Reviews of Modern Physics, 36(1):3–7, January 1964.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.3;   http://rmp.aps.org/abstract/RMP/v36/
i1/p3_1.

Graben:1964:TVC

[GP64] H. W. Graben and R. D. Present. Third virial coefficient for the Sutherland (∞, ν) potential. Reviews of Modern Physics, 36(4): 1025–1033, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.36.1025;   http://rmp.
aps.org/abstract/RMP/v36/i4/p1025_1.

Goldemberg:1966:MES

[GP66] J. Goldemberg and R. H. Pratt. Magnetic electron scattering from nuclei. Reviews of Modern Physics, 38(2):311–329, April 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.311;   http://rmp.aps.org/abstract/RMP/
v38/i2/p311_1.

Grad:1960:RPM

[Gra60]  Harold Grad.  Reducible problems in magneto-fluid dy-

namic steady flows. Reviews of Modern Physics, 32(4):830– 847, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.830;   http://
rmp.aps.org/abstract/RMP/v32/i4/p830_1.

Green:1963:DES

[Gre63] L. C. Green. Discussion on electronic structure of atoms. Reviews of Modern Physics, 35(3):488–490, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.488;   http://rmp.aps.org/abstract/RMP/
v35/i3/p488_1.

Griffin:1968:BLS

[Gri68] Allan Griffin. Brillouin light scattering from crystals in the hydrodynamic region. Reviews of Modern Physics, 40(1): 167–205, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.167;   http:

//rmp.aps.org/abstract/RMP/v40/i1/p167_1. See erratum

[Gri69].

Griffin:1969:EBS

[Gri69] A. Griffin. Erratum: Brillouin Scattering from Crystals in the Hydrodynamic Region. Reviews of Modern Physics, 41 (1):274, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.274;   http://
rmp.aps.org/abstract/RMP/v41/i1/p274_1. See [Gri68].

Gross:1965:SIS

[Gro65] Robert A. Gross. Strong ionizing shock waves. Reviews of Modern Physics, 37(4):724–743, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.724;   http:
//rmp.aps.org/abstract/RMP/v37/i4/p724_1.

Green:1967:MTI

[GS67] A. E. S. Green and T. Sawada. Meson theoretic N–N interactions for nuclear physics. Reviews of Modern Physics, 39(3):594–610, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.594;   http://
rmp.aps.org/abstract/RMP/v39/i3/p594_1.

Gudder:1968:HVQ

[Gud68] Stanley P. Gudder. Hidden variables in quantum mechanics reconsidered. Reviews of Modern Physics, 40(1):229– 231, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.229;   http://
rmp.aps.org/abstract/RMP/v40/i1/p229_1.

Grotch:1969:EPM

[GY69] H. Grotch and D. R. Yennie. Effective potential model for calculating nuclear corrections to the energy levels of hydrogen. Reviews of Modern Physics, 41(2):350–374, April 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.41.350;   http://rmp.aps.org/abstract/RMP/
v41/i2/p350_1.

Hake:1964:TPH

[Hak64] R. R. Hake. Thermodynamic properties of a high-field superconductor. Reviews of Modern Physics, 36(1):124–127, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.124;   http://rmp.aps.org/abstract/RMP/
v36/i1/p124_1.

Hameka:1962:TMP

[Ham62] H. F. Hameka. Theory of magnetic properties of molecules with particular emphasis on the hydrogen molecule. Reviews of Modern Physics, 34(1):87–101, January 1962. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
34.87;   http://rmp.aps.org/abstract/RMP/v34/i1/p87_1.

Hanamura:1968:LIA

[Han68] Eiichi Hanamura. Lattice instability associated with metalsemiconductor transitions. Reviews of Modern Physics, 40 (4):744–747, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.744;   http://
rmp.aps.org/abstract/RMP/v40/i4/p744_1.

Harris:1960:GD

[Har60] D. L. Harris III. General discussion. Reviews of Modern Physics, 32(4):1026–1028, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.1026;   http:/
/rmp.aps.org/abstract/RMP/v32/i4/p1026_1.

Harris:1963:GWF

[Har63] Frank E. Harris. Gaussian wave functions for polyatomic molecules. Reviews of Modern Physics, 35(3):558–569, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.558;   http://rmp.aps.org/abstract/RMP/
v35/i3/p558_1.

Harrison:1967:NMV

[Har67] E. R. Harrison. Normal modes of vibrations of the universe. Reviews of Modern Physics, 39(4):862–882, October 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.862;   http://rmp.aps.org/abstract/RMP/
v39/i4/p862_1.

Hasimoto:1960:MWV

[Has60] Hidenori Hasimoto. Magnetohydrodynamic wakes in a viscous conducting fluid. Reviews of Modern Physics, 32(4): 860–866, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.860;   http://
rmp.aps.org/abstract/RMP/v32/i4/p860_1.

Havas:1964:FDF

[Hav64] Peter Havas. Four-dimensional formulations of Newtonian mechanics and their relation to the special and the General Theory of Relativity. Reviews of Modern Physics, 36 (4):938–965, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.938;   http://
rmp.aps.org/abstract/RMP/v36/i4/p938_1.

Hawkins:1964:CC

[Haw64] David Hawkins. On chance and choice. Reviews of Modern Physics, 36(2):512–517, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.512;   http:
//rmp.aps.org/abstract/RMP/v36/i2/p512_1.

Hayward:1963:PLN

[Hay63] Evans Hayward. Photodisintegration of light nuclei. Reviews of Modern Physics, 35(2):324–331, April 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.324;   http://rmp.aps.org/abstract/RMP/
v35/i2/p324_1.

Heise:1964:ATT

[Hei64] B. H. Heise. Axial torque in trained superconducting wires in a transverse magnetic field. Reviews of Modern Physics, 36(1):64–66, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.64;   http://
rmp.aps.org/abstract/RMP/v36/i1/p64_1.

Heller:1967:ITN

[Hel67] Leon Heller. Interaction of two nucleons at low energies. I = 1. Reviews of Modern Physics, 39(3):584–590, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.584;   http://rmp.aps.org/abstract/RMP/
v39/i3/p584_1.

Herring:1962:CHL

[Her62] Conyers Herring. Critique of the Heitler–London method of calculating spin couplings at large distances. Reviews of Modern Physics, 34(4):631–645, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.631;   http:
//rmp.aps.org/abstract/RMP/v34/i4/p631_1.

Hensel:1968:MNT

[HF68] F. Hensel and E. U. Franck. Metal-nonmetal transition in dense mercury vapor. Reviews of Modern Physics, 40(4): 697–703, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.697;   http://
rmp.aps.org/abstract/RMP/v40/i4/p697_1.

Hein:1964:SNS

[HGB64] R. A. Hein, J. W. Gibson, and R. D. Blaugher. Superconductivity in the NbMo system. Reviews of Modern Physics, 36 (1):149–152, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.149;   http://
rmp.aps.org/abstract/RMP/v36/i1/p149_1.

Hijikata:1960:ELF

[Hij60] Katsunori Hijikata. Energy levels of F_{2} and F_{2}^{+}. Reviews of Modern Physics, 32(2):445–446, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.445;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p445_1.

Hirschfelder:1963:DIF

[Hir63] J. O. Hirschfelder. Discussion on intermolecular forces. Reviews of Modern Physics, 35(3):616–618, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.616;   http://rmp.aps.org/abstract/RMP/
v35/i3/p616_1.

Hitchcock:1964:PMN

[Hit64] Harley C. Hitchcock. Pseudoreversible magnetization of Nb. Reviews of Modern Physics, 36(1):61–64, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.61;   http://rmp.aps.org/abstract/RMP/v36/i1/p61_1.

Hite:1969:RDR

[Hit69] Gerald E. Hite. Recent developments of the Regge pole model. Reviews of Modern Physics, 41(4):669–723, October 1969. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.41.669;   http://rmp.aps.org/abstract/RMP/
v41/i4/p669_1.

Heims:1962:TGE

[HJ62] S. P. Heims and E. T. Jaynes. Theory of gyromagnetic effects and some related magnetic phenomena. Reviews of Modern Physics, 34(2):143–165, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.143;   http://
rmp.aps.org/abstract/RMP/v34/i2/p143_1.

Hammond:1964:NMR

[HK64] R. H. Hammond and G. M. Kelly. Nuclear magnetic resonance and relaxation in superconducting aluminum. Reviews of Modern Physics, 36(1):185–187, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.185;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p185_1.

Herdan:1960:DET

[HL60] R. Herdan and B. S. Liley. Dynamical equations and transport relationships for a thermal plasma. Reviews of Modern Physics, 32(4):731–741, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.731;   http://
rmp.aps.org/abstract/RMP/v32/i4/p731_1.

Hegstrom:1968:PCS

[HL68] Roger A. Hegstrom and William N. Lipscomb. Paramagnetism in closed-shell molecules. Reviews of Modern Physics, 40(2):354–358, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.354;   http://
rmp.aps.org/abstract/RMP/v40/i2/p354_1.

Hain:1960:HWF

[HLS60] K. Hain, R. Lüst, and A. Schlüter. Hydromagnetic waves of finite amplitude in a plasma with isotropic and nonisotropic pressure perpendicular to a magnetic field. Reviews of Modern Physics, 32(4):967–972, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.967;   http://
rmp.aps.org/abstract/RMP/v32/i4/p967_1.

Huby:1965:DWB

[HM65] R. Huby and J. R. Mines. Distorted-wave Born approximation for stripping to virtual levels. Reviews of Modern Physics, 37(3):406–408, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.406;   http://
rmp.aps.org/abstract/RMP/v37/i3/p406_1.

Hand:1963:EMF

[HMW63] L. N. Hand, D. G. Miller, and Richard Wilson. Electric and magnetic form factors of the nucleon. Reviews of Modern Physics, 35(2):335–349, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.335;   http://
rmp.aps.org/abstract/RMP/v35/i2/p335_1.

Harris:1963:DNE

[HNLD63] Frank E. Harris, Robert K. Nesbet, Per-Olov Löwdin, and Ernest R. Davidson. Discussion on natural expansions and properties of the chemical bond. Reviews of Modern Physics, 35(3):629–630, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.629;   http://
rmp.aps.org/abstract/RMP/v35/i3/p629_1.

Hochstrasser:1962:LOM

[Hoc62] Robin M. Hochstrasser. The luminescence of organic molecular crystals. Reviews of Modern Physics, 34(3):531–550, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.531;   http://rmp.aps.org/abstract/RMP/
v34/i3/p531_1.

Hoh:1962:LTP

[Hoh62] F. C. Hoh. Low-temperature plasma diffusion in a magnetic field. Reviews of Modern Physics, 34(2):267–286, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.267;   http://rmp.aps.org/abstract/RMP/
v34/i2/p267_1.

Honig:1968:NET

[Hon68] J. M. Honig. Nature of the electrical transition in Ti_{2}O_{3}. Reviews of Modern Physics, 40(4):748–751, October 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.748;   http://rmp.aps.org/abstract/RMP/
v40/i4/p748_1.

Horwitz:1968:OLC

[Hor68] Gerald Horwitz. Onset of local correlation in a narrow s-band model. Reviews of Modern Physics, 40(4):807–809, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.807;   http://rmp.aps.org/abstract/RMP/
v40/i4/p807_1.

Hoshizaki:1967:PAP

[Hos67a] Norio Hoshizaki. Phenomenological analysis of proton-proton scattering in the inelastic region. Reviews of Modern Physics, 39(3):700–705, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.700;   http://
rmp.aps.org/abstract/RMP/v39/i3/p700_1.

Hoshizaki:1967:RMI

[Hos67b] Norio Hoshizaki. Recent p–p measurements at INS, Tokyo, and p–p phase shifts at 50 MeV-preliminary results. Reviews of Modern Physics, 39(3):548–549, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.548;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p548_1.

Hide:1960:HFD

[HR60] R. Hide and P. H. Roberts. Hydromagnetic flow due to an oscillating plane. Reviews of Modern Physics, 32(4):799– 806, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.799;   http://
rmp.aps.org/abstract/RMP/v32/i4/p799_1.

Halperin:1968:PAS

[HR68] B. I. Halperin and T. M. Rice. Possible anomalies at a semimetalsemiconductor transistion. Reviews of Modern Physics, 40 (4):755–766, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.755;   http://
rmp.aps.org/abstract/RMP/v40/i4/p755_1.

Hagstrom:1963:NTE

[HS63] Stanley Hagstrom and Harrison Shull. The nature of the two-electron chemical bond. III. Natural orbitals for H_{2}. Reviews of Modern Physics, 35(3):624–629, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.624;   http://rmp.aps.org/abstract/RMP/
v35/i3/p624_1.

Hauser:1964:STF

[HT64] J. J. Hauser and H. C. Theuerer. Superconducting tantalum films. Reviews of Modern Physics, 36(1):80–83, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.80;   http://rmp.aps.org/abstract/RMP/v36/i1/p80_1.

Hubbard:1961:QMS

[Hub61] Paul S. Hubbard. Quantum-mechanical and semiclassical forms of the density operator theory of relaxation. Reviews of Modern Physics, 33(2):249–264, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.249;   http:
//rmp.aps.org/abstract/RMP/v33/i2/p249_1.

Huebner:1961:RBP

[Hue61] Walter F. Huebner. Relations between plasma physics and astrophysics. Comment about comet tails. Reviews of Modern Physics, 33(3):498, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.498;   http://
rmp.aps.org/abstract/RMP/v33/i3/p498_1.

Hurley:1960:ESB

[Hur60] A. C. Hurley. Electronic structure and binding energy of carbon monoxide. Reviews of Modern Physics, 32(2):400–411, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.400;   http://rmp.aps.org/abstract/RMP/
v32/i2/p400_1.

Hurley:1963:EAE

[Hur63] A. C. Hurley. Elimination of atomic errors from molecular calculations. Reviews of Modern Physics, 35(3):448–456, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.448;   http://rmp.aps.org/abstract/RMP/
v35/i3/p448_1.

Huber:1966:RBF

[HV66] D. L. Huber and J. H. Van Vleck. The role of Boltzmann factors in line shape. Reviews of Modern Physics, 38(1): 187–204, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.187;   http://
rmp.aps.org/abstract/RMP/v38/i1/p187_1.

Houtappel:1965:CBU

[HVW65] R. M. F. Houtappel, H. Van Dam, and E. P. Wigner. The conceptual basis and use of the geometric invariance principles. Reviews of Modern Physics, 37(4):595–632, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.595;   http://rmp.aps.org/abstract/RMP/
v37/i4/p595_1.

Hamilton:1963:DPN

[HW63] J. Hamilton and W. S. Woolcock. Determination of pionnucleon parameters and phase shifts by dispersion relations. Reviews of Modern Physics, 35(4):737–787, October 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.737;   http://rmp.aps.org/abstract/RMP/
v35/i4/p737_1.

Houk:1967:MNP

[HW67] Theodore L. Houk and Richard Wilson. Measurements of the neutron-proton and neutron-carbon total cross sections at electron-volt energies. Reviews of Modern Physics, 39 (3):546–547, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.546;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p546_1. See erratum

[HW68].

Houk:1968:EMN

[HW68] Theodore L. Houk and Richard Wilson. Erratum: Measurements of the Neutron-Proton and Neutron-Carbon Cross Sections at Electron Volt Energies. Reviews of Modern Physics, 40(3):672, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.672;   http://
rmp.aps.org/abstract/RMP/v40/i3/p672_1. See [HW67].

Hylleraas:1963:REQ

[Hyl63] Egil A. Hylleraas. Reminiscences from early quantum mechanics of two-electron atoms. Reviews of Modern Physics, 35(3):421–430, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.421;   http://
rmp.aps.org/abstract/RMP/v35/i3/p421_1.

Hyland:1968:SRE

[Hyl68] G. J. Hyland. Some remarks on electronic phase transitions and on the nature of the “metallic” state in VO_{2}. Reviews of Modern Physics, 40(4):739–743, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.739;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p739_1.

IHaya:1963:DAM

[I’H63] Y. J. I’Haya. Discussion on atomic and molecular calculations. Reviews of Modern Physics, 35(3):442–444, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.442;   http://rmp.aps.org/abstract/RMP/
v35/i3/p442_1.

Imai:1960:FCF

[Ima60] Isao Imai. On flows of conducting fluids past bodies. Reviews of Modern Physics, 32(4):992–999, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.992;   http://rmp.aps.org/abstract/RMP/
v32/i4/p992_1.

Itzykson:1966:UGR

[IN66] C. Itzykson and M. Nauenberg. Unitary groups: Representations and decompositions. Reviews of Modern Physics, 38 (1):95–120, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.38.95;   http://
rmp.aps.org/abstract/RMP/v38/i1/p95_1.

Inglis:1961:CQM

[Ing61] D. R. Inglis. Completeness of quantum mechanics and chargeconjugation correlations of theta particles. Reviews of Modern Physics, 33(1):1–7, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.1;   http://
rmp.aps.org/abstract/RMP/v33/i1/p1_1.

Inglis:1962:NMS

[Ing62] David R. Inglis. Nuclear models and surfaces. Reviews of Modern Physics, 34(2):165–172, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.165;   http:
//rmp.aps.org/abstract/RMP/v34/i2/p165_1.

Ingber:1967:RNF

[Ing67] L. Ingber. Realistic nuclear forces. Reviews of Modern Physics, 39(3):648, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.648;   http://
rmp.aps.org/abstract/RMP/v39/i3/p648_1.

Itoh:1965:DNH

[Ito65] Takashi Itoh. Derivation of nonrelativistic Hamiltonian for electrons from quantum electrodynamics. Reviews of Modern Physics, 37(1):159–165, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.159;   http://
rmp.aps.org/abstract/RMP/v37/i1/p159_1.

Jacchia:1963:VEU

[Jac63] Luigi G. Jacchia. Variations in the Earth’s upper atmosphere as revealed by satellite drag. Reviews of Modern Physics, 35 (4):973–991, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.973;   http://
rmp.aps.org/abstract/RMP/v35/i4/p973_1.

Jackson:1965:PAQ

[Jac65a] Daphne F. Jackson. Problems in the analysis of quasifree scattering. Reviews of Modern Physics, 37(3):393–395, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.393;   http://rmp.aps.org/abstract/RMP/
v37/i3/p393_1.

Jackson:1965:PPD

[Jac65b] J. D. Jackson. Peripheral production and decay correlations of resonances. Reviews of Modern Physics, 37(3):484–500, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.484;   http://rmp.aps.org/abstract/RMP/
v37/i3/p484_1.

Jaffe:1969:WAF

[Jaf69]  Arthur M. Jaffe. Whither axiomatic field theory?  Re-

views of Modern Physics, 41(4):576–580, October 1969. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.41.576;   http://rmp.aps.org/abstract/RMP/
v41/i4/p576_1.

Jastrow:1960:GET

[Jas60] Robert Jastrow. Geophysical effects of the trapped particle layer. Reviews of Modern Physics, 32(4):947–950, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.947;   http://rmp.aps.org/abstract/RMP/
v32/i4/p947_1.

Joiner:1964:MRT

[JB64] W. C. H. Joiner and R. D. Blaugher. Magnetic and resistive transitions of some Mo–Re alloys. Reviews of Modern Physics, 36(1):67–69, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.67;   http://
rmp.aps.org/abstract/RMP/v36/i1/p67_1.

Jones:1965:SER

[JBJ^{+}65] C. M. Jones, J. K. Bair, C. H. Johnson, H. B. Willard, and M. Reeves III. Sequential emission in the reaction ^{7}Li(d, n)2α. Reviews of Modern Physics, 37(3):437–441, July 1965. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.437;   http://rmp.aps.org/abstract/RMP/
v37/i3/p437_1.

Jerome:1968:ENP

[Jér68] D. Jérome. Effects on the NMR properties of the metal-nonmetal transistion in impure semiconductors. Reviews of Modern Physics, 40(4):830–832, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.830;   http://
rmp.aps.org/abstract/RMP/v40/i4/p830_1.

Jackson:1962:ESE

[JF62] H. W. Jackson and E. Feenberg. Energy spectrum of elementary excitations in helium II. Reviews of Modern Physics, 34 (4):686–693, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.686;   http://
rmp.aps.org/abstract/RMP/v34/i4/p686_1.

Jones:1964:UCF

[JHC64] C. K. Jones, J. K. Hulm, and B. S. Chandrasekhar. Upper critical field of solid solution alloys of the transition elements. Reviews of Modern Physics, 36(1):74–76, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.74;   http://rmp.aps.org/abstract/RMP/v36/i1/p74_1.

Jacob:1966:QFS

[JM66] Gerhard Jacob and Th. A. J. Maris. Quasi-free scattering and nuclear structure. Reviews of Modern Physics, 38(1): 121–142, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.121;   http://
rmp.aps.org/abstract/RMP/v38/i1/p121_1.

James:1963:DMM

[JN63] H. James and R. K. Nesbet. Discussion on molecules and molecular integrals. Reviews of Modern Physics, 35(3):550– 551, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.550;   http://
rmp.aps.org/abstract/RMP/v35/i3/p550_1.

Jordan:1962:GCD

[Jor62] P. Jordan. Geophysical consequences of Dirac’s hypothesis. Reviews of Modern Physics, 34(4):596–600, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.596;   http://rmp.aps.org/abstract/RMP/
v34/i4/p596_1.

Josephson:1964:CS

[Jos64a] B. D. Josephson. Coupled superconductors. Reviews of Modern Physics, 36(1):216–220, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.216;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p216_1.

Jost:1964:PBU

[Jos64b] Res Jost. Poisson brackets (an unpedagogical lecture). Reviews of Modern Physics, 36(2):572–579, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.572;   http://rmp.aps.org/abstract/RMP/
v36/i2/p572_1.

Joseph:1965:GC

[Jos65] D. W. Joseph. Generalized covariance. Reviews of Modern Physics, 37(1):225–226, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.225;   http:
//rmp.aps.org/abstract/RMP/v37/i1/p225_1.

Joseph:1967:SAL

[Jos67] A. Joseph. Self-adjoint ladder operators (I). Reviews of Modern Physics, 39(4):829–837, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.829;   http:
//rmp.aps.org/abstract/RMP/v39/i4/p829_1.

Joseph:1968:SAL

[Jos68] A. Joseph. Self-adjoint ladder operators. III. Reviews of Modern Physics, 40(4):845–871, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.845;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p845_1.

Jauch:1968:HVR

[JP68] J. M. Jauch and C. Piron. Hidden variables revisited. Reviews of Modern Physics, 40(1):228–229, January 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.228;   http://rmp.aps.org/abstract/RMP/

v40/i1/p228_1. See [BB66b, BB68].

Jordan:1961:LGD

[JS61] Thomas F. Jordan and E. C. G. Sudarshan. Lie group dynamical formalism and the relation between quantum mechanics and classical mechanics. Reviews of Modern Physics, 33 (4):515–524, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.515;   http://
rmp.aps.org/abstract/RMP/v33/i4/p515_1. See errata [JS62].

Jordan:1962:ELG

[JS62] Thomas F. Jordan and E. C. G. Sudarshan. Errata: Lie-Group Dynamical Formalism and the Relation between Quantum Mechanics and Classical Mechanics. Reviews of Modern Physics, 34(2):377, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.377;   http://
rmp.aps.org/abstract/RMP/v34/i2/p377_1. See [JS61].

Jungclaus:1960:TDB

[Jun60] Günther Jungclaus. Two-dimensional boundary layers and jets in magneto-fluid dynamics. Reviews of Modern Physics, 32 (4):823–827, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.823;   http://
rmp.aps.org/abstract/RMP/v32/i4/p823_1.

Kacser:1965:AED

[KA65] Claude Kacser and I. J. R. Aitchison. Apparent, but energydependent pseudo-resonances due to rescattering in sequential processes. Reviews of Modern Physics, 37(3):350–353, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.350;   http://rmp.aps.org/abstract/RMP/
v37/i3/p350_1.

Kaplan:1960:TPD

[Kap60] S. A. Kaplan. On the theory of propagation and decay of hydromagnetic waves in an anisotropic medium. Reviews of Modern Physics, 32(4):881, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.881;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p881_1.

Karplus:1960:WIM

[Kar60] Martin Karplus. Weak interactions in molecular quantum mechanics. Reviews of Modern Physics, 32(2):455–460, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.455;   http://rmp.aps.org/abstract/RMP/
v32/i2/p455_1.

Kazarinov:1967:EIN

[Kaz67a] Yu. M. Kazarinov. Experimental investigations of the nucleonnucleon interaction at 600–650 MeV. Reviews of Modern Physics, 39(3):509–512, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.509;   http://
rmp.aps.org/abstract/RMP/v39/i3/p509_1.

Kazarinov:1967:PSAb

[Kaz67b] Yu. M. Kazarinov. Phase-shift analysis of data and the determination of the scattering amplitude in the nucleon energy region above pion-production threshold. Reviews of Modern Physics, 39(3):706–709, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.706;   http://
rmp.aps.org/abstract/RMP/v39/i3/p706_1.

Kane:1967:EPO

[KB67a] P. P. Kane and G. Basavaraju. Erratum: Possibilities of Observation of Nonlinear Quantum Electrodynamic Effects in Vacuum. Reviews of Modern Physics, 39(4):942, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.942;   http://rmp.aps.org/abstract/RMP/
v39/i4/p942_1. See [KB67b].

Kane:1967:PON

[KB67b] P. P. Kane and G. Basavaraju. Possibilities of observation of nonlinear quantum electrodynamic effects in vacuum. Reviews of Modern Physics, 39(1):52–58, January 1967. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
39.52;   http://rmp.aps.org/abstract/RMP/v39/i1/p52_1.

See erratum [KB67a].

Kemeny:1968:SCP

[KC68] G. Kemeny and Laurent G. Caron. Self-consistent pair correlations in narrow energy bands. Reviews of Modern Physics, 40 (4):790–795, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.790;   http://
rmp.aps.org/abstract/RMP/v40/i4/p790_1.

Kieffer:1966:EII

[KD66] L. J. Kieffer and Gordon H. Dunn. Electron impact ionization cross-section data for atoms, atomic ions, and diatomic molecules: I. Experimental data. Reviews of Modern Physics, 38(1):1–35, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.38.1;   http://
rmp.aps.org/abstract/RMP/v38/i1/p1_1.

Kerth:1961:KNI

[Ker61] Leroy T. Kerth. K-nucleon interaction. Reviews of Modern Physics, 33(3):389–392, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.389;   http:
//rmp.aps.org/abstract/RMP/v33/i3/p389_1.

Kessler:1969:ESP

[Kes69] J. Kessler. Electron spin polarization by low-energy scattering from unpolarized targets. Reviews of Modern Physics, 41(1):3–25, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.41.3;   http://
rmp.aps.org/abstract/RMP/v41/i1/p3_1.

Kadanoff:1967:SPN

[KGH^{+}67] Leo P. Kadanoff, Wolfgang Götze, David Hamblen, Robert Hecht, E. A. S. Lewis, V. V. Palciauskas, Martin Rayl, J. Swift, David Aspnes, and Joseph Kane. Static phenomena near critical points: Theory and experiment. Reviews of Modern Physics, 39(2):395–431, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.395;   http://
rmp.aps.org/abstract/RMP/v39/i2/p395_1.

Kim:1964:RSH

[KHS64] Y. B. Kim, C. F. Hempstead, and A. R. Strnad. Resistive states of hard superconductors. Reviews of Modern Physics, 36(1):43–45, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.43;   http://
rmp.aps.org/abstract/RMP/v36/i1/p43_1.

Katsura:1960:GLM

[KI60] Shigetoshi Katsura and Sakari Inawashiro. On Green and Leipnik’s method for solution of the association problem. Reviews of Modern Physics, 32(4):1031–1032, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
32.1031;   http://rmp.aps.org/abstract/RMP/v32/i4/p1031_
1.

Kienle:1964:RDR

[Kie64] P. Kienle. Recent developments in rare-earth Mössbauer studies. I. Reviews of Modern Physics, 36(1):372–378, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.372;   http://rmp.aps.org/abstract/RMP/
v36/i1/p372_1.

Kim:1965:SWT

[KK65] Y. E. Kim and J. V. Kane. Some work on the theoretical interpretation of the reaction D + P → P + P + N. Reviews of Modern Physics, 37(3):519–522, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.519;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p519_1.

Kamke:1965:MRB

[KKR65] D. Kamke, J. Krug, and F.-W. Richter. Mechanism of the reaction ^{11}B(p, α)2α below and in the vicinity of the 675-keV resonance. Reviews of Modern Physics, 37(3):453–454, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.453;   http://rmp.aps.org/abstract/RMP/
v37/i3/p453_1.

Kulikovskii:1960:GIM

[KL60] A. G. Kulikovskii and G. A. Lyubimov. On gas-ionizing magnetohydrodynamic shock waves. Reviews of Modern Physics, 32 (4):977–979, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.977;   http://
rmp.aps.org/abstract/RMP/v32/i4/p977_1.

Klein:1962:EPP

[Kle62] Claude A. Klein. Electrical properties of pyrolytic graphites. Reviews of Modern Physics, 34(1):56–79, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
34.56;   http://rmp.aps.org/abstract/RMP/v34/i1/p56_1.

Klein:1968:MCP

[Kle68] J. J. Klein. Motion of a charged particle in a uniform magnetic field. Reviews of Modern Physics, 40(3):523–530, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.523;   http://rmp.aps.org/abstract/RMP/
v40/i3/p523_1.

Kazarinov:1967:PSAa

[KLJ67] Yu. M. Kazarinov, F. Lehar, and Z. Janout. Phase-shift analysis of data on nucleon-nucleon scattering at low energies. Reviews of Modern Physics, 39(3):571–573, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.571;   http://rmp.aps.org/abstract/RMP/
v39/i3/p571_1.

Kinsel:1964:MTP

[KLS64] T. Kinsel, E. A. Lynton, and B. Serin. Magnetic and thermal properties of second-kind superconductors. I. Magnetization curves. Reviews of Modern Physics, 36(1):105–109, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.105;   http://rmp.aps.org/abstract/RMP/
v36/i1/p105_1.

Kihara:1960:IPP

[KM60] Taro Kihara and Yukio Midzuno. Irreversible processes in plasmas in a strong magnetic field. Reviews of Modern Physics, 32 (4):722–730, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.722;   http://
rmp.aps.org/abstract/RMP/v32/i4/p722_1.

Kotani:1960:SPT

[KMKI60] Masao Kotani, Yukio Mizuno, Kunihusa Kayama, and Eiichi Ishiguro. Some problems in the theory of homonuclear diatomic molecules. Reviews of Modern Physics, 32(2):266–271, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.266;   http://rmp.aps.org/abstract/RMP/
v32/i2/p266_1.

Kolb:1960:MCP

[Kol60] Alan C. Kolb. Magnetic compression of plasmas. Reviews of Modern Physics, 32(4):748–757, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.748;   http://rmp.aps.org/abstract/RMP/
v32/i4/p748_1.

Kopp:1967:MNC

[Köp67] G. Köpp. Meson-nucleon coupling constants from low-energy proton-proton scattering. Reviews of Modern Physics, 39 (3):640–647, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.640;   http://
rmp.aps.org/abstract/RMP/v39/i3/p640_1.

Kotani:1963:ESI

[Kot63] Masao Kotani. On the electronic state of iron in hemoglobins, with a short introduction to problems of quantum biophysics. Reviews of Modern Physics, 35(3):717–720, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.717;   http://rmp.aps.org/abstract/RMP/
v35/i3/p717_1.

Kovasznay:1960:PT

[Kov60] Leslie S. G. Kovásznay. Plasma turbulence. Reviews of Modern Physics, 32(4):815–822, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.815;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p815_1.

Khuri:1964:SPP

[KP64a] N. N. Khuri and A. Pais. Singular potentials and Peratization. I. Reviews of Modern Physics, 36(2):590–595, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.590;   http://rmp.aps.org/abstract/RMP/
v36/i2/p590_1.

Koide:1964:IBM

[KP64b] S. Koide and M. Peter. Interaction between magnetic ions and conduction electrons. Reviews of Modern Physics, 36(1): 160–162, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.160;   http://
rmp.aps.org/abstract/RMP/v36/i1/p160_1.

Krishnaji:1966:EMQ

[KP66]  Krishnaji and Vinod Prakash.  Evaluation of molecular

quadrupole moments. Reviews of Modern Physics, 38(4): 690–709, October 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.690;   http://
rmp.aps.org/abstract/RMP/v38/i4/p690_1.

Kolos:1960:AEW

[KR60a] W. Kolos and C. C. J. Roothaan. Accurate electronic wave functions for the H_{2} molecule. Reviews of Modern Physics, 32(2):219–232, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.219;   http://
rmp.aps.org/abstract/RMP/v32/i2/p219_1.

Kolos:1960:COG

[KR60b] W. Kolos and C. C. J. Roothaan. Correlated orbitals for the ground state of the hydrogen molecule. Reviews of Modern Physics, 32(2):205–210, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.205;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p205_1.

Kramer:1965:AMP

[Kra65] Peter Kramer. Angular momenta and Pauli principle in nuclear reactions a + A → C^{∗} → b_{1} + b_{2} + b_{3}. Reviews of Modern Physics, 37(3):346–349, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.346;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p346_1.

Kolos:1960:GSS

[KRS60] W. Kolos, C. C. J. Roothaan, and R. A. Sack. Ground state of systems of three particles with Coulomb interaction. Reviews of Modern Physics, 32(2):178–179, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.178;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p178_1.

Kisslinger:1963:SNS

[KS63a] Leonard S. Kisslinger and Raymond A. Sorensen. Spherical nuclei with simple residual forces. Reviews of Modern Physics, 35 (4):853–915, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.853;   http:
//rmp.aps.org/abstract/RMP/v35/i4/p853_1. See erratum

[KS64].

Knight:1963:TEA

[KS63b] Robert E. Knight and Charles W. Scherr. Two-electron atoms II. A perturbation study of some excited states. Reviews of Modern Physics, 35(3):431–435, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.431;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p431_1.

Kisslinger:1964:ESN

[KS64] Leonard S. Kisslinger and Raymond A. Sorensen. Erratum: Spherical Nuclei with Simple Residual Forces. Reviews of Modern Physics, 36(4):1104, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.36.1104;
http://rmp.aps.org/abstract/RMP/v36/i4/p1104_2.  See

[KS63a].

Klemens:1964:RLT

[KT64] P. G. Klemens and L. Tewordt. Reduction of the lattice thermal conductivity of superconductors due to point defects. Reviews of Modern Physics, 36(1):118–120, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.118;   http://rmp.aps.org/abstract/RMP/
v36/i1/p118_1.

Kleinman:1964:PAT

[KTB64] Leonard Kleinman, B. N. Taylor, and E. Burstein. Phononassisted tunneling in superconductors. Reviews of Modern Physics, 36(1):208–210, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.208;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p208_1.

Kunzler:1961:SHM

[Kun61] J. E. Kunzler. Superconductivity in high magnetic fields at high current densities. Reviews of Modern Physics, 33(4): 501–509, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.501;   http://
rmp.aps.org/abstract/RMP/v33/i4/p501_1.

Kolos:1963:NTD

[KW63] W. Kolos and L. Wolniewicz. Nonadiabatic theory for diatomic molecules and its application to the hydrogen molecule. Reviews of Modern Physics, 35(3):473–483, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.473;   http://rmp.aps.org/abstract/RMP/
v35/i3/p473_1.

Kasuya:1968:ATP

[KY68] T. Kasuya and A. Yanase. Anomalous transport phenomena in Eu-chalcogenide alloys. Reviews of Modern Physics, 40 (4):684–696, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.684;   http://
rmp.aps.org/abstract/RMP/v40/i4/p684_1.

Liu:1968:BTS

[LA68] T. L. Liu and H. Amar. Band-theoretic study of the metallic character of the alkali-noble metal alloys. Reviews of Modern Physics, 40(4):782–789, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.782;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p782_1.

Lane:1960:RWI

[Lan60] A. M. Lane. Reduced widths of individual nuclear energy levels. Reviews of Modern Physics, 32(3):519–566, July 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.519;   http://rmp.aps.org/abstract/RMP/
v32/i3/p519_1.

Lanczos:1962:SRT

[Lan62] C. Lanczos. The splitting of the Riemann tensor. Reviews of Modern Physics, 34(3):379–389, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//adsabs.harvard.edu/abs/1962RvMP...34..379L;   http:
//link.aps.org/doi/10.1103/RevModPhys.34.379;   http:
//rmp.aps.org/abstract/RMP/v34/i3/p379_1.

Lapidus:1967:INE

[Lap67] L. I. Lapidus. Ideas on new experiments on polarized target. Reviews of Modern Physics, 39(3):689–694, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.689;   http://rmp.aps.org/abstract/RMP/
v39/i3/p689_1.

Lax:1960:FNS

[Lax60] Melvin Lax. Fluctuations from the nonequilibrium steady state. Reviews of Modern Physics, 32(1):25–64, January 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
32.25;   http://rmp.aps.org/abstract/RMP/v32/i1/p25_1.

Lax:1966:CNIa

[Lax66a] Melvin Lax. Classical noise III: Nonlinear Markoff processes. Reviews of Modern Physics, 38(2):359–379, April 1966. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.359;   http://rmp.aps.org/abstract/RMP/
v38/i2/p359_1.

Lax:1966:CNIb

[Lax66b] Melvin Lax. Classical noise IV: Langevin methods. Reviews of Modern Physics, 38(3):541–566, July 1966. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.541;   http://rmp.aps.org/abstract/RMP/
v38/i3/p541_1.

London:1964:STF

[LC64] H. London and G. R. Clarke. Superconductivity of thin films of niobium. Reviews of Modern Physics, 36(1):320– 323, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.320;   http://
rmp.aps.org/abstract/RMP/v36/i1/p320_1.

Lerner:1968:EBP

[LCW68] Lawrence S. Lerner, Kermit F. Cuff, and Lloyd R. Williams. Energy-band parameters and relative band-edge motions in the Bi–Sb alloy system near the semimetal-semiconductor transition. Reviews of Modern Physics, 40(4):770–775, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.770;   http://rmp.aps.org/abstract/RMP/
v40/i4/p770_1.

Leader:1966:PPS

[Lea66] E. Leader. Present phenomenological status of the Regge pole theory. Reviews of Modern Physics, 38(3):476–482, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.476;   http://rmp.aps.org/abstract/RMP/
v38/i3/p476_1.

Leader:1967:NNI

[Lea67] E. Leader. Nucleon-nucleon interaction at high energies. Reviews of Modern Physics, 39(3):663–669, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.663;   http://rmp.aps.org/abstract/RMP/
v39/i3/p663_1.

Lehnert:1960:EBC

[Leh60] B. Lehnert. Energy balance and confinement of a magnetized plasma. Reviews of Modern Physics, 32(4):1012–1019, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1012;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1012_1.

Lomon:1967:NNI

[LF67] E. Lomon and H. Feshbach. A nucleon-nucleon interaction consistent with theory and experiment. Reviews of Modern Physics, 39(3):611–621, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.611;   http://
rmp.aps.org/abstract/RMP/v39/i3/p611_1.

Lie:1966:TPE

[LI66] T. J. Lie and Yoshi H. Ichikawa. Thermodynamic properties of an electron plasma. Reviews of Modern Physics, 38 (4):680–689, October 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.680;   http://
rmp.aps.org/abstract/RMP/v38/i4/p680_1.

Liehr:1960:STV

[Lie60] Andrew D. Liehr. Semiempirical theory of vibronic interactions in some simple conjugated hydrocarbons. Reviews of Modern Physics, 32(2):436–439, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.436;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p436_1.

Little:1964:LRC

[Lit64] W. A. Little. Long-range correlations in superconductivity. Reviews of Modern Physics, 36(1):264–268, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.264;   http://rmp.aps.org/abstract/RMP/
v36/i1/p264_1.

Livingston:1964:DMH

[Liv64] J. D. Livingston. Defects and magnetic hysteresis in Type II superconductors. Reviews of Modern Physics, 36(1):54–58, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.54;   http://rmp.aps.org/abstract/
RMP/v36/i1/p54_1.

Lorquet:1960:IPA

[Lor60] J. C. Lorquet. Ionization potential of ammonia — some implications concerning Koopman’s theorem. Reviews of Modern Physics, 32(2):312, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.312;   http://
rmp.aps.org/abstract/RMP/v32/i2/p312_1.

Lowdin:1960:ETT

[Löw60] Per-Olov Löwdin. Expansion theorems for the total wave function and extended Hartree–Fock schemes. Reviews of Modern Physics, 32(2):328–334, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.328;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p328_1.

Lowdin:1962:ECS

[Löw62a] Per-Olov Löwdin. Exchange, correlation, and spin effects in molecular and solid-state theory. Reviews of Modern Physics, 34(1):80–87, January 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.34.80;   http://
rmp.aps.org/abstract/RMP/v34/i1/p80_1.

Lowdin:1962:NCM

[Löw62b] Per-Olov Löwdin. The normal constants of motion in quantum mechanics treated by projection technique. Reviews of Modern Physics, 34(3):520–530, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.520;   http:
//rmp.aps.org/abstract/RMP/v34/i3/p520_1.

Lowdin:1963:ISAb

[Löw63a] Per-Olov Löwdin. International Symposium on Atomic and Molecular Quantum Mechanics: Proceedings in Honor of Professor Egil A. Hylleraas and arranged by the University of Florida at Sanibel Island, 14–19 January 1963: Introduction. Reviews of Modern Physics, 35(3):415, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.415;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p415_1.

Lowdin:1963:PTD

[Löw63b] Per-Olov Löwdin. Proton tunneling in DNA and its biological implications. Reviews of Modern Physics, 35(3):724–732, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.724;   http://rmp.aps.org/abstract/RMP/
v35/i3/p724_1.

Lowdin:1963:WRO

[Löw63c] Per-Olov Löwdin. Wave and reaction operators in the quantum theory of many-particle systems. Reviews of Modern Physics, 35(3):702–707, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.702;   http://
rmp.aps.org/abstract/RMP/v35/i3/p702_1.

Lowdin:1964:AMW

[Löw64] Per-Olov Löwdin. Angular momentum wavefunctions constructed by projection operators. Reviews of Modern Physics, 36(4): 966–976, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.966;   http://
rmp.aps.org/abstract/RMP/v36/i4/p966_1.

Lowdin:1967:GAC

[Löw67] Per-Olov Löwdin. Group algebra, convolution algebra, and applications to quantum mechanics. Reviews of Modern Physics, 39(2):259–287, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

falcon.kvac.uu.se/swe/personal/PerOlovLowdinPub.html;

 http://link.aps.org/doi/10.1103/RevModPhys.39.259;
http://rmp.aps.org/abstract/RMP/v39/i2/p259_1.

Lassila:1967:PIN

[LP67] K. E. Lassila and E. I. Peltola. Pion instability in nucleon-nucleon scattering. Reviews of Modern Physics, 39(3):591–592, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.591;   http://rmp.aps.org/abstract/RMP/
v39/i3/p591_1.

Lykos:1963:DHF

[LPLH63] P. Lykos, G. W. Pratt, P.-O. Löwdin, and Gilda M. Harris. Discussion on “The Hartree–Fock Approximation”. Reviews of Modern Physics, 35(3):496–501, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.496;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p496_1.

Lehnert:1960:SHM

[LS60] B. Lehnert and G. Sjögren. Stability of a hollow mercury jet. Reviews of Modern Physics, 32(4):813–814, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.813;   http://rmp.aps.org/abstract/RMP/
v32/i4/p813_1.

Lee:1964:PMD

[LS64a] T. D. Lee and A. Sirlin. Possible method of determining the moment of charge of ν_{e}. Reviews of Modern Physics, 36 (2):666–669, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.666;   http://
rmp.aps.org/abstract/RMP/v36/i2/p666_1.

Love:1964:UAS

[LS64b] R. E. Love and R. W. Shaw. Ultrasonic attenuation in superconducting lead. Reviews of Modern Physics, 36(1):260– 264, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.260;   http://
rmp.aps.org/abstract/RMP/v36/i1/p260_1.

Ludford:1960:IFP

[Lud60] G. S. S. Ludford. Inviscid flow past a body at low magnetic Reynolds number. Reviews of Modern Physics, 32(4):1000–1003, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1000;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1000_1.

Lust:1960:MFD

[Lüs60] R. Lüst. Magneto-fluid dynamic shock waves. Reviews of Modern Physics, 32(4):706–709, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.706;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p706_1.

Lykoudis:1960:TLT

[Lyk60] Paul S. Lykoudis. Transition from laminar to turbulent flow in magneto-fluid mechanic channels. Reviews of Modern Physics, 32(4):796–798, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.796;   http://
rmp.aps.org/abstract/RMP/v32/i4/p796_1.

Lynch:1961:FSA

[Lyn61] Gerald R. Lynch. Final states of the antiproton-proton system. Reviews of Modern Physics, 33(3):395–401, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.395;   http://rmp.aps.org/abstract/RMP/
v33/i3/p395_1.

Lyth:1965:FVD

[Lyt65] D. H. Lyth. Fixed variable dispersion relations for the pion-nucleon system. Reviews of Modern Physics, 37(4):709– 723, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.709;   http://
rmp.aps.org/abstract/RMP/v37/i4/p709_1.

Macfarlane:1962:RPW

[Mac62] A. J. Macfarlane. Relativistic partial wave analysis. Reviews of Modern Physics, 34(1):41–55, January 1962. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
34.41;   http://rmp.aps.org/abstract/RMP/v34/i1/p41_1.

Macdonald:1966:SSI

[Mac66] J. Ross Macdonald. Some simple isothermal equations of state. Reviews of Modern Physics, 38(4):669–679, October 1966. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.669;   http://rmp.aps.org/abstract/RMP/
v38/i4/p669_1.

Macgregor:1967:DES

[Mac67] Malcolm H. Macgregor. Determination of the p–p elastic scattering matrix. Reviews of Modern Physics, 39(3):556–559, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.556;   http://rmp.aps.org/abstract/RMP/
v39/i3/p556_1.

Macdonald:1969:RSE

[Mac69] J. Ross Macdonald. Review of some experimental and analytical equations of state. Reviews of Modern Physics, 41 (2):316–349, April 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.316;   http://
rmp.aps.org/abstract/RMP/v41/i2/p316_1.

Manchester:1967:SLA

[Man67] F. D. Manchester. Solid layers in the adsorbed helium film and mobility of the film at low coverages. Reviews of Modern Physics, 39(2):383–394, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.383;   http://
rmp.aps.org/abstract/RMP/v39/i2/p383_1.

Marion:1961:AEC

[Mar61a] Jerry B. Marion. Accelerator energy calibrations. Reviews of Modern Physics, 33(2):139–147, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.33.139;   http:
//rmp.aps.org/abstract/RMP/v33/i2/p139_1.  See errata

[Mar61b].

Marion:1961:EAE

[Mar61b] Jerry B. Marion. Errata: Accelerator Energy Calibrations. Reviews of Modern Physics, 33(4):623, October 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.623;   http://rmp.aps.org/abstract/RMP/
v33/i4/p623_1. See [Mar61a].

Mariwalla:1962:ERD

[Mar62] K. H. Mariwalla. Explicit representation of discrete-symmetry operators in quantum theory of free fields. Reviews of Modern Physics, 34(2):215–226, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.215;   http:
//rmp.aps.org/abstract/RMP/v34/i2/p215_1.

Maradudin:1964:LDA

[Mar64a] A. A. Maradudin. Lattice dynamical aspects of the resonance absorption of gamma rays by nuclei bound in a crystal. Reviews of Modern Physics, 36(1):417–432, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.417;   http://rmp.aps.org/abstract/RMP/
v36/i1/p417_1.

Marcus:1964:ESG

[Mar64b] Paul M. Marcus. Exact solution of the Ginzburg–Landau equations for slabs in tangential magnetic fields. Reviews of Modern Physics, 36(1):294–299, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.294;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p294_1.

Marion:1966:ACE

[Mar66] Jerry B. Marion. Accelerator calibration energies. Reviews of Modern Physics, 38(4):660–668, October 1966. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.660;   http://rmp.aps.org/abstract/RMP/
v38/i4/p660_1.

Mason:1965:SZS

[Mas65] D. E. Mason. Stimulation of zinc sulfide and similar inorganic phosphors. Reviews of Modern Physics, 37(4):743– 762, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.743;   http://
rmp.aps.org/abstract/RMP/v37/i4/p743_1.

Matthias:1961:CSM

[Mat61] B. T. Matthias. Conclusions from superconducting molybdenum technetium alloys. Reviews of Modern Physics, 33(4): 499–500, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.499;   http://
rmp.aps.org/abstract/RMP/v33/i4/p499_1.

Matsen:1963:DTS

[Mat63] F. A. Matsen. Discussion on treatment of symmetry properties. Reviews of Modern Physics, 35(3):645, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.645;   http://rmp.aps.org/abstract/RMP/
v35/i3/p645_1.

Maximon:1969:CRC

[Max69] L. C. Maximon. Comments on radiative corrections. Reviews of Modern Physics, 41(1):193–204, January 1969. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.41.193;   http://rmp.aps.org/abstract/RMP/
v41/i1/p193_1.

Macdonald:1963:RRS

[MB63] J. Ross Macdonald and Carl A. Barlow, Jr. Relaxation, retardation, and superposition. Reviews of Modern Physics, 35 (4):940–946, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.940;   http://
rmp.aps.org/abstract/RMP/v35/i4/p940_1.

Mclachlan:1964:TDH

[MB64] A. D. Mclachlan and M. A. Ball. Time-dependent Hartree– Fock theory for molecules. Reviews of Modern Physics, 36 (3):844–855, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.844;   http://
rmp.aps.org/abstract/RMP/v36/i3/p844_1.

Mccarthy:1965:IOP

[Mcc65] I. E. Mccarthy. Information obtainable from (p, 2 p) reactions. Reviews of Modern Physics, 37(3):388–392, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.388;   http://rmp.aps.org/abstract/RMP/
v37/i3/p388_1.

Mccumber:1966:HFC

[Mcc66] D. E. Mccumber. High-frequency conductivity of a solid-state plasma with dynamically screened interactions. I. Thermal equilibrium. Reviews of Modern Physics, 38(3):494–512, July 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.494;   http://rmp.aps.org/abstract/RMP/
v38/i3/p494_1.

McDowell:1963:SRH

[McD63] Charles A. McDowell. Spin resonance and hyperfine interaction. Reviews of Modern Physics, 35(3):528–537, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.528;   http://rmp.aps.org/abstract/RMP/
v35/i3/p528_1.

Marshall:1964:MDA

[MCJR64] W. Marshall, T. E. Cranshaw, C. E. Johnson, and M. S. Ridout. Magnetism of dilute alloys. Reviews of Modern Physics, 36(1):399–406, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.399;   http://
rmp.aps.org/abstract/RMP/v36/i1/p399_1.

McKinley:1963:IFP

[McK63] John M. McKinley. Interpolative formulas for pion-nucleon scattering phase shifts. Reviews of Modern Physics, 35(4): 788–794, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.788;   http://
rmp.aps.org/abstract/RMP/v35/i4/p788_1.

McMaster:1961:MRP

[McM61] William H. McMaster. Matrix representation of polarization. Reviews of Modern Physics, 33(1):8–28, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
33.8;   http://rmp.aps.org/abstract/RMP/v33/i1/p8_1.

Mcvoy:1965:SGP

[Mcv65] Kirk W. Mcvoy. Symmetry groups in physics. Reviews of Modern Physics, 37(1):84–104, January 1965. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
37.84;   http://rmp.aps.org/abstract/RMP/v37/i1/p84_1.

McWeeny:1960:SRA

[McW60] R. McWeeny. Some recent advances in density matrix theory. Reviews of Modern Physics, 32(2):335–369, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.335;   http://rmp.aps.org/abstract/RMP/
v32/i2/p335_1.

McWeeny:1963:DST

[McW63] R. McWeeny. Discussion on statistical theory of atoms. Reviews of Modern Physics, 35(3):510–511, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.510;   http://rmp.aps.org/abstract/RMP/
v35/i3/p510_1.

Massot:1967:GGM

[MEBL67] J.-N. Massot, E. El-Baz, and J. Lafoucrière. A general graphical method for angular momentum. Reviews of Modern Physics, 39(2):288–305, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.288;   http://
rmp.aps.org/abstract/RMP/v39/i2/p288_1.

Moazed:1965:R

[MEHW65a] C. Moazed, J. E. Etter, H. D. Holmgren, and M. A. Waggoner. The Be^{9} (He^{3}, ααα) reaction. Reviews of Modern Physics, 37(3):354–357, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.354;   http://
rmp.aps.org/abstract/RMP/v37/i3/p354_1.

Moazed:1965:PPA

[MEHW65b] C. Moazed, J. E. Etter, H. D. Holmgren, and M. A. Waggoner. Particle-particle angular correlation for the Be^{9}(He^{3}, α)Be^{8} → αα reaction. Reviews of Modern Physics, 37(3):441–443, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.441;   http://rmp.aps.org/abstract/RMP/
v37/i3/p441_1.

Meiklejohn:1964:TDC

[Mei64] W. H. Meiklejohn. Temperature dependence of the critical current of tantalum and indium-tin alloy films. Reviews of Modern Physics, 36(1):302–305, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.302;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p302_1.

Melvin:1960:EPS

[Mel60] M. A. Melvin. Elementary particles and symmetry principles. Reviews of Modern Physics, 32(3):477–518, July 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.477;   http://rmp.aps.org/abstract/RMP/

v32/i3/p477_1. See errata [Mel62].

Melvin:1962:ENE

[Mel62] M. A. Melvin. Errata and note: Elementary particles and symmetry principles. Reviews of Modern Physics, 34(3): 582–584, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.582;   http://
rmp.aps.org/abstract/RMP/v34/i3/p582_1. See [Mel60].

Mendelssohn:1964:DTS

[Men64a] K. Mendelssohn. On different types of superconductivity. Reviews of Modern Physics, 36(1):50–51, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.50;   http://rmp.aps.org/abstract/RMP/v36/i1/p50_1.

Mendelssohn:1964:PWS

[Men64b] K. Mendelssohn. Prewar work on superconductivity as seen from Oxford. Reviews of Modern Physics, 36(1):7–12, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.7;   http://rmp.aps.org/abstract/RMP/v36/
i1/p7_1.

Menzel:1964:GRI

[Men64c] Donald H. Menzel. Generalized radial integrals with hydrogenic functions. Reviews of Modern Physics, 36(2):613–617, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.613;   http://rmp.aps.org/abstract/RMP/
v36/i2/p613_1.

Merriam:1964:BZE

[Mer64] M. F. Merriam. Brillouin zone effects in superconductivity. Reviews of Modern Physics, 36(1):152–155, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.152;   http://rmp.aps.org/abstract/RMP/
v36/i1/p152_1.

Meyer:1960:MHF

[Mey60] Rudolf X. Meyer. Magnetohydrodynamic-hypersonic flow in the quasi-Newtonian approximation. Reviews of Modern Physics, 32 (4):1004–1007, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.1004;   http:/
/rmp.aps.org/abstract/RMP/v32/i4/p1004_1.

Meyerhof:1965:CES

[Mey65] W. E. Meyerhof. Comparison of excited state of helium 4 from two different reactions. Reviews of Modern Physics, 37 (3):512–515, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.512;   http://
rmp.aps.org/abstract/RMP/v37/i3/p512_1.

Macfarlane:1960:SRS

[MF60] M. H. Macfarlane and J. B. French. Stripping reactions and the structure of light and intermediate nuclei. Reviews of Modern Physics, 32(3):567–691, July 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.567;   http:
//rmp.aps.org/abstract/RMP/v32/i3/p567_1.

Meissner:1964:SIN

[MF64a] Hans Meissner and Robert V. Fanelli. Surface impedance of normal conducting films deposited on superconducting substrates. Reviews of Modern Physics, 36(1):194–197, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.194;   http://rmp.aps.org/abstract/RMP/
v36/i1/p194_1.

Moorjani:1964:EPA

[MF64b] Kishin Moorjani and Charles Feldman. Electrical properties of amorphous semiconductors. Reviews of Modern Physics, 36(4): 1042–1049, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.36.1042;   http://rmp.
aps.org/abstract/RMP/v36/i4/p1042_1.

Macklin:1965:NCD

[MG65] R. L. Macklin and J. H. Gibbons. Neutron capture data at stellar temperatures. Reviews of Modern Physics, 37(1): 166–176, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.166;   http://
rmp.aps.org/abstract/RMP/v37/i1/p166_1.

Matthias:1963:ES

[MGC63a] B. T. Matthias, T. H. Geballe, and V. B. Compton. Errata: Superconductivity. Reviews of Modern Physics, 35 (2):414, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.414;   http://
rmp.aps.org/abstract/RMP/v35/i2/p414_2. See [MGC63b].

Matthias:1963:S

[MGC63b] B. T. Matthias, T. H. Geballe, and V. B. Compton. Superconductivity. Reviews of Modern Physics, 35(1):1–22, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.1;   http://rmp.aps.org/abstract/RMP/v35/

i1/p1_1. See errata [MGC63a].

Matthias:1964:SRI

[MGC^{+}64] B. T. Matthias, T. H. Geballe, V. B. Compton, E. Corenzwit, and G. W. Hull, Jr. Superconductivity in the Y–Rh and Y–Ir systems. Reviews of Modern Physics, 36(1):155–156, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.155;   http://rmp.aps.org/abstract/RMP/
v36/i1/p155_1.

Mcvoy:1967:OAP

[MHB67] K. W. Mcvoy, L. Heller, and M. Bolsterli. Optical analysis of potential well resonances. Reviews of Modern Physics, 39 (1):245–258, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.245;   http://
rmp.aps.org/abstract/RMP/v39/i1/p245_1.

Michalowicz:1967:SER

[Mic67] A. Michalowicz. A survey of experimental results on nucleonnucleon scattering work done by the Orsay Group. Reviews of Modern Physics, 39(3):523–525, July 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.523;   http://rmp.aps.org/abstract/RMP/
v39/i3/p523_1.

Mikoshiba:1968:MMN

[Mik68] Nobuo Mikoshiba. Model for the metal-nonmetal transition in impure semiconductors. Reviews of Modern Physics, 40 (4):833–838, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.833;   http://
rmp.aps.org/abstract/RMP/v40/i4/p833_1.

Mitescu:1964:CTS

[Mit64] Catalin D. Mitescu. Currents in thin superconducting films: Nonlinear effects. Reviews of Modern Physics, 36(1):305– 308, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.305;   http://
rmp.aps.org/abstract/RMP/v36/i1/p305_1.

Mahaffey:1968:MDC

[MJ68] Derek W. Mahaffey and David A. Jerde. Microwave dielectric constants of sodium-ammonia solutions. Reviews of Modern Physics, 40(4):710–713, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.710;   http://
rmp.aps.org/abstract/RMP/v40/i4/p710_1.

Mcnamee:1964:TCG

[MJC64] P. Mcnamee, S. J., and Frank Chilton. Tables of Clebsch–Gordan coefficients of SU_{3}. Reviews of Modern Physics, 36(4):1005–1024, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.1005;   http://rmp.aps.org/
abstract/RMP/v36/i4/p1005_1.

Miller:1964:TMS

[MKQ64] P. B. Miller, B. W. Kington, and D. J. Quinn. Transverse magnetization of In–Sn films. Reviews of Modern Physics, 36 (1):70–73, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.36.70;   http://
rmp.aps.org/abstract/RMP/v36/i1/p70_1.

Motz:1964:ESA

[MOK64] J. W. Motz, Haakon Olsen, and H. W. Koch. Electron scattering without atomic or nuclear excitation. Reviews of Modern Physics, 36(4):881–928, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.881;   http:
//rmp.aps.org/abstract/RMP/v36/i4/p881_1.

Motz:1969:PPP

[MOK69] J. W. Motz, Haakon A. Olsen, and H. W. Koch. Pair production by photons. Reviews of Modern Physics, 41(4): 581–639, October 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.581;   http://
rmp.aps.org/abstract/RMP/v41/i4/p581_1.

Moldauer:1964:ACN

[Mol64] P. A. Moldauer. Average compound-nucleus cross sections. Reviews of Modern Physics, 36(4):1079–1084, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.1079;   http://rmp.aps.org/abstract/RMP/
v36/i4/p1079_1.

Moore:1960:NRC

[Moo60] R. G. Moore, Jr. Nuclear reaction cross-section theory. Reviews of Modern Physics, 32(1):101–116, January 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.101;   http://rmp.aps.org/abstract/RMP/
v32/i1/p101_1.

Morrison:1964:TCS

[Mor64] P. Morrison. A thermodynamic characterization of selfreproduction. Reviews of Modern Physics, 36(2):517–524, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.517;   http://rmp.aps.org/abstract/
RMP/v36/i2/p517_1.

Moravcsik:1967:TNI

[Mor67] Michael J. Moravcsik. Two-nucleon interaction as a research tool in particle physics. Reviews of Modern Physics, 39 (3):670–680, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.670;   http://
rmp.aps.org/abstract/RMP/v39/i3/p670_1.

Moscowitz:1960:ARD

[Mos60] Albert Moscowitz. Analysis of rotatory dispersion curves. Reviews of Modern Physics, 32(2):440–443, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.440;   http://rmp.aps.org/abstract/RMP/
v32/i2/p440_1.

Moshinsky:1962:WCG

[Mos62] Marcos Moshinsky. Wigner coefficients for the SU_{3} group and some applications. Reviews of Modern Physics, 34(4): 813–828, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.813;   http://
rmp.aps.org/abstract/RMP/v34/i4/p813_1.

Moser:1963:DDO

[Mos63] C. Moser. Discussion on different orbitals for different spins. Reviews of Modern Physics, 35(3):637–640, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.637;   http://rmp.aps.org/abstract/RMP/
v35/i3/p637_1.

Mossbauer:1964:GAN

[Mös64] R. L. Mössbauer. General aspects of nuclear hyperfine interactions in salts of the rare earths. Reviews of Modern Physics, 36(1):362–371, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.362;   http://
rmp.aps.org/abstract/RMP/v36/i1/p362_1.

Moszkowski:1967:NSN

[Mos67] Steven A. Moszkowski. Nuclear structure and the nucleonnucleon interaction. Reviews of Modern Physics, 39(3): 657–662, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.657;   http://
rmp.aps.org/abstract/RMP/v39/i3/p657_1.

Mott:1968:MIT

[Mot68] N. F. Mott. Metal-insulator transition. Reviews of Modern Physics, 40(4):677–683, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.677;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p677_1.

Mountain:1966:SDS

[Mou66] Raymond D. Mountain. Spectral distribution of scattered light in a simple fluid. Reviews of Modern Physics, 38(1): 205–214, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.205;   http://
rmp.aps.org/abstract/RMP/v38/i1/p205_1.

Moyer:1961:ESE

[Moy61] Burton J. Moyer. Π^{−}-p elastic scattering in the energy region 500– 1500 MeV. Reviews of Modern Physics, 33(3):367–373, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.367;   http://rmp.aps.org/abstract/RMP/
v33/i3/p367_1.

Munschy:1963:SCE

[MP63] Gilbert Munschy and Philippe Pluvinage. On the spatial correlation of electrons in the ground state of helium. Reviews of Modern Physics, 35(3):494–495, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.494;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p494_1.

Marshalek:1963:SDA

[MPS63] Eugene Marshalek, Lucy Wu Person, and Raymond K. Sheline. Systematics of deformations of atomic nuclei. Reviews of Modern Physics, 35(1):108–116, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.108;   http:
//rmp.aps.org/abstract/RMP/v35/i1/p108_1.

Mossbauer:1964:TRS

[MS64a] R. L. Mössbauer and D. H. Sharp. On the theory of resonant scattering of gamma rays by nuclei bound in crystals. Reviews of Modern Physics, 36(1):410–417, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.410;   http://rmp.aps.org/abstract/RMP/
v36/i1/p410_1.

Muller:1964:IPS

[MS64b] C. B. Müller and E. J. Saur. Influence of pressure on the superconductivity of some high-field superconductors. Reviews of Modern Physics, 36(1):103–105, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.103;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p103_1.

Moiseiwitsch:1968:EIE

[MS68] B. L. Moiseiwitsch and S. J. Smith. Electron impact excitation of atoms. Reviews of Modern Physics, 40(2):238–353, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.238;   http://rmp.aps.org/abstract/RMP/

v40/i2/p238_1. See erratum [MS69].

Moiseiwitsch:1969:EEI

[MS69] B. L. Moiseiwitsch and S. J. Smith. Erratum: Electron Impact Excitation of Atoms. Reviews of Modern Physics, 41 (3):574, July 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.574;   http://
rmp.aps.org/abstract/RMP/v41/i3/p574_1. See [MS68].

Mo:1969:RCE

[MT69] L. W. Mo and Y. S. Tsai. Radiative corrections to elastic and inelastic ep and μp scattering. Reviews of Modern Physics, 41 (1):205–235, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.41.205;   http://
rmp.aps.org/abstract/RMP/v41/i1/p205_1.

Mulliken:1960:SCF

[Mul60] R. S. Mulliken. Self-consistent field atomic and molecular orbitals and their approximations as linear combinations of Slater-type orbitals. Reviews of Modern Physics, 32(2):232–238, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.232;   http://rmp.aps.org/abstract/RMP/
v32/i2/p232_1.

Musher:1967:AWF

[Mus67] Jeremy I. Musher. Approximate wave functions for molecules and crystals using perturbation theory. Reviews of Modern Physics, 39(1):203–227, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.203;   http://
rmp.aps.org/abstract/RMP/v39/i1/p203_1.

Maradudin:1968:SPN

[MV68] A. A. Maradudin and S. H. Vosko. Symmetry properties of the normal vibrations of a crystal. Reviews of Modern Physics, 40(1):1–37, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.1;   http://
rmp.aps.org/abstract/RMP/v40/i1/p1_1.

Mandel:1965:CPO

[MW65] L. Mandel and E. Wolf. Coherence properties of optical fields. Reviews of Modern Physics, 37(2):231–287, April 1965. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.231;   http://rmp.aps.org/abstract/RMP/
v37/i2/p231_1.

McLean:1960:CIH

[MWY60] A. D. McLean, A. Weiss, and M. Yoshimine. Configuration interaction in the hydrogen molecule — the ground state. Reviews of Modern Physics, 32(2):211–218, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.211;   http://rmp.aps.org/abstract/RMP/
v32/i2/p211_1.

Nakagawa:1960:SRH

[Nak60] Yoshinari Nakagawa. Some results on heat transport by convection in the presence of a magnetic field. Reviews of Modern Physics, 32(4):916–918, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.916;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p916_1.

Napolitano:1960:TMF

[Nap60] L. G. Napolitano. On turbulent magneto-fluid dynamic boundary layers. Reviews of Modern Physics, 32(4):785–795, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.785;   http://rmp.aps.org/abstract/RMP/
v32/i4/p785_1.

NeEman:1965:EST

[Ne’65] Yuval Ne’Eman. Embedded space-time and particle symmetries. Reviews of Modern Physics, 37(1):227–230, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.227;   http://rmp.aps.org/abstract/RMP/
v37/i1/p227_1.

Nesbet:1960:DMP

[Nes60] R. K. Nesbet. Diatomic molecule project at RIAS and Boston University. Reviews of Modern Physics, 32(2):272– 273, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.272;   http://
rmp.aps.org/abstract/RMP/v32/i2/p272_1.

Nesbet:1961:AMQ

[Nes61] R. K. Nesbet. Approximate methods in the quantum theory of many-fermion systems. Reviews of Modern Physics, 33 (1):28–36, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.33.28;   http://
rmp.aps.org/abstract/RMP/v33/i1/p28_1.

Nesbet:1963:CPE

[Nes63] R. K. Nesbet. Computer programs for electronic wavefunction calculations. Reviews of Modern Physics, 35(3): 552–557, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.552;   http://
rmp.aps.org/abstract/RMP/v35/i3/p552_1.

Nelson:1968:SQP

[NG68] T. J. Nelson and R. H. Good, Jr. Second-quantization process for particles with any spin and with internal symmetry.

Reviews of Modern Physics, 40(3):508–522, July 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.508;   http://rmp.aps.org/abstract/RMP/
v40/i3/p508_1.

Nieto:1968:SLT

[Nie68] Michael Martin Nieto. Study of leptonic three-body decays of baryons. Reviews of Modern Physics, 40(1):140– 152, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.140;   http://
rmp.aps.org/abstract/RMP/v40/i1/p140_1.

Nigam:1963:TNI

[Nig63] B. P. Nigam. The two-nucleon interaction. Reviews of Modern Physics, 35(1):117–129, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.117;   http:
//rmp.aps.org/abstract/RMP/v35/i1/p117_1.

Noer:1964:NMR

[NK64] R. J. Noer and W. D. Knight. Nuclear magnetic resonance and relaxation in superconducting vanadium. Reviews of Modern Physics, 36(1):177–185, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.177;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p177_1.

Norbeck:1965:LR

[NMC65] E. Norbeck, M. D. Mancusi, and R. R. Carlson. The 2 Li^{6} → 3α reaction. Reviews of Modern Physics, 37(3):455–457, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.455;   http://rmp.aps.org/abstract/RMP/
v37/i3/p455_1.

Nambu:1964:MFP

[NT64] Y. Nambu and S. F. Tuan. Magnetic field and phase transition in superconducting thin films. Reviews of Modern Physics, 36(1):288–291, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.288;   http://
rmp.aps.org/abstract/RMP/v36/i1/p288_1.

OKonski:1963:DSR

[OA63] C. T. O’Konski and L. C. Allen. Discussion on spin resonance and hyperfine splitting. Reviews of Modern Physics, 35(3):537–540, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.537;   http://
rmp.aps.org/abstract/RMP/v35/i3/p537_1.

Olsen:1964:SUP

[OBL^{+}64] J. L. Olsen, E. Bucher, M. Levy, J. Muller, E. Corenzwit, and T. Geballe. Superconductivity under pressure. Reviews of Modern Physics, 36(1):168–170, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.168;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p168_1.

OKonski:1963:DPT

[O’K63] Chester T. O’Konski. Discussion on the proton-tunneling hypothesis for dielectric relaxation in ice. Reviews of Modern Physics, 35(3):732–733, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.732;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p732_1.

Oster:1960:LTP

[Ost60] L. Oster. Linearized theory of plasma oscillations. Reviews of Modern Physics, 32(1):141–168, January 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.141;   http://rmp.aps.org/abstract/RMP/
v32/i1/p141_1.

Oster:1961:EAC

[Ost61] Ludwig Oster. Emission, absorption, and conductivity of a fully ionized gas at radio frequencies. Reviews of Modern Physics, 33(4):525–543, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.525;   http://
rmp.aps.org/abstract/RMP/v33/i4/p525_1.

Pai:1960:WMS

[Pai60] S. I. Pai. Wave motions of small amplitude in a fully ionized plasma without external magnetic field. Reviews of Modern Physics, 32(4):882–887, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.882;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p882_1.

Pais:1966:DSP

[Pai66] A. Pais. Dynamical symmetry in particle physics. Reviews of Modern Physics, 38(2):215–255, April 1966. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.38.215;   http://rmp.aps.org/abstract/RMP/
v38/i2/p215_1.

Palazzi:1968:HEB

[Pal68] Giordano Diambrini Palazzi. High-energy Bremsstrahlung and electron pair production in thin crystals. Reviews of Modern Physics, 40(3):611–631, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.611;   http:
//rmp.aps.org/abstract/RMP/v40/i3/p611_1.

Parr:1960:PCM

[Par60] Robert G. Parr. Papers from the Conference on Molecular Quantum Mechanics, Held at the University of Colorado, Boulder, Colorado, June 21–27, 1959. Introductory note. Reviews of Modern Physics, 32(2):169, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.169;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p169_1.

Park:1964:STT

[Par64] J. G. Park. Superconducting transition of tin alloys in an alternating field. Reviews of Modern Physics, 36(1):87–90, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.87;   http://rmp.aps.org/abstract/
RMP/v36/i1/p87_1.

Parkinson:1965:PGC

[PB65] E. R. Parkinson and D. Bodansky. Proton–gamma competition in the Cu^{63} (p, 2p)Ni^{62} reaction at 10.5 MeV. Reviews of Modern Physics, 37(3):372–374, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.372;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p372_1.

Powell:1968:PV

[PBS68] R. J. Powell, C. N. Berglund, and W. E. Spicer. Photoemission from VO_{2}. Reviews of Modern Physics, 40(4): 737–738, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.737;   http://
rmp.aps.org/abstract/RMP/v40/i4/p737_2.

Penrose:1965:RPP

[Pen65] Roger Penrose. A remarkable property of plane waves in General Relativity. Reviews of Modern Physics, 37(1):215– 220, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.215;   http://
rmp.aps.org/abstract/RMP/v37/i1/p215_1.

Perdrisat:1966:SSS

[Per66] Charles F. Perdrisat. Survey of some systematic properties of the nuclear E 1 transition probability. Reviews of Modern Physics, 38(1):41–94, January 1966. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
38.41;   http://rmp.aps.org/abstract/RMP/v38/i1/p41_1.

Perring:1967:PSA

[Per67] J. K. Perring. Phase-shift analysis at Harwell. Reviews of Modern Physics, 39(3):550–555, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.550;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p550_1.

Peterson:1967:FTN

[Pet67] Robert L. Peterson. Formal theory of nonlinear response. Reviews of Modern Physics, 39(1):69–77, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
39.69;   http://rmp.aps.org/abstract/RMP/v39/i1/p69_1.

Phelps:1968:RVE

[Phe68] A. V. Phelps. Rotational and vibrational excitation of molecules by low-energy electrons. Reviews of Modern Physics, 40 (2):399–410, April 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.399;   http://
rmp.aps.org/abstract/RMP/v40/i2/p399_1.

Phillips:1964:CMT

[Phi64] G. C. Phillips. The cluster model, time delay, and threebody breakup. Reviews of Modern Physics, 36(4):1085–1094, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.1085;   http://rmp.aps.org/
abstract/RMP/v36/i4/p1085_1.

Phillips:1965:SDT

[Phi65] G. C. Phillips. Sequential decay of ^{12}C into three α particles. Reviews of Modern Physics, 37(3):409–417, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.409;   http://rmp.aps.org/abstract/RMP/
v37/i3/p409_1.

Phillips:1967:AF

[Phi67] R. J. N. Phillips. Antinuclear forces. Reviews of Modern Physics, 39(3):681–688, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.681;   http://
rmp.aps.org/abstract/RMP/v39/i3/p681_1.

Pipkin:1964:SFD

[Pip64a] A. C. Pipkin. Small finite deformations of viscoelastic solids. Reviews of Modern Physics, 36(4):1034–1041, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.1034;   http://rmp.aps.org/
abstract/RMP/v36/i4/p1034_1.

Pippard:1964:UPS

[Pip64b] A. B. Pippard. Unsolved problems of superconductivity. Reviews of Modern Physics, 36(1):328–331, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.328;   http://rmp.aps.org/abstract/RMP/
v36/i1/p328_1.

Pitzer:1963:DCR

[Pit63] K. S. Pitzer. Discussion on chemical reactivity. Reviews of Modern Physics, 35(3):619–623, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.619;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p619_1.

Pohl:1963:DPB

[PK63] H. A. Pohl and W. Kauzmann. Discussion on problems of biophysics and biochemistry. Reviews of Modern Physics, 35 (3):721–723, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.721;   http://
rmp.aps.org/abstract/RMP/v35/i3/p721_1.

Plass:1961:CEE

[Pla61] Gilbert N. Plass. Classical electrodynamic equations of motion with radiative reaction. Reviews of Modern Physics, 33 (1):37–62, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.33.37;   http://
rmp.aps.org/abstract/RMP/v33/i1/p37_1.

Phillips:1964:LHC

[PLG64] Norman E. Phillips, Marcel H. Lambert, and William R. Gardner. Lattice heat capacity of superconducting mercury and lead. Reviews of Modern Physics, 36(1):131–134, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.131;   http://rmp.aps.org/abstract/RMP/
v36/i1/p131_1.

Parks:1964:DEQ

[PM64] R. D. Parks and J. M. Mochel. Direct evidence for quantized vortices in a thin superconductor. Reviews of Modern Physics, 36(1):284–288, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.284;   http://
rmp.aps.org/abstract/RMP/v36/i1/p284_1.

Polanyi:1962:TKB

[Pol62] Michael Polanyi. Tacit knowing: Its bearing on some problems of philosophy. Reviews of Modern Physics, 34(4):601– 616, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.601;   http://
rmp.aps.org/abstract/RMP/v34/i4/p601_1.

Pollack:1964:SSR

[Pol64] Gerald L. Pollack. The solid state of rare gases. Reviews of Modern Physics, 36(3):748–791, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.748;   http:
//rmp.aps.org/abstract/RMP/v36/i3/p748_1.

Pollack:1969:KR

[Pol69] Gerald L. Pollack. Kapitza resistance. Reviews of Modern Physics, 41(1):48–81, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.41.48;   http:/
/rmp.aps.org/abstract/RMP/v41/i1/p48_1.

Porter:1964:SPC

[Por64] C. E. Porter. Statistical propoerties of compound states. Reviews of Modern Physics, 36(4):1094–1100, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.1094;   http://rmp.aps.org/abstract/RMP/v36/i4/p1094_
1.

Post:1967:SE

[Pos67] E. J. Post. Sagnac effect. Reviews of Modern Physics, 39 (2):475–493, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.475;   http://
rmp.aps.org/abstract/RMP/v39/i2/p475_1.

Pullman:1960:SEA

[PP60] Bernard Pullman and Alberte Pullman. Some electronic aspects of biochemistry. Reviews of Modern Physics, 32(2): 428–436, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.428;   http://
rmp.aps.org/abstract/RMP/v32/i2/p428_1.

Paalman:1963:FAX

[PP63] H. H. Paalman and C. J. Pings. Fourier analysis of X-ray diffraction data from liquids. Reviews of Modern Physics, 35 (2):389–399, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.389;   http://
rmp.aps.org/abstract/RMP/v35/i2/p389_1.

Perlow:1964:MEX

[PP64] G. J. Perlow and M. R. Perlow. The Mössbauer effect in xenon compounds. Reviews of Modern Physics, 36(1):353– 357, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.353;   http://
rmp.aps.org/abstract/RMP/v36/i1/p353_1.

Pratt:1963:GSP

[Pra63] G. W. Pratt, Jr. A generalized single-particle equation. Reviews of Modern Physics, 35(3):502–505, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.502;   http://rmp.aps.org/abstract/RMP/
v35/i3/p502_1.

Preuss:1963:CAE

[Pre63] Heinzwerner Preuss. Calculation of adiabatic energy surfaces for molecules using the method of atomic associations. Reviews of Modern Physics, 35(3):646–667, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.646;   http://rmp.aps.org/abstract/RMP/
v35/i3/p646_1.

Primas:1963:GPT

[Pri63] H. Primas. Generalized perturbation theory in operator form. Reviews of Modern Physics, 35(3):710–711, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.710;   http://rmp.aps.org/abstract/RMP/
v35/i3/p710_1.

Rosenblum:1964:DUC

[RAG64] E. S. Rosenblum, S. H. Autler, and K. H. Gooen. The dependence of the upper critical field of niobium on temperature and resistivity. Reviews of Modern Physics, 36(1):77–80, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.77;   http://rmp.aps.org/abstract/
RMP/v36/i1/p77_1.

Ransil:1960:SMSa

[Ran60a] Bernard J. Ransil. Studies in molecular structure. I. Scope and summary of the diatomic molecule program. Reviews of Modern Physics, 32(2):239–244, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.32.239;   http:
//rmp.aps.org/abstract/RMP/v32/i2/p239_1.

Ransil:1960:SMSb

[Ran60b] Bernard J. Ransil. Studies in molecular structure. II. LCAO–MO– SCF wave functions for selected first-row diatomic molecules. Reviews of Modern Physics, 32(2):245–254, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.245;   http://rmp.aps.org/abstract/RMP/
v32/i2/p245_1.

Rank:1962:PIS

[Ran62] D. H. Rank. Precision infrared spectroscopy. Reviews of Modern Physics, 34(3):577–581, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.577;   http:
//rmp.aps.org/abstract/RMP/v34/i3/p577_1.

Rastall:1964:QR

[Ras64] Peter Rastall. Quaternions in Relativity. Reviews of Modern Physics, 36(3):820–832, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.820;   http:
//rmp.aps.org/abstract/RMP/v36/i3/p820_1.

Rajaraman:1967:TBP

[RB67a] R. Rajaraman and H. A. Bethe. Three-body problem in nuclear matter. Reviews of Modern Physics, 39(4):745–770, October 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.745;   http://rmp.aps.org/abstract/RMP/
v39/i4/p745_1.

Rose:1967:ADG

[RB67b] H. J. Rose and D. M. Brink. Angular distributions of gamma rays in terms of phase-defined reduced matrix elements. Reviews of Modern Physics, 39(2):306–347, April 1967. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.306;   http://rmp.aps.org/abstract/RMP/
v39/i2/p306_1.

Rosenfeld:1964:DEP

[RBGB^{+}64] Arthur H. Rosenfeld, Angela Barbaro-Galtieri, Walter H. Barkas, Pierre L. Bastien, Janos Kirz, and Matts Roos. Data on elementary particles and resonant states. Reviews of Modern Physics, 36(4):977–1004, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.977;   http://
rmp.aps.org/abstract/RMP/v36/i4/p977_1.

Rosenfeld:1965:DPR

[RBGB^{+}65] Arthur H. Rosenfeld, Angela Barbaro-Galtieri, Walter H. Barkas, Pierre L. Bastien, Janos Kirz, and Matts Roos. Data on particles and resonant states. Reviews of Modern Physics, 37 (4):633–651, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.633;   http://
rmp.aps.org/abstract/RMP/v37/i4/p633_1.

Rosenfeld:1967:DPR

[RBGP^{+}67] Arthur H. Rosenfeld, Angela Barbaro-Galtieri, William J. Podolsky, Leroy R. Price, Paul Soding, Charles G. Wohl, Matts Roos, and William J. Willis. Data on particles and resonant states. Reviews of Modern Physics, 39(1):1–51, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
39.1;   http://rmp.aps.org/abstract/RMP/v39/i1/p1_1.

Rosenfeld:1968:DPR

[RBSBG^{+}68] Arthur H. Rosenfeld, Naomi Barash-Schmidt, Angela Barbaro-Galtieri, Leroy R. Price, Paul Söding, Charles G. Wohl, Matts Roos, and William J. Willis. Data on particles and resonant states. Reviews of Modern Physics, 40(1):77–128, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.40.77;   http://rmp.aps.org/abstract/
RMP/v40/i1/p77_1.

Rhodes:1967:GST

[RC67] William Rhodes and Malcolm Chase. Generalized susceptibility theory I. Theories of hypochromism. Reviews of Modern Physics, 39(2):348–361, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.348;   http:
//rmp.aps.org/abstract/RMP/v39/i2/p348_1.

Rogers:1961:MPC

[RCF^{+}61] Ernest H. Rogers, Owen Chamberlain, James H. Foote, Herbert M. Steiner, Clyde Wiegand, and Tom Ypsilantis. 310 MeV Π^{+}-p polarization and cross-section experiments. Phase-shift analysis. Reviews of Modern Physics, 33(3):356–361, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.356;   http://rmp.aps.org/abstract/RMP/
v33/i3/p356_1.

Redlich:1968:FTS

[Red68] Otto Redlich. Fundamental thermodynamics since Carathéodory. Reviews of Modern Physics, 40(3):556–563, July 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.40.556;   http://rmp.aps.org/abstract/RMP/
v40/i3/p556_1.

Resler:1960:SRH

[Res60] E. L. Resler, Jr. Some remarks on hydromagnetic waves for finite conductivity. Reviews of Modern Physics, 32(4): 866–867, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.866;   http://
rmp.aps.org/abstract/RMP/v32/i4/p866_1.

Robinson:1960:RER

[RF60] Berol L. Robinson and Richard W. Fink. Recent experimental results on orbital electron capture. Reviews of Modern Physics, 32(1):117–128, January 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.117;   http://
rmp.aps.org/abstract/RMP/v32/i1/p117_1.

Rose:1968:ENT

[RHW68] H. J. Rose, O. Häusser, and E. K. Warburton. Evidence for a nuclear tensor force from mass-14 beta- and gamma-ray data. Reviews of Modern Physics, 40(3):591–610, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.591;   http://rmp.aps.org/abstract/RMP/
v40/i3/p591_1.

Richardson:1960:IMO

[Ric60] James W. Richardson. Improved molecular orbital wave functions and nuclear quadrupole coupling in CO, N_{2}, and Li_{2}. Reviews of Modern Physics, 32(2):461–465, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.461;   http://rmp.aps.org/abstract/RMP/
v32/i2/p461_1.

Riou:1965:INS

[Rio65] M. Riou. Information on nuclear strucure from (p, 2p) and other knock-out reactions. Reviews of Modern Physics, 37 (3):375–387, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.375;   http://
rmp.aps.org/abstract/RMP/v37/i3/p375_1.

Resler:1960:SES

[RM60] E. L. Resler, Jr. and J. E. McCune. Some exact solutions in linearized magnetoaerodynamics for arbitrary magnetic Reynolds numbers. Reviews of Modern Physics, 32(4): 848–854, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.848;   http://
rmp.aps.org/abstract/RMP/v32/i4/p848_1.

Robinson:1965:SEP

[RN65] I. Robinson and Y. Ne’Eman. Seminar on the embedding problem introduction. Reviews of Modern Physics, 37(1): 201, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.201;   http://
rmp.aps.org/abstract/RMP/v37/i1/p201_1.

Reid:1963:CRW

[RÖ63] Charles E. Reid and Yngve Öhrn. On the calculations of real wave functions in natural form for two-electron systems. Reviews of Modern Physics, 35(3):445–447, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.445;   http://rmp.aps.org/abstract/RMP/
v35/i3/p445_1.

Romain:1963:TMA

[Rom63] J. E. Romain. Time measurements in accelerated frames of reference. Reviews of Modern Physics, 35(2):376–388, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.376;   http://rmp.aps.org/abstract/RMP/
v35/i2/p376_1.

Roothaan:1960:SCF

[Roo60] C. C. J. Roothaan. Self-consistent field theory for open shells of electronic systems. Reviews of Modern Physics, 32 (2):179–185, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.179;   http://
rmp.aps.org/abstract/RMP/v32/i2/p179_1.

Roos:1963:TEP

[Roo63] Matts Roos. Tables of elementary particles and resonant states. Reviews of Modern Physics, 35(2):314–323, April 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.314;   http://rmp.aps.org/abstract/RMP/
v35/i2/p314_1.

Rossow:1960:FDC

[Ros60] Vernon J. Rossow. Flow in direct-current electromagnetic pumps. Reviews of Modern Physics, 32(4):987–991, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.987;   http://rmp.aps.org/abstract/RMP/
v32/i4/p987_1.

Rosen:1965:EVR

[Ros65] Joe Rosen. Embedding of various relativistic Riemannian spaces in pseudo-Euclidean spaces. Reviews of Modern Physics, 37 (1):204–214, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.204;   http://
rmp.aps.org/abstract/RMP/v37/i1/p204_1.

Rowell:1964:TBS

[Row64] J. M. Rowell. Tunneling between superconductors. Reviews of Modern Physics, 36(1):199–200, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.199;   http://rmp.aps.org/abstract/RMP/
v36/i1/p199_1.

Rowe:1968:EMM

[Row68] D. J. Rowe. Equations-of-motion method and the extended shell model. Reviews of Modern Physics, 40(1):153–166, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.153;   http://rmp.aps.org/abstract/RMP/
v40/i1/p153_1.

Roothaan:1960:ASC

[RSW60] C. C. J. Roothaan, Lester M. Sachs, and A. W. Weiss. Analytical self-consistent field functions for the atomic configurations 1s^{2}, 1s^{2}2s, and 1s^{2}2s^{2}. Reviews of Modern Physics, 32(2):186–194, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.186;   http://
rmp.aps.org/abstract/RMP/v32/i2/p186_1.

Rudge:1968:TIA

[Rud68] M. R. H. Rudge. Theory of the ionization of atoms by electron impact. Reviews of Modern Physics, 40(3):564–590, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.564;   http://rmp.aps.org/abstract/RMP/
v40/i3/p564_1.

Ruedenberg:1962:PNC

[Rue62] Klaus Ruedenberg. The physical nature of the chemical bond. Reviews of Modern Physics, 34(2):326–376, April 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.326;   http://rmp.aps.org/abstract/RMP/
v34/i2/p326_1.

Ruedenberg:1963:DCS

[Rue63] K. Ruedenberg. Discussion on conjugated systems. Reviews of Modern Physics, 35(3):599–603, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.599;   http://rmp.aps.org/abstract/RMP/
v35/i3/p599_1.

Ruelle:1964:CPC

[Rue64] David Ruelle. Cluster property of the correlation functions of classical gases. Reviews of Modern Physics, 36(2): 580–584, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.580;   http://
rmp.aps.org/abstract/RMP/v36/i2/p580_1.

Roothaan:1960:COG

[RW60] C. C. J. Roothaan and A. W. Weiss. Correlated orbitals for the ground state of heliumlike systems. Reviews of Modern Physics, 32(2):194–205, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.194;   http://
rmp.aps.org/abstract/RMP/v32/i2/p194_1.

Reif:1964:EPA

[RW64] F. Reif and Michael A. Woolf. Effect of paramagnetic atoms on the energy gap of superconductors. Reviews of Modern Physics, 36(1):238–240, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.238;   http://
rmp.aps.org/abstract/RMP/v36/i1/p238_1.

Salam:1961:SSN

[Sal61] Abdus Salam. Some speculations on the new resonances. Reviews of Modern Physics, 33(3):426–430, July 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.426;   http://rmp.aps.org/abstract/RMP/
v33/i3/p426_1.

Silverman:1967:CSA

[SB67] Jeremiah N. Silverman and George H. Brigman. A comparative study of atomic variational wave functions. Reviews of Modern Physics, 39(1):228–244, January 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.228;   http:
//rmp.aps.org/abstract/RMP/v39/i1/p228_1.

Simpson:1965:EAC

[SBJP65] W. D. Simpson, J. D. Bronson, W. R. Jackson, and G. C. Phillips. Energy-angular correlations and sequential decay in the ^{2}H(p, 2p)n and ^{1}H(d, 2p)n reactions. Reviews of Modern Physics, 37(3):523–528, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.523;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p523_1.

Schwartz:1964:DSI

[SC64] Brian B. Schwartz and Leon N. Cooper. Does superconductivity imply flux quantization? Reviews of Modern Physics, 36 (1):280–284, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.280;   http://
rmp.aps.org/abstract/RMP/v36/i1/p280_1.

Scalapino:1964:SST

[Sca64] D. J. Scalapino. Singularities in superconducting tunnel junction I–V characteristics. Reviews of Modern Physics, 36(1): 205–207, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.205;   http://
rmp.aps.org/abstract/RMP/v36/i1/p205_1.

Shiffman:1964:SHM

[SCGP64] C. A. Shiffman, J. F. Cochran, M. Garber, and G. W. Pearsall. Specific heat measurements and proximity effects in tin-lead eutectic alloys. Reviews of Modern Physics, 36 (1):127–130, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.127;   http://
rmp.aps.org/abstract/RMP/v36/i1/p127_1.

Schatzman:1960:SPF

[Sch60] E. Schatzman. On the stability of a periodic force-free field with neutral lines. Reviews of Modern Physics, 32(4):903– 906, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.903;   http://
rmp.aps.org/abstract/RMP/v32/i4/p903_1.

Schick:1961:MSA

[Sch61] L. H. Schick. Multiple-scattering analysis of the three-body scattering problem. Reviews of Modern Physics, 33(4):608– 622, October 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.608;   http://
rmp.aps.org/abstract/RMP/v33/i4/p608_1.

Schwartz:1963:STF

[Sch63] Judah L. Schwartz. Stability of Thomas–Fermi molecules. Reviews of Modern Physics, 35(3):734, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.734;   http://rmp.aps.org/abstract/RMP/
v35/i3/p734_1.

Schiff:1964:OBM

[Sch64a] L. I. Schiff. Observational basis of Mach’s Principle. Reviews of Modern Physics, 36(2):510–511, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.510;   http://rmp.aps.org/abstract/RMP/
v36/i2/p510_1.

Schiffer:1964:RWS

[Sch64b] John P. Schiffer. Reduced widths and strength functions. Reviews of Modern Physics, 36(4):1065–1070, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.1065;   http://rmp.aps.org/abstract/RMP/v36/i4/p1065_
2.

Schrieffer:1964:TET

[Sch64c] J. R. Schrieffer. Theory of electron tunneling. Reviews of Modern Physics, 36(1):200–204, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.200;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p200_1.

Schwinger:1964:NAV

[Sch64d] Julian Schwinger. Non-abelian vector gauge fields and the electromagnetic field. Reviews of Modern Physics, 36(2): 609–613, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.609;   http://
rmp.aps.org/abstract/RMP/v36/i2/p609_1.

Schnitzer:1965:DRD

[Sch65] Howard J. Schnitzer. Dispersion relations and direct nuclear reactions. Reviews of Modern Physics, 37(4):666–678, October 1965.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.666;   http://rmp.aps.org/abstract/RMP/
v37/i4/p666_1.

Schwarz:1968:KAC

[Sch68] Gerhard Schwarz. Kinetic analysis by chemical relaxation methods. Reviews of Modern Physics, 40(1):206–218, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.206;   http://rmp.aps.org/abstract/RMP/
v40/i1/p206_1.

Sciama:1964:EPS

[Sci64a] D. W. Sciama. Erratum: The Physical Structure of General Relativity. Reviews of Modern Physics, 36(4):1103, October 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.1103;   http://rmp.aps.org/
abstract/RMP/v36/i4/p1103_1. See [Sci64b].

Sciama:1964:PSG

[Sci64b] D. W. Sciama. The physical structure of General Relativity. Reviews of Modern Physics, 36(1):463–469, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.463;   http://rmp.aps.org/abstract/RMP/

v36/i1/p463_1. See erratum [Sci64a].

Scott:1962:RGR

[Sco62] G. G. Scott. Review of gyromagnetic ratio experiments. Reviews of Modern Physics, 34(1):102–109, January 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.102;   http://rmp.aps.org/abstract/RMP/
v34/i1/p102_1.

Scott:1963:TSA

[Sco63] William T. Scott. The theory of small-angle multiple scattering of fast charged particles. Reviews of Modern Physics, 35(2):231–313, April 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.231;   http://
rmp.aps.org/abstract/RMP/v35/i2/p231_1.

Saxena:1961:TDB

[SD61] S. C. Saxena and S. M. Dave. Thermal diffusion of binary gas mixtures. Reviews of Modern Physics, 33(2):148–152, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.148;   http://rmp.aps.org/abstract/RMP/
v33/i2/p148_1.

Signell:1967:NNS

[SD67] Peter Signell and John Durso. Nucleon-nucleon scattering, pionnucleon scattering, and the nucleon form factors. Reviews of Modern Physics, 39(3):635–639, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.635;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p635_1.

Seraphim:1964:FEC

[SDH64] D. P. Seraphim, F. M. D’Heurle, and W. R. Heller. Free energy of composite wires in the superconducting state. Reviews of Modern Physics, 36(1):323–327, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.323;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p323_1.

Sears:1960:SRA

[Sea60] W. R. Sears. Some remarks about flow past bodies. Reviews of Modern Physics, 32(4):701–705, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.701;   http://rmp.aps.org/abstract/RMP/
v32/i4/p701_1.

Seitz:1962:EIM

[Sei62] Frederick Seitz. The effects of irradiation on metals. Reviews of Modern Physics, 34(4):656–666, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.656;   http://rmp.aps.org/abstract/RMP/
v34/i4/p656_1.

Selove:1965:SAF

[Sel65] W. Selove. Some aspects of final-state interaction studies at high energies. Reviews of Modern Physics, 37(3):460–472, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.460;   http://rmp.aps.org/abstract/RMP/
v37/i3/p460_1.

Serber:1964:HEP

[Ser64] Robert Serber. High-energy proton-proton scattering. Reviews of Modern Physics, 36(2):649–655, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.649;   http://rmp.aps.org/abstract/RMP/
v36/i2/p649_1.

Sewell:1964:CIF

[Sew64] G. L. Sewell. Concept of an “inner field” in the theory of superconductivity. Reviews of Modern Physics, 36(1):247– 251, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.247;   http://
rmp.aps.org/abstract/RMP/v36/i1/p247_1.

Silsbee:1964:OAM

[SF64] R. H. Silsbee and D. B. Fitchen. Optical analogs of the Mössbauer effect in solids. Reviews of Modern Physics, 36 (1):432–436, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.432;   http://
rmp.aps.org/abstract/RMP/v36/i1/p432_1.

Saxena:1963:TCM

[SG63] S. C. Saxena and J. M. Gandhi. Thermal conductivity of multicomponent mixtures of inert gases. Reviews of Modern Physics, 35(4):1022–1032, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.1022;   http:/
/rmp.aps.org/abstract/RMP/v35/i4/p1022_1.

Sheline:1960:VSD

[She60a] Raymond K. Sheline. Vibrational states in deformed even-even nuclei. Reviews of Modern Physics, 32(1):1–24, January 1960.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.1;   http://rmp.aps.org/abstract/RMP/v32/
i1/p1_1.

Shercliff:1960:SGO

[She60b] J. A. Shercliff. Some generalizations in one-dimensional magneto-gas dynamics. Reviews of Modern Physics, 32(4): 980–986, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.980;   http://
rmp.aps.org/abstract/RMP/v32/i4/p980_1.

Sheldon:1963:ACI

[She63] Eric Sheldon. Angular correlation in inelastic nucleon scattering. Reviews of Modern Physics, 35(4):795–852, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.795;   http://rmp.aps.org/abstract/RMP/
v35/i4/p795_1.

Shirley:1964:AII

[Shi64] D. A. Shirley. Application and interpretation of isomer shifts. Reviews of Modern Physics, 36(1):339–351, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.339;   http://rmp.aps.org/abstract/RMP/
v36/i1/p339_1.

Shore:1967:STA

[Sho67] Bruce W. Shore. Scattering theory of absorption-line profiles and refractivity. Reviews of Modern Physics, 39(2): 439–462, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.39.439;   http://
rmp.aps.org/abstract/RMP/v39/i2/p439_1.

Silvert:1964:SDP

[Sil64] William Silvert. Spatial dependence of pair correlation functions in nonhomogeneous superconductors. Reviews of Modern Physics, 36(1):251–253, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.251;   http://
rmp.aps.org/abstract/RMP/v36/i1/p251_1.

Simmons:1967:LEN

[Sim67] J. E. Simmons. Low-energy neutron-proton polarization experiments. Reviews of Modern Physics, 39(3):542–545, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.542;   http://rmp.aps.org/abstract/RMP/
v39/i3/p542_1.

Sinanoglu:1963:DPT

[Sin63a] O. Sinanoglu. Discussion on perturbation theory and reaction operators. Reviews of Modern Physics, 35(3):708–709, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.708;   http://rmp.aps.org/abstract/RMP/
v35/i3/p708_1.

Sinanoglu:1963:MAM

[Sin63b] Oktay Sinanoğlu. A method for the analysis of manyelectron wave functions. Reviews of Modern Physics, 35(3): 517–519, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.517;   http://
rmp.aps.org/abstract/RMP/v35/i3/p517_1.

Shapiro:1964:EMJ

[SJH64] Sidney Shapiro, Andre R. Janus, and Sandor Holly. Effect of microwaves on Josephson currents in superconducting tunneling. Reviews of Modern Physics, 36(1):223–225, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.223;   http://rmp.aps.org/abstract/RMP/
v36/i1/p223_1.

Stonehill:1962:PPI

[SK62] D. L. Stonehill and H. L. Kraybill. Pion-pion interactions in π^{+}–p collisions. Reviews of Modern Physics, 34(3):503–507, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.503;   http://rmp.aps.org/abstract/RMP/
v34/i3/p503_1.

Scherr:1963:TEA

[SK63] Charles W. Scherr and Robert E. Knight. Two-electron atoms III. A sixth-order perturbation study of the 1^{1}S ground state. Reviews of Modern Physics, 35(3):436–442, July 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.436;   http://rmp.aps.org/abstract/RMP/
v35/i3/p436_1.

Sherrington:1968:SAG

[SK68a] D. Sherrington and W. Kohn. Speculations about grey tin. Reviews of Modern Physics, 40(4):767–769, October 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.767;   http://rmp.aps.org/abstract/RMP/
v40/i4/p767_1.

Sparks:1968:MST

[SK68b] Joseph T. Sparks and Ted Komoto. Metal-to-semiconductor transition in hexagonal NiS. Reviews of Modern Physics, 40 (4):752–754, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.752;   http://
rmp.aps.org/abstract/RMP/v40/i4/p752_1.

Slater:1963:DAS

[Sla63a] J. C. Slater. Discussion on atomic spectra. Reviews of Modern Physics, 35(3):520–521, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.520;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p520_1.

Slater:1963:ESA

[Sla63b] J. C. Slater. The electronic structure of atoms-the Hartree– Fock method and correlation. Reviews of Modern Physics, 35(3):484–487, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.484;   http://
rmp.aps.org/abstract/RMP/v35/i3/p484_1.

Slater:1965:SGW

[Sla65] J. C. Slater. Space groups and wave-function symmetry in crystals. Reviews of Modern Physics, 37(1):68–83, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.37.68;   http://rmp.aps.org/abstract/
RMP/v37/i1/p68_1.

Slaus:1967:NNI

[Sla67] Ivo Slaus. Neutron-neutron interaction. Reviews of Modern Physics, 39(3):575–583, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.575;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p575_1.

Slepian:1960:CRH

[Sle60] Joseph Slepian. Comment on “Review of High-Temperature Rotating-Plasma Experiments”. Reviews of Modern Physics, 32 (4):1032–1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.32.1032;   http:
//rmp.aps.org/abstract/RMP/v32/i4/p1032_1. See [Wil59,

Wil60].

Steele:1962:CSE

[SLV62] Derek Steele, Ellis R. Lippincott, and Joseph T. Vanderslice. Comparative study of empirical internuclear potential functions. Reviews of Modern Physics, 34(2):239–251, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.239;   http://rmp.aps.org/abstract/RMP/
v34/i2/p239_1.

Saxena:1965:TDB

[SM65] S. C. Saxena and B. P. Mathur. Thermal diffusion in binary gas mixtures and intermolecular forces. Reviews of Modern Physics, 37(2):316–325, April 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.316;   http://
rmp.aps.org/abstract/RMP/v37/i2/p316_1.

Saxena:1966:TDI

[SM66] S. C. Saxena and B. P. Mathur. Thermal diffusion in isotopic gas mixtures and intermolecular forces. Reviews of Modern Physics, 38(2):380–390, April 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.380;   http://
rmp.aps.org/abstract/RMP/v38/i2/p380_1.

Smith:1964:SSS

[Smi64] Cyril Stanley Smith. Structure, substructure, and superstructure. Reviews of Modern Physics, 36(2):524–532, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.524;   http://rmp.aps.org/abstract/RMP/
v36/i2/p524_1.

Schultz:1964:TDI

[SML64] T. D. Schultz, D. C. Mattis, and E. H. Lieb. Two-dimensional Ising model as a soluble problem of many fermions. Reviews of Modern Physics, 36(3):856–871, July 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.856;   http://rmp.aps.org/abstract/RMP/
v36/i3/p856_1.

Strongin:1964:ASM

[SMR64] Myron Strongin, E. Maxwell, and T. B. Reed. Ac susceptibility measurements on transition metal superconductors containing rare earth and ferromagnetic metal solutes. Reviews of Modern Physics, 36(1):164–168, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.36.164;   http:
//rmp.aps.org/abstract/RMP/v36/i1/p164_1.

Stern:1969:LEE

[SPB69] R. M. Stern, J. J. Perry, and D. S. Boudreaux. Low-energy electron-diffraction dispersion surfaces and band structure in three-dimensional mixed Laue and Bragg reflections. Reviews of Modern Physics, 41(2):275–295, April 1969. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.41.275;   http://rmp.aps.org/abstract/RMP/
v41/i2/p275_1.

Spitzer:1960:PMF

[Spi60] Lyman Spitzer, Jr. Problems in magneto-fluid dynamics. Reviews of Modern Physics, 32(4):696–700, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.696;   http://rmp.aps.org/abstract/RMP/
v32/i4/p696_1.

Salant:1961:PPI

[SPRM61] E. O. Salant, E. Pickup, D. K. Robinson, and B. A. Munir. Pion-pion interactions and pion production branching ratios at the third resonance. Reviews of Modern Physics, 33 (3):435–436, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.435;   http://
rmp.aps.org/abstract/RMP/v33/i3/p435_1.

Saxena:1962:TPT

[SR62] S. C. Saxena and S. Raman. Theory and performance of thermal-diffusion column. Reviews of Modern Physics, 34 (2):252–266, April 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.252;   http://
rmp.aps.org/abstract/RMP/v34/i2/p252_1.

Sherrill:1964:MNT

[SR64a] M. D. Sherrill and K. Rose. Microwave nonlinearities in thin superconducting films. Reviews of Modern Physics, 36 (1):312–315, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.312;   http://
rmp.aps.org/abstract/RMP/v36/i1/p312_1.

Silcox:1964:HHS

[SR64b] J. Silcox and R. W. Rollins. Hysteresis in hard superconductors. Reviews of Modern Physics, 36(1):52–54, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
36.52;   http://rmp.aps.org/abstract/RMP/v36/i1/p52_1.

Smith:1964:FQP

[SR64c] T. I. Smith and H. E. Rorschach, Jr. Flux quantization and persistent currents. Reviews of Modern Physics, 36(1): 277–280, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.277;   http://
rmp.aps.org/abstract/RMP/v36/i1/p277_1.

Snow:1961:MH

[SS61] G. A. Snow and M. M. Shapiro. Mesons and hyperons. Reviews of Modern Physics, 33(2):231–238, April 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.231;   http://rmp.aps.org/abstract/RMP/
v33/i2/p231_1.

Smolinsky:1963:TCS

[SSW63] G. Smolinsky, L. C. Snyder, and E. Wasserman. A theoretical correlation of substituent effects in ground-state triplet nitrene EPR spectra. Reviews of Modern Physics, 35(3):576–577, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.576;   http://rmp.aps.org/abstract/RMP/
v35/i3/p576_1.

Stecki:1965:AEB

[ST65] J. Stecki and H. S. Taylor. On the areas of equivalence of the Bogoliubov theory and the prigogine theory of irreversible processes in classical gases. Reviews of Modern Physics, 37 (4):762–773, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.762;   http://
rmp.aps.org/abstract/RMP/v37/i4/p762_1.

Stapp:1962:AMT

[Sta62] Henry P. Stapp. Axiomatic S-matrix theory. Reviews of Modern Physics, 34(3):390–394, July 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.390;   http:
//rmp.aps.org/abstract/RMP/v34/i3/p390_1.

Stewartson:1960:MBT

[Ste60] K. Stewartson. Motion of bodies through conducting fluids. Reviews of Modern Physics, 32(4):855–859, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.855;   http://rmp.aps.org/abstract/RMP/
v32/i4/p855_1.

Stoneham:1969:SIB

[Sto69] A. M. Stoneham. Shapes of inhomogeneously broadened resonance lines in solids. Reviews of Modern Physics, 41(1): 82–108, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.41.82;   http://
rmp.aps.org/abstract/RMP/v41/i1/p82_1.

Stromgren:1964:COT

[Str64] Bengt Strömgren. Comparison of observed and theoretically calculated intensities in the continuous spectra of main-sequence B stars. Reviews of Modern Physics, 36(2):532–545, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.532;   http://rmp.aps.org/abstract/RMP/
v36/i2/p532_1.

Strocchi:1966:CCQ

[Str66] F. Strocchi. Complex coordinates and quantum mechanics. Reviews of Modern Physics, 38(1):36–40, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
38.36;   http://rmp.aps.org/abstract/RMP/v38/i1/p36_1.

Sheldon:1966:CIN

[SV66] Eric Sheldon and Douglas M. Van Patter. Compound inelastic nucleon and gamma-ray angular distributions for evenand odd-mass nuclei. Reviews of Modern Physics, 38(1): 143–186, January 1966. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.38.143;   http://
rmp.aps.org/abstract/RMP/v38/i1/p143_1.

Sauer:1960:TPN

[SW60] J. A. Sauer and A. E. Woodward. Transitions in polymers by nuclear magnetic resonance and dynamic mechanical methods. Reviews of Modern Physics, 32(1):88–101, January 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
32.88;   http://rmp.aps.org/abstract/RMP/v32/i1/p88_1.

Swan:1965:SDT

[Swa65] P. Swan. Sequential decay theory for the B^{11}(p, 3α) reaction. Reviews of Modern Physics, 37(3):336–345, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.336;   http://rmp.aps.org/abstract/RMP/
v37/i3/p336_1.

Talmi:1962:EIC

[Tal62] Igal Talmi. Effective interactions and coupling schemes in nuclei. Reviews of Modern Physics, 34(4):704–722, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.704;   http://rmp.aps.org/abstract/RMP/
v34/i4/p704_1.

Tamura:1965:ASN

[Tam65] Taro Tamura. Analyses of the scattering of nuclear particles by collective nuclei in terms of the coupled-channel calculation. Reviews of Modern Physics, 37(4):679–708, October 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.679;   http://rmp.aps.org/abstract/RMP/
v37/i4/p679_1.

Tamagaki:1967:NNI

[Tam67] Ryozo Tamagaki. Nucleon-nucleon interaction at small distances. Reviews of Modern Physics, 39(3):629–634, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.629;   http://rmp.aps.org/abstract/RMP/
v39/i3/p629_1.

Tatsumi:1960:ESM

[Tat60] T. Tatsumi. Energy spectra in magneto-fluid dynamic turbulence. Reviews of Modern Physics, 32(4):807–812, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.807;   http://rmp.aps.org/abstract/RMP/
v32/i4/p807_1.

Tayler:1960:STM

[Tay60] R. J. Tayler. Stability of twisted magnetic fields in a fluid of finite electrical conductivity. Reviews of Modern Physics, 32 (4):907–913, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.907;   http://
rmp.aps.org/abstract/RMP/v32/i4/p907_1.

Tayler:1961:EPI

[Tay61] R. J. Tayler. Errata: Proceedings of the International Symposium on Magneto-Fluid Dynamics [Rev. Mod. Phys 32 (October 1960), entire issue]. Reviews of Modern Physics, 33 (2):353–354, April 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.353;   http://
rmp.aps.org/abstract/RMP/v33/i2/p353_1.

Taylor:1969:SVR

[TB69] Raymond L. Taylor and Steven Bitterman. Survey of vibrational relaxation data for processes important in the CO_{2}–N_{2} laser system. Reviews of Modern Physics, 41(1):26–47, January 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.41.26;   http://rmp.aps.org/abstract/
RMP/v41/i1/p26_1.

Teller:1962:SMT

[Tel62] Edward Teller. On the stability of molecules in the Thomas–Fermi theory. Reviews of Modern Physics, 34(4):627–631, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.627;   http://rmp.aps.org/abstract/RMP/
v34/i4/p627_1.

Thompson:1960:LRF

[TH60] W. B. Thompson and J. Hubbard. Long-range forces and the diffusion coefficients of a plasma. Reviews of Modern Physics, 32(4):714–718, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.714;   http://
rmp.aps.org/abstract/RMP/v32/i4/p714_1.

Tsuji:1960:BAL

[THH60] Mikio Tsuji, Sigeru Huzinaga, and Tasuke Hasino. Bond alternation in long polyenes. Reviews of Modern Physics, 32 (2):425–427, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.425;   http://
rmp.aps.org/abstract/RMP/v32/i2/p425_1.

Thomas:1963:STA

[Tho63] L. H. Thomas. Statistical theory of atoms. Reviews of Modern Physics, 35(3):508–510, July 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.508;   http:
//rmp.aps.org/abstract/RMP/v35/i3/p508_1.

Thorndike:1967:EAN

[Tho67] Edward H. Thorndike. Experimental aspects of nucleon-nucleon scattering near 210 MeV — a critical review. Reviews of Modern Physics, 39(3):513–522, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.513;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p513_1.

Thompson:1968:ILP

[Tho68a] J. C. Thompson. Introduction and list of participants. Reviews of Modern Physics, 40(4):673–676, October 1968. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.673;   http://rmp.aps.org/abstract/RMP/
v40/i4/p673_1.

Thompson:1968:MNT

[Tho68b] J. C. Thompson. Metal-nonmetal transition in metal-ammonia solutions. Reviews of Modern Physics, 40(4):704–710, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.40.704;   http://rmp.aps.org/abstract/RMP/
v40/i4/p704_1.

Tinkham:1964:CFQ

[Tin64] M. Tinkham. Consequences of fluxoid quantization in the transitions of superconducting films. Reviews of Modern Physics, 36(1):268–276, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.268;   http://
rmp.aps.org/abstract/RMP/v36/i1/p268_1.

Tisza:1963:CSP

[Tis63] Laszlo Tisza. The conceptual structure of physics. Reviews of Modern Physics, 35(1):151–184, January 1963. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.151;   http://rmp.aps.org/abstract/RMP/
v35/i1/p151_1.

Tolstoy:1963:TWS

[Tol63] Ivan Tolstoy. The theory of waves in stratified fluids including the effects of gravity and rotation. Reviews of Modern Physics, 35(1):207–230, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.207;   http://
rmp.aps.org/abstract/RMP/v35/i1/p207_1.

Toxen:1964:TDC

[Tox64] Arnold M. Toxen. Temperature dependence of the critical fields of thin superconducting films. Reviews of Modern Physics, 36 (1):308–312, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.308;   http://
rmp.aps.org/abstract/RMP/v36/i1/p308_1.

Taylor:1969:DUM

[TPL69] B. N. Taylor, W. H. Parker, and D. N. Langenberg. Determination of e/h, using macroscopic quantum phase coherence in superconductors: Implications for quantum electrodynamics and the fundamental physical constants. Reviews of Modern Physics, 41(3):375–496, July 1969. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.41.375;   http:
//rmp.aps.org/abstract/RMP/v41/i3/p375_1.  See errata

[TPL73].

Taylor:1973:EDU

[TPL73] B. N. Taylor, W. H. Parker, and D. N. Langenberg. Erratum: Determination of e/h, Using Macroscopic Quantum Phase Coherence in Superconductors: Implications for Quantum Electrodynamics and the Fundamental Physical Constants. Reviews of Modern Physics, 45(1):109, January 1973. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.45.109;   http://rmp.aps.org/abstract/RMP/
v45/i1/p109_1. See [TPL69].

Tutsch:1968:CTB

[Tut68] Jerald H. Tutsch. Collapse time for the Bohm–Bub Hidden Variable Theory. Reviews of Modern Physics, 40(1): 232–234, January 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.232;   http://
rmp.aps.org/abstract/RMP/v40/i1/p232_1.

Thomas:1963:PVM

[TW63] G. Thomas and J. Washburn. Precipitation of vacancies in metals. Reviews of Modern Physics, 35(4):992–1011, October 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.992;   http://rmp.aps.org/abstract/RMP/
v35/i4/p992_1.

Udo:1965:DFD

[Udo65] F. Udo. Disintegration in flight of deuterons under the influence of nuclear fields. Reviews of Modern Physics, 37 (3):365–368, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.37.365;   http://
rmp.aps.org/abstract/RMP/v37/i3/p365_1.

VanVleck:1962:NIB

[Van62] J. H. Van Vleck. Note on the interactions between the spins of magnetic ions or nuclei in metals. Reviews of Modern Physics, 34(4):681–686, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.681;   http://
rmp.aps.org/abstract/RMP/v34/i4/p681_1.

VanHove:1964:HEC

[Van64] L. Van Hove. High-energy collisions of strongly interacting particles. Reviews of Modern Physics, 36(2):655–665, April 1964.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.655;   http://rmp.aps.org/abstract/RMP/
v36/i2/p655_1.

Vasvari:1968:EBF

[Vas68] B. Vasvári. Energy band and Fermi surface of divalent metals under pressure. Reviews of Modern Physics, 40(4):776– 781, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.40.776;   http://
rmp.aps.org/abstract/RMP/v40/i4/p776_1.

Verleur:1968:OPV

[VBB68] Hans W. Verleur, A. S. Barker, Jr., and C. N. Berglund. Optical properties of VO_{2} between 0.25 and 5 eV. Reviews of Modern Physics, 40(4):737, October 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.737;   http:
//rmp.aps.org/abstract/RMP/v40/i4/p737_1.

Vestine:1960:PMA

[Ves60] E. H. Vestine. Polar, magnetic, auroral, and ionospheric phenomena. Reviews of Modern Physics, 32(4):1020–1025, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1020;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1020_1.

VanKranendonk:1968:TRV

[VK68] J. Van Kranendonk and G. Karl. Theory of the rotational and vibrational excitations in solid parahydrogen, and frequency analysis of the infrared and Raman spectra. Reviews of Modern Physics, 40(3):531–555, July 1968. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.40.531;   http:
//rmp.aps.org/abstract/RMP/v40/i3/p531_1.

Vanderslice:1960:QMC

[VM60] Joseph T. Vanderslice and Edward A. Mason. Quantum mechanical calculations of short-range intermolecular forces. Reviews of Modern Physics, 32(2):417–421, April 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.417;   http://rmp.aps.org/abstract/RMP/
v32/i2/p417_1.

Vogt:1962:TLE

[Vog62] Erich Vogt. Theory of low energy nuclear reactions. Reviews of Modern Physics, 34(4):723–747, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.723;   http://rmp.aps.org/abstract/RMP/
v34/i4/p723_1.

Varshni:1963:AHM

[VS63] Yatendra Pal Varshni and Ramesh Chandra Shukla. Alkali hydride molecules: Potential energy curves and the nature of their binding. Reviews of Modern Physics, 35(1): 130–144, January 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.35.130;   http://
rmp.aps.org/abstract/RMP/v35/i1/p130_1.

Wada:1964:EQD

[Wad64] Yasushi Wada. The effect of quasiparticle damping on the ratio between the energy gap and the transition temperature of lead. Reviews of Modern Physics, 36(1):253–257, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.253;   http://rmp.aps.org/abstract/RMP/
v36/i1/p253_1.

Waldram:1964:SIS

[Wal64] J. R. Waldram. Surface impedance of superconductors. Reviews of Modern Physics, 36(1):187–193, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.187;   http://rmp.aps.org/abstract/RMP/
v36/i1/p187_1.

Wallace:1965:LDE

[Wal65] Duane C. Wallace. Lattice dynamics and elasticity of stressed crystals. Reviews of Modern Physics, 37(1):57–67, January 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.37.57;   http://rmp.aps.org/abstract/
RMP/v37/i1/p57_1.

Walker:1967:RPP

[Wal67] W. D. Walker. A review of pi–pi phase shifts. Reviews of Modern Physics, 39(3):695–699, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.39.695;   http:
//rmp.aps.org/abstract/RMP/v39/i3/p695_1.

Wannier:1962:DBE

[Wan62] Gregory H. Wannier. Dynamics of band electrons in electric and magnetic fields. Reviews of Modern Physics, 34(4): 645–655, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.34.645;   http://
rmp.aps.org/abstract/RMP/v34/i4/p645_1.

Ward:1965:CNO

[War65] J. F. Ward. Calculation of nonlinear optical susceptibilities using diagrammatic perturbation theory. Reviews of Modern Physics, 37(1):1–18, January 1965. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
37.1;   http://rmp.aps.org/abstract/RMP/v37/i1/p1_1.

Warren:1968:FCS

[War68] J. L. Warren. Further considerations on the symmetry properties of the normal vibrations of a crystal. Reviews of Modern Physics, 40(1):38–76, January 1968. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
40.38;   http://rmp.aps.org/abstract/RMP/v40/i1/p38_1.

Wasserman:1960:TOA

[Was60] Edel Wasserman. Theory of the optical activity of bimesityl derivatives. Reviews of Modern Physics, 32(2):443– 444, April 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.443;   http://
rmp.aps.org/abstract/RMP/v32/i2/p443_1.

Weber:1959:M

[Web59] J. Weber. Masers. Reviews of Modern Physics, 31(3): 681–710, July 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.31.681;   http:
//rmp.aps.org/abstract/RMP/v31/i3/p681_1. See erratum

[Web60].

Weber:1960:EM

[Web60] J. Weber. Erratum: Masers. Reviews of Modern Physics, 32(4): 1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.32.1033;   http://rmp.aps.org/
abstract/RMP/v32/i4/p1033_2. See [Web59].

Weger:1964:EBS

[Weg64] M. Weger. The electronic band structure of V_{3}Si and V_{3}Ga. Reviews of Modern Physics, 36(1):175–177, January 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.175;   http://rmp.aps.org/abstract/RMP/
v36/i1/p175_1.

Waggoner:1965:RP

[WEHM65] M. A. Waggoner, J. E. Etter, H. D. Holmgren, and C. Moazed. The Be^{10} (He^{3}, ααα) reaction process. Reviews of Modern Physics, 37(3):358–361, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.358;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p358_1.

Weidenmuller:1961:FFB

[Wei61] Hans A. Weidenmüller. First-forbidden beta decay. Reviews of Modern Physics, 33(4):574–607, October 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.574;   http://rmp.aps.org/abstract/RMP/
v33/i4/p574_1.

Weinstein:1963:DUL

[Wei63] A. Weinstein. Discussion on upper and lower bounds for expectation values. Reviews of Modern Physics, 35(3):715–716, July 1963.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.35.715;   http://rmp.aps.org/abstract/RMP/
v35/i3/p715_1.

Werthamer:1964:TLS

[Wer64] N. R. Werthamer. Theory of a local superconductor in a magnetic field. Reviews of Modern Physics, 36(1):292–294, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.292;   http://rmp.aps.org/abstract/RMP/
v36/i1/p292_1.

Wetherell:1961:RWS

[Wet61] A. M. Wetherell. Recent work on strong interactions at CERN. Reviews of Modern Physics, 33(3):382–389, July 1961. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.33.382;   http://rmp.aps.org/abstract/RMP/
v33/i3/p382_1.

Wheeler:1961:GPM

[Whe61] John A. Wheeler. Geometrodynamics and the problem of motion. Reviews of Modern Physics, 33(1):63–78, January 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
33.63;   http://rmp.aps.org/abstract/RMP/v33/i1/p63_1.

Wheeler:1962:PFB

[Whe62] John Archibald Wheeler. Problems on the frontiers between General Relativity and differential geometry. Reviews of Modern Physics, 34(4):873–892, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.34.873;   http:
//rmp.aps.org/abstract/RMP/v34/i4/p873_1.

Whitehead:1967:PFW

[Whi67] C. Whitehead. Present and future work with the Harwell synchrocyclotron. Reviews of Modern Physics, 39(3):538–541, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.538;   http://rmp.aps.org/abstract/RMP/
v39/i3/p538_1.

Wightman:1962:LQM

[Wig62] A. S. Wightman. On the localizability of quantum mechanical systems. Reviews of Modern Physics, 34(4):845–872, October 1962. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.845;   http://rmp.aps.org/abstract/RMP/
v34/i4/p845_1.

Wilcox:1959:RHT

[Wil59] John M. Wilcox. Review of high-temperature rotating-plasma experiments. Reviews of Modern Physics, 31(4):1045–1051, October 1959. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.31.1045;   http://rmp.aps.org/
abstract/RMP/v31/i4/p1045_1. See comment [Sle60] and reply

[Wil60].

Wilcox:1960:RCR

[Wil60] John M. Wilcox. Reply to comment on “Review of High-Temperature Rotating-Plasma Experiments”. Reviews of Modern Physics, 32(4):1033, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.32.1033;
http://rmp.aps.org/abstract/RMP/v32/i4/p1033_1.  See

[Wil59, Sle60].

Williams:1964:EFH

[Wil64] Robert W. Williams. Empirical formula for high-energy nuclear cross sections. Reviews of Modern Physics, 36(3): 815–819, July 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.815;   http://
rmp.aps.org/abstract/RMP/v36/i3/p815_1.

Wipf:1964:ISE

[Wip64] S. L. Wipf. The interphase surface energy in pure and impure tantalum. Reviews of Modern Physics, 36(1):83–86, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.36.83;   http://rmp.aps.org/abstract/
RMP/v36/i1/p83_1.

Woltjer:1960:THE

[Wol60] L. Woltjer. On the theory of hydromagnetic equilibrium. Reviews of Modern Physics, 32(4):914–915, October 1960. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.32.914;   http://rmp.aps.org/abstract/RMP/
v32/i4/p914_1.

Wong:1967:MTN

[Won67] David Y. Wong. Meson theory of nucleon-nucleon interaction. Reviews of Modern Physics, 39(3):622–628, July 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.622;   http://rmp.aps.org/abstract/RMP/
v39/i3/p622_1.

Wolf:1961:CLK

[WSL^{+}61] Sanford E. Wolf, Norbert Schmitz, Lester J. Lloyd, William Laskar, Frank S. Crawford, Jr., Janice Button, Jared A. Anderson, and Gideon Alexander. The cusp in lambda–K production at sigma–K threshold. Reviews of Modern Physics, 33(3):439–447, July 1961. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.33.439;   http://
rmp.aps.org/abstract/RMP/v33/i3/p439_1.

Wu:1964:UFI

[Wu64] C. S. Wu. The universal Fermi interaction and the conserved vector current in beta decay. Reviews of Modern Physics, 36 (2):618–632, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.36.618;   http://
rmp.aps.org/abstract/RMP/v36/i2/p618_1.

Wyder:1964:MFP

[Wyd64] P. Wyder. Mean free paths in the normal and the superconducting state. Reviews of Modern Physics, 36(1):116–118, January 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.116;   http://rmp.aps.org/abstract/RMP/
v36/i1/p116_1.

Yang:1962:CDL

[Yan62] C. N. Yang. Concept of off-diagonal long-range order and the quantum phases of liquid He and of superconductors. Reviews of Modern Physics, 34(4):694–704, October 1962. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.34.694;   http://rmp.aps.org/abstract/RMP/
v34/i4/p694_1.

Yuasa:1965:QES

[YH65] Toshiko Yuasa and Eid Hourani. Quasi-elastic p–α scattering in C^{12}. Reviews of Modern Physics, 37(3):399–401, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.399;   http://rmp.aps.org/abstract/RMP/
v37/i3/p399_1.

Young:1965:LR

[YJE^{+}65] F. C. Young, K. S. Jayaraman, J. E. Etter, H. D. Holmgren, and M. A. Waggoner. The Li^{6} (He^{3}, pαα) reaction. Reviews of Modern Physics, 37(3):362–364, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.37.362;   http:
//rmp.aps.org/abstract/RMP/v37/i3/p362_1.

Zernik:1967:TIO

[Zer67] Wolfgang Zernik. Theory of infrared optical properties of fluorite crystals. Reviews of Modern Physics, 39(2):432–439, April 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.39.432;   http://rmp.aps.org/abstract/RMP/
v39/i2/p432_1.

Zhigulev:1960:SMF

[Zhi60] V. N. Zhigulev. Some magneto-fluid dynamic effects in a finitely conducting medium. Reviews of Modern Physics, 32 (4):828–830, October 1960. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.32.828;   http://
rmp.aps.org/abstract/RMP/v32/i4/p828_1.

Zupancic:1965:SSM

[Zup65] C. Zupancic. A survey of simple mechanisms of three-particle reactions. Reviews of Modern Physics, 37(3):330–335, July 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.37.330;   http://rmp.aps.org/abstract/RMP/
v37/i3/p330_1.